### v1

# Majorization via positivity of Jack and Macdonald polynomial differences

Abstract.
Majorization inequalities have a long history, going back to Maclaurin and Newton. They were recently studied for several families of symmetric functions, including by Cuttler–Greene–Skandera (2011), Sra (2016), Khare–Tao (2021), McSwiggen–Novak (2022), and Chen–Sahi (2024+) among others. Here we extend the inequalities by these authors to Jack and Macdonald polynomials, and obtain conjectural characterizations of majorization and of weak majorization of the underlying partitions. We prove these characterizations for several cases of partitions, including all partitions with two parts. In fact, we upgrade – and prove in the above cases – the characterization of majorization, to containment of Jack and Macdonald differences lying in the Muirhead semiring.
Key words and phrases:
Majorization, weak majorization, Jack polynomials, Macdonald polynomials, Schur polynomials, Muirhead semiring
2020 Mathematics Subject Classification:
Primary 05E05, 33D52; Secondary 06A07, 26D05
1. Introduction
In this section, we first recall some of the history of symmetric functions and of inequalities associated to them.
1.1. Symmetric functions
Symmetric polynomials and symmetric functions have a long history and are ubiquitous in mathematics and physics, especially in algebra and representation theory, combinatorics, probability, and statistics. We give a very brief history of the symmetric functions below. For more details, see [14].
One of the most classical results is the relation between the elementary symmetric polynomials and the power sums, studied by Newton around 1666. These relations, together with Vieta’s formulas, enable one to find the power sums of the roots of a univariate polynomial, without finding the roots explicitly.
Since the work of Jacobi in the mid-nineteenth century and the papers of Frobenius, Schur, Weyl, MacMahon, and Young in the early twentieth century, Schur polynomials have been well studied, and play a significant role in the representation theory of the symmetric group $S_{n}$ and the complex general linear group $GL_{n}(\mathbb{C})$.
Around 1960, Hall and Littlewood independently introduced a one-parameter generalization of the Schur functions, now known as the Hall–Littlewood polynomials. By the work of Green and Macdonald, these functions are related to the representation theory of $GL_{n}$ over finite and $p$-adic fields.
In the late 1960s, Jack discovered Jack polynomials, unifying Schur polynomials and zonal polynomials – the latter being related to the representation theory of $GL_{n}(\mathbb{R})$ and multivariate statistics.
Hall–Littlewood polynomials and Jack polynomials are quite different generalizations of Schur polynomials. In the 1980s, Macdonald unified these developments into a two-parameter family of symmetric polynomials, the Macdonald polynomials. Hall–Littlewood polynomials can be obtained from Macdonald polynomials by specializing one of the parameters to 0; while Jack polynomials arise as the limiting case when both parameters approach 1.
1.2. Symmetric function inequalities
The notion of positivity has an even longer history: it begin with mathematics itself – counting and measuring objects. Let us restrict our attention to inequalities about symmetric functions.
Most families of symmetric polynomials are indexed by integer partitions. An integer partition (of length at most $n$) is an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of integers such that $\lambda_{1}\geqslant\cdots\geqslant\lambda_{n}\geqslant 0$. Denote the set of these partitions by $\mathcal{P}_{n}$.
Recall that for partitions $\lambda$ and $\mu$ in $\mathcal{P}_{n}$, one has the following partial orders:
  • $\lambda$ contains $\mu$, denoted by $\lambda\supseteq\mu$, if $\lambda_{i}\geqslant\mu_{i}$ for $i=1,\dots,n$;
  • $\lambda$ weakly majorizes $\mu$, denoted by $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, if $\sum_{i=1}^{k}\lambda_{i}\geqslant\sum_{i=1}^{k}\mu_{i}$, for $k=1,\dots,n$;
  • $\lambda$ majorizes $\mu$, denoted by $\lambda\mathrel{\succcurlyeq}\mu$, if $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $|\lambda|=|\mu|$.
Throughout the paper, $\bm{1}=\bm{1}_{n}\coloneqq(1,\dots,1)$ ($n$ times).
We now recall some symmetric function inequalities related to the three partial orders. The notation used here and below is found in Section 2.
1.2.1. Containment-type and expansion positivity
Of great importance in representation theory and algebraic combinatorics is the notion of expansion positivity.
We say a symmetric polynomial is Schur positive if it can be written as a sum of Schur polynomials with non-negative (integer) coefficients. For example, by the celebrated Littlewood–Richardson rule, the product of two Schur polynomials is Schur positive: $s_{\lambda}\cdot s_{\mu}=\sum_{\nu}c_{\lambda\mu}^{\nu}s_{\nu}$, where the Littlewood–Richardson coefficients $c_{\lambda\mu}^{\nu}\geqslant 0$. Similarly, the Schur polynomial $s_{\lambda}$ is monomial positive: $s_{\lambda}=\sum_{\mu\mathrel{\succcurlyeq}\lambda}K_{\lambda\mu}m_{\mu}$, where the Kostka numbers $K_{\lambda\mu}\geqslant 0$.
Expansion positivity naturally suggests underlying combinatorial objects or structure. For example, Littlewood–Richardson coefficients have various combinatorial interpretations: Littlewood–Richardson tableaux [27, 18], Berenstein–Zelevinsky patterns [3] and Knutson–Tao honeycombs [13]; and Kostka numbers count the number of semi-standard Young tableaux (SSYT) [27, 18].
For Jack polynomials $P_{\lambda}^{(\alpha)}(x)$, it was conjectured in Macdonald’s book [18, VI.(10.26?)] and proved by Knop and Sahi [12] that their expansions in terms of the monomial basis $(m_{\lambda})$ are positive (and integral, under certain normalizations) in the parameter $\alpha$.
The following binomial formula for Jack polynomials is due to Okounkov and Olshanski [21]:
$\frac{P_{\lambda}^{(\alpha)}(x+\bm{1})}{P_{\lambda}^{(\alpha)}(\bm{1})}$ $=\sum_{\nu\subseteq\lambda}\binom{\lambda}{\nu}_{\alpha}\ \frac{P_{\nu}^{(\alpha)}(x)}{P_{\nu}^{(\alpha)}(\bm{1})}.$ (1.1)
The coefficient $\binom{\lambda}{\nu}_{\alpha}$ is known as the generalized binomial coefficient. Sahi [25] showed that $\binom{\lambda}{\nu}_{\alpha}$ is non-negative (i.e. lies in a positive cone $\mathbb{F}_{\geqslant 0}$ defined in Eq. 2.11) and hence $\frac{P_{\lambda}^{(\alpha)}(x+\bm{1})}{P_{\lambda}^{(\alpha)}(\bm{1})}$ is Jack positive. Recently, Chen and Sahi proved [4] that the coefficient $\binom{\lambda}{\nu}_{\alpha}$ is monotone in $\lambda$:
$\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\in\mathbb{F}_{\geqslant 0},\quad\lambda\supseteq\mu$ (1.2)
which, in turn, leads to the Jack positivity of $\frac{P_{\lambda}^{(\alpha)}(x+\bm{1})}{P_{\lambda}^{(\alpha)}(\bm{1})}-\frac{P_{\mu}^{(\alpha)}(x+\bm{1})}{P_{\mu}^{(\alpha)}(\bm{1})}$ for $\lambda\supseteq\mu$:
$\frac{P_{\lambda}^{(\alpha)}(x+\bm{1})}{P_{\lambda}^{(\alpha)}(\bm{1})}-\frac{P_{\mu}^{(\alpha)}(x+\bm{1})}{P_{\mu}^{(\alpha)}(\bm{1})}=\sum_{\nu\subseteq\lambda}\left(\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\right)\frac{P_{\nu}^{(\alpha)}(x)}{P_{\nu}^{(\alpha)}(\bm{1})}.$ (1.3)
The study of generalized binomial coefficients originated in two distinct contexts. In statistics, their zonal case ($\alpha=2$) was first examined by Constantine in the 1960s, while Lascoux [15] explored their Schur case ($\alpha=1$) to compute Chern classes for exterior and symmetric squares of vector bundles (see also [18, p. 47, Example 10]). Seminal contributions emerged in the 1990s through the foundational work of Lassalle [16, 17], Kaneko [9], and Okounkov–Olshanski [22, 21].
Generalized binomial coefficients can be explicitly computed via evaluating interpolation Jack polynomials, first studied by Knop–Sahi [24, 11] and Okounkov–Olshanski [21]. Recently, the $BC$-symmetry analogue of binomial coefficients and interpolation polynomials were studied in [4].
In what follows, we write Jack polynomials with parameter $\tau=1/\alpha$, that is, our $P_{\lambda}(x;\tau)$ is equal to Macdonald’s $P_{\lambda}^{(1/\tau)}(x)$.
Theorem 1.1 ([4, Theorem 6.1]).
The following statements are equivalent:
  1. (1)
    $\lambda$ contains $\mu$, $\lambda\supseteq\mu$.
  2. (2)
    For any fixed $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
  3. (3)
    The difference $\frac{P_{\lambda}(x+\bm{1};\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x+\bm{1};\tau)}{P_{\mu}(\bm{1};\tau)}$ is Jack positive over $\mathbb{F}_{\geqslant 0}$, namely, can be written as an $\mathbb{F}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau)$.
It is well-known, see [18, Chapters I, VI, VII], that Jack polynomials $P_{\lambda}(x;\tau)$ specialize to many other symmetric polynomials: monomial symmetric polynomials $m_{\lambda}$ when $\tau=0$, zonal polynomials $Z_{\lambda}$ when $\tau=1/2$ (spherical functions for $GL_{n}(\mathbb{R})/O(n)$) and $\tau=2$ (for $GL_{n}(\mathbb{H})/Sp(n)$), Schur polynomials $s_{\lambda}$ when $\tau=1$, and elementary symmetric polynomials $e_{\lambda^{\prime}}$ when $\tau=\infty$ (where $\lambda^{\prime}$ is the conjugate of $\lambda$). Hence, we also have the following characterization, which somewhat surprisingly also holds for power-sums:
Theorem 1.2 ([4, Theorems 6.2, 6.4]).
The following statements are equivalent:
  1. (1)
    $\lambda$ contains $\mu$.
  2. (2)
    The difference $\frac{m_{\lambda}(x+\bm{1})}{m_{\lambda}(\bm{1})}-\frac{m_{\mu}(x+\bm{1})}{m_{\mu}(\bm{1})}$ is monomial positive.
  3. (3)
    The difference $\frac{Z_{\lambda}(x+\bm{1})}{Z_{\lambda}(\bm{1})}-\frac{Z_{\mu}(x+\bm{1})}{Z_{\mu}(\bm{1})}$ is zonal positive.
  4. (4)
    The difference $\frac{s_{\lambda}(x+\bm{1})}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x+\bm{1})}{s_{\mu}(\bm{1})}$ is Schur positive.
  5. (5)
    The difference $\frac{e_{\lambda}^{\prime}(x+\bm{1})}{e_{\lambda}^{\prime}(\bm{1})}-\frac{e_{\mu}^{\prime}(x+\bm{1})}{e_{\mu}^{\prime}(\bm{1})}$ is elementary positive.
  6. (6)
    The difference $\frac{p_{\lambda}(x+\bm{1})}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}(x+\bm{1})}{p_{\mu}(\bm{1})}$ is power sum positive.
1.2.2. Majorization-type and evaluation positivity
As the name suggests, evaluation positivity is about evaluating real polynomials, which is more commonly seen in classical analysis. We begin with the special case of majorization: $(k+1,k-1)\mathrel{\succcurlyeq}(k,k)$ for $k\geqslant 1$.
It is a classical result, dating back to Newton, that polynomials with non-positive real roots have ultra log-concave coefficients. More precisely, for
$f(t)=\prod_{i=1}^{n}(t+x_{i})=\sum_{k=0}^{n}e_{k}(x)\,t^{n-k},$
Newton’s inequality asserts
$\frac{e_{k+1}(x)}{e_{k+1}(\bm{1})}\cdot\frac{e_{k-1}(x)}{e_{k-1}(\bm{1})}\leqslant\left(\frac{e_{k}(x)}{e_{k}(\bm{1})}\right)^{2},\qquad x\in[0,\infty)^{n}.$ (1.4)
Equivalently, the normalized elementary symmetric polynomials $a_{k}={e_{k}(x)}/{e_{k}(\bm{1})}$ form a log-concave sequence: $a_{k+1}a_{k-1}\leqslant a_{k}^{2}$. A familiar instance of log-concavity is given by the binomial coefficients $\binom{n}{0},\binom{n}{1},\dots,\binom{n}{n}$. Log-concavity also plays a central role in combinatorial Hodge theory, as developed by Adiprasito–Huh–Katz [1].
In 1959, long after Newton, Gantmacher [6] proved a similar inequality for power sums:
$p_{k+1}(x)p_{k-1}(x)\geqslant p_{k}(x)^{2},\quad x\in[0,\infty)^{n}.$ (1.5)
Note that $p_{k}(\bm{1}_{n})=n$ for any $k$, and hence one can normalize Gantmacher’s inequality in the same way as Newton’s.
For partitions $\lambda$ and $\mu$ with $|\lambda|=|\mu|$, Muirhead showed in [20] that for the monomial symmetric polynomial $m_{\lambda}$
$\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}\geqslant\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})},\quad x\in(0,\infty)^{n}$ (1.6)
if and only if $\lambda\mathrel{\succcurlyeq}\mu$. (Note that (1.6) holds even when $\lambda\in\mathbb{R}^{n}$, in which case, $m_{\lambda}$ is defined as the orbit-sum of $x^{\eta}=x_{1}^{\eta_{1}}\cdots x_{n}^{\eta_{n}}$, where $\eta$ runs over the orbit $S_{n}\cdot\lambda$, acting by permutation.)
Inspired by the inequalities above (and some others), Cuttler, Greene, and Skandera [5] generalized the inequalities of Newton and Gantmacher to elementary symmetric polynomials and power sums indexed by partitions, and conjectured an inequality for Schur polynomials, which was later proved by Sra [26] (again, for $\lambda,\mu$ real $n$-tuples). We recall all these results.
Theorem 1.3.
Let $\lambda$ and $\mu$ be integer partitions with $|\lambda|=|\mu|$. The following are equivalent:
  1. (1)
    $\lambda$ majorizes $\mu$, $\lambda\mathrel{\succcurlyeq}\mu$.
  2. (2)
    ([20], Muirhead’s inequality) The following difference is positive:
    $\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}-\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.7)
  3. (3)
    ([5, Theorem 3.2], generalized Newton’s inequality) The following difference is positive:
    $\frac{e_{\lambda^{\prime}}(x)}{e_{\lambda^{\prime}}(\bm{1})}-\frac{e_{\mu^{\prime}}(x)}{e_{\mu^{\prime}}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.8)
  4. (4)
    ([5, Theorem 4.2], generalized Gantmacher’s inequality) The following difference is positive:
    $\frac{p_{\lambda}(x)}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}(x)}{p_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.9)
  5. (5)
    ([5, Conjecture 7.4, Theorem 7.5] and [26], Cuttler–Greene–Skandera and Sra’s inequality) The following difference is positive:
    $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.10)
Since Jack polynomials and Macdonald polynomials generalize monomial symmetric, elementary symmetric, and Schur polynomials, the first and third authors made some similar conjectures in [4]. In this paper, we extend these conjectures and prove them in some cases.
As we will see below – as also in [5, 10, 19, 26] – the more challenging implication is to show that if $\lambda\mathrel{\succcurlyeq}\mu$ then the normalized differences of various symmetric polynomials are non-negative on $0,\infty)^{n}$. An alternate, stronger approach would be to show that these differences “decompose” non-negatively into non-negative quantities, like monomials and differences of monomials as in part (2). This is known for $e_{\lambda^{\prime}}$ and for $p_{\lambda}$, as we now state.
Let $M_{\lambda}\coloneqq m_{\lambda}/m_{\lambda}(\bm{1})$ be the normalized monomial. Recall [5, Section 6] that the Muirhead cone $\mathcal{M}_{C}(\mathcal{C})$ and the Muirhead semiring $\mathcal{M}_{S}(\mathcal{C})$ over a cone $\mathcal{C}$ are defined as follows: $\mathcal{M}_{C}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of Muirhead differences $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{M}_{S}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of products of functions in $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,M_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$. In particular, when evaluated at $x\in[0,\infty)^{n}$, functions in $\mathcal{M}_{C}(\mathcal{C})$ and $\mathcal{M}_{S}(\mathcal{C})$ take values in the cone $\mathcal{C}$.
Theorem 1.4 ([5, Theorem 6.1]).
Given partitions $\lambda\mathrel{\succcurlyeq}\mu$, the following differences
$\frac{e_{\lambda^{\prime}}(x)}{e_{\lambda^{\prime}}(\bm{1})}-\frac{e_{\mu^{\prime}}(x)}{e_{\mu^{\prime}}(\bm{1})},\quad\frac{p_{\lambda}(x)}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}(x)}{p_{\mu}(\bm{1})}$ (1.11)
lie in the Muirhead semiring $\mathcal{M}_{S}$ over $\mathbb{Q}_{\geqslant 0}$.
It should be noted that the Muirhead semiring $\mathcal{M}_{S}$ is strictly larger than the Muirhead cone $\mathcal{M}_{C}$. Some power sum differences belong to $\mathcal{M}_{S}\setminus\mathcal{M}_{C}$. For example, when $n=3$, $\lambda=(33)$ and $\mu=(321)$, we have
$\frac{p_{\lambda}}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}}{p_{\mu}(\bm{1})}=\frac{2}{3}M_{(3)}\left(M_{(3)}-M_{(21)}\right)=\frac{2}{9}\left(M_{(6)}-M_{(51)}-M_{(42)}+2M_{(33)}-M_{(321)}\right),$
which is in $\mathcal{M}_{S}$ but not in $\mathcal{M}_{C}$. However, it seems that the differences of elementary symmetric polynomials lie in the Muirhead cone $\mathcal{M}_{C}$.
1.2.3. Weak majorization-type and evaluation positivity
These inequalities (specifically, the final one in the next result) were first recorded in [10, Theorem 1.14] for Schur polynomials, followed by the other variants that were shown in [4].
Theorem 1.5 ([4, Theorem 6.5]).
The following statements are equivalent:
  1. (1)
    $\lambda$ weakly majorizes $\mu$, $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$.
  2. (2)
    The following difference is positive:
    $\frac{m_{\lambda}(x+\bm{1})}{m_{\lambda}(\bm{1})}-\frac{m_{\mu}(x+\bm{1})}{m_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.12)
  3. (3)
    The following difference is positive:
    $\frac{e_{\lambda^{\prime}}(x+\bm{1})}{e_{\lambda^{\prime}}(\bm{1})}-\frac{e_{\mu^{\prime}}(x+\bm{1})}{e_{\mu^{\prime}}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.13)
  4. (4)
    The following difference is positive:
    $\frac{p_{\lambda}(x+\bm{1})}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}(x+\bm{1})}{p_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.14)
  5. (5)
    (see also [10]) The following difference is positive:
    $\frac{s_{\lambda}(x+\bm{1})}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x+\bm{1})}{s_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.15)
As noted in [4, Theorem 6.5], these inequalities follow easily from their containment- and majorization- analogues, via the following easy lemma that connects the three partial orders.
Lemma 1.6 (see [4]).
If $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, then there exists some $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
As mentioned above, [10] provides a different proof of Theorem 1.5 for Schur polynomials.
1.3. Our work
This work was motivated by the goal of extending the above characterizations to Jack and even Macdonald polynomials. We refer the reader to:
  • Theorems 3.1, 3.2, 3.4 and 4.8, connecting (weak) majorization and normalized Jack differences;
  • Theorem 4.4 for the Muirhead semiring and Jack polynomial differences;
  • Theorems 5.6 and 5.8 for (weak) majorization via Macdonald polynomials.
2. Preliminaries
In this section, we recall the definitions of the symmetric functions that appear in this paper, mostly following [18, 4], with minor modifications.
2.1. Partitions
An (integer) partition is an infinite tuple $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of weakly-decreasing non-negative integers with finitely many nonzero entries. The non-zero $\lambda_{i}$’s are called the parts of $\lambda$. The zeros are usually omitted, and we use the exponential to indicate repeated entries. The length of $\lambda$, denoted by $\ell(\lambda)$, is the number of parts; and the size is the sum of the parts, $|\lambda|=\lambda_{1}+\cdots+\lambda_{\ell(\lambda)}$. For example, $\lambda=(2^{2}1^{3})=(2,2,1,1,1,0,\dots)$ has length 5 and size 7.
The conjugate of a partition $\lambda$ is denoted by $\lambda^{\prime}$, given by
$\lambda_{j}^{\prime}=\#\{\,i\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda_{i}\geqslant j\,\},\quad j\geqslant 1.$ (2.1)
Note that $\lambda^{\prime\prime}=\lambda$, and $\lambda_{1}^{\prime}=\ell(\lambda)$.
Denote by $\mathcal{P}_{n}$ the set of partitions of length at most $n$ (note, this differs from Macdonald’s notation), and $\mathcal{P}_{n}^{\prime}$ the set of the conjugates of $\mathcal{P}_{n}$, namely, the partitions $\lambda$ with $\lambda_{1}\leqslant n$.
2.2. Symmetric polynomials
Let $\Lambda_{n}$ be the algebra of symmetric polynomials in $n$ variables $x=(x_{1},\dots,x_{n})$ over the field $\mathbb{Q}$. The monomial symmetric polynomial is defined as
$m_{\lambda}(x)\coloneqq\sum_{\eta}x^{\eta},$ (2.2)
where $x^{\eta}\coloneqq x_{1}^{\eta_{1}}\cdots x_{n}^{\eta_{n}}$ and the sum is over distinct permutations $\eta$ of $\lambda=(\lambda_{1},\dots,\lambda_{n})$. Note that $m_{\lambda}$ is a well-defined function for any real $n$-tuple $\lambda$.
For any $k\geqslant 1$, the $k$-th elementary symmetric polynomial and power sum are defined by
$e_{k}(x)$ $\coloneqq m_{(1^{k})}(x)=\sum_{1\leqslant i_{1}<\dots<i_{k}\leqslant n}x_{i_{1}}\cdots x_{i_{k}},$ (2.3)
$p_{k}(x)$ $\coloneqq m_{(k)}(x)=\sum_{1\leqslant i\leqslant n}x_{i}^{k}.$ (2.4)
(If $k>n$, $e_{k}(x_{1},\dots,x_{n}):=0$.) Next, define $e_{\lambda}$ and $p_{\lambda}$ by multiplication:
$e_{\lambda}\coloneqq e_{\lambda_{1}}\cdots e_{\lambda_{l}},\quad p_{\lambda}\coloneqq p_{\lambda_{1}}\cdots p_{\lambda_{l}},\qquad l=\ell(\lambda).$ (2.5)
The Schur polynomial is defined by
$s_{\lambda}(x)=\frac{\det(x_{i}^{\lambda_{j}+n-j})_{1\leqslant i,j\leqslant n}}{\det(x_{i}^{n-j})_{1\leqslant i,j\leqslant n}}.$ (2.6)
Note that the denominator is the Vandermonde determinant
$\det(x_{i}^{n-j})_{1\leqslant i,j\leqslant n}=\prod_{1\leqslant i<j\leqslant n}(x_{i}-x_{j}),$ (2.7)
and that the numerator (and hence $s_{\lambda}$) is a well-defined function for any real $n$-tuple $\lambda$.
Each of $\{\,m_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,e_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,p_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ and $\{\,s_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ is a $\mathbb{Q}$-basis of $\Lambda_{n}$ [18, Chapter I].
2.3. Jack polynomials and Macdonald polynomials
Let $\alpha=\frac{1}{\tau}$, $q$ and $t$ be indeterminates over $\mathbb{Q}$ and consider the fields of rational functions $\mathbb{F}=\mathbb{Q}(\alpha)=\mathbb{Q}(\tau)$ and $\mathbb{F}=\mathbb{Q}(q,t)$. Let $\Lambda_{n,\tau}\coloneqq\Lambda_{n}\otimes\mathbb{Q}(\tau)$ and $\Lambda_{n,q,t}\coloneqq\Lambda_{n}\otimes\mathbb{Q}(q,t)$ be the algebras of symmetric polynomials over the larger fields.
Recall the $\tau$-Hall inner product over $\Lambda_{n,\tau}$ and the $q,t$-Hall inner product over $\Lambda_{n,q,t}$:
$\langle p_{\lambda},p_{\mu}\rangle_{\tau}\coloneqq\delta_{\lambda}z_{\lambda}\tau^{-\ell(\lambda)},\quad\langle p_{\lambda},p_{\mu}\rangle_{q,t}\coloneqq\delta_{\lambda}z_{\lambda}\prod_{i=1}^{\ell(\lambda)}\frac{1-q^{\lambda_{i}}}{1-t^{\lambda_{i}}},$ (2.8)
where $\lambda,\mu\in\mathcal{P}_{n}^{\prime}$, and $z_{\lambda}$ is a constant given in [18, p. 24]. Then the Jack polynomial $P_{\lambda}(\tau)$ and the Macdonald polynomial $P_{\lambda}(q,t)$ can be defined as the unique polynomials satisfying:
$\langle P_{\lambda}(*),P_{\mu}(*)\rangle_{*}=0,\quad\lambda\neq\mu,$ (2.9)
$P_{\lambda}(*)=\sum_{\lambda\mathrel{\succcurlyeq}\mu}K_{\lambda\mu}(*)m_{\mu},\quad K_{\lambda\lambda}=1;$ (2.10)
here * is $\tau$ for Jack polynomials and $(q,t)$ for Macdonald polynomials. Note that our Jack polynomial $P_{\lambda}(x;\tau)$ is equal to Macdonald’s $P_{\lambda}^{(\alpha=1/\tau)}(x)$.
We provide additional definitions later when computing the polynomials.
2.4. The cone of positivity
In the Jack case, we define the cone of positivity $\mathbb{F}_{\geqslant 0}$ and its real extension $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ as
$\mathbb{F}_{\geqslant 0}\coloneqq\left\{\,\frac{f}{g}\mathclose{}\nonscript\;\middle|\penalty 500\nonscript\;\mathopen{}f,g\in\mathbb{Z}_{\geqslant 0}[\tau],\ g\neq 0\,\right\},\quad\mathbb{F}_{\geqslant 0}^{\mathbb{R}}\coloneqq\left\{\,\frac{f}{g}\mathclose{}\nonscript\;\middle|\penalty 500\nonscript\;\mathopen{}f,g\in\mathbb{R}_{\geqslant 0}[\tau],\ g\neq 0\,\right\}.$ (2.11)
In the Macdonald case, we define the cone of positivity $\mathbb{F}_{\geqslant 0}$ and its real extension $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ as
$\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\},\quad\mathbb{F}_{\geqslant 0}^{\mathbb{R}}\coloneqq\{\,f\in\mathbb{R}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\}.$ (2.12)
3. Jack polynomial conjectures and their partial resolutions
Having explained the previous literature, we now come to the main thrust of the present work. Notice that Theorem 1.2 above is a “manifestation” of Theorem 1.1 for special real values of the parameter $\tau$. Given similar manifestations of (weak) majorization in Theorems 1.3 and 1.5, it is natural to expect the analogues of Theorem 1.1 to also hold for (weak) majorization. At the same time, one would expect Theorem 3.2 to also generalize. Thus, we begin with the following upgradation of [4, Conjecture 1(1)] to majorization. For reasons of exposition, we write the assertion that $\lambda\mathrel{\succcurlyeq}\mu$ at the end:
Conjecture 1 (CGS Conjecture for Jack polynomials).
Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
  1. (1)
    We have
    $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (3.1)
  2. (2)
    For some fixed $\tau_{0}\in[0,\infty]$, we have
    $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (3.2)
  3. (3)
    For some fixed $\tau_{0}\in[0,\infty]$, we have
    $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in(0,1)^{n}\cup(1,\infty)^{n}.$ (3.3)
  4. (4)
    $\lambda$ majorizes $\mu$.
In exact parallel, we next upgrade [4, Conjecture 1(2)] to weak majorization.
Conjecture 2 (KT Conjecture for Jack polynomials).
The following are equivalent for partitions $\lambda$ and $\mu$:
  1. (1)
    We have
    $\frac{P_{\lambda}(x+\bm{1};\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x+\bm{1};\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (3.4)
  2. (2)
    For some fixed $\tau_{0}\in[0,\infty]$, we have
    $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (3.5)
  3. (3)
    $\lambda$ weakly majorizes $\mu$.
The next result shows “most of” the implications in these conjectures, leaving one unproved implication.
Theorem 3.1.
[…]
3.1. The proof, using majorization via log-positive and log-negative sequences
Before showing Theorem 3.1, we take a closer look at Theorem 1.3. If $\lambda\mathrel{\succcurlyeq}\mu$, then the inequalities in the final four parts of the theorem hold over the entire orthant. However, to deduce $\lambda\mathrel{\succcurlyeq}\mu$, it suffices to work with just two sequences $x_{n}$ – one in the “log-positive” orthant $(1,\infty)^{n}$ and the other in the log-negative orthant $(0,1)^{n}$ – as opposed to all $x\in[0,\infty)^{n}$. This provides additional characterizations of (weak) majorization, and is the content of (the proof of) the next result.
Theorem 3.2.
If any of the final four inequalities in Theorem 1.3 holds for all $x$ in $(0,1)^{n}\cup(1,\infty)^{n}$, then $\lambda\mathrel{\succcurlyeq}\mu$. (The converse follows from Theorem 1.3.)
Proof.
The assertion for normalized Schur differences was shown in [10]. In a similar vein, we consider these inequalities for the monomial symmetric polynomials. Since $M_{\lambda}\geqslant M_{\mu}$ on $(1,\infty)^{n}$ (and these functions are continuous), it follows from Theorem 1.5 that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$. Specifically, if $\lambda$ does not weakly majorize $\mu$, then there is some $k$ such that $\sum_{i=1}^{k}\lambda_{i}<\sum_{i=1}^{k}\mu_{i}$. Now fix scalars $u_{i}>1$, and consider the sequence of vectors
$x(t)\coloneqq(tu_{1},\dots,tu_{k},u_{k+1},\dots,u_{n})\in(1,\infty)^{n}$
for a sequence $1\leqslant t=t_{m}\uparrow\infty$. Then as polynomials in $t$, $M_{\lambda}(x(t))$ has degree $\sum_{i=1}^{k}\lambda_{i}$, while $M_{\mu}(x(t))$ has degree $\sum_{i=1}^{k}\mu_{i}$, so $M_{\lambda}(x(t_{m}))-M_{\mu}(x(t_{m}))\to$-$\infty$ as $m\to\infty$, which is false.
Similarly, since $M_{\lambda}\geqslant M_{\mu}$ on $(0,1)^{n}$, we claim that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$. Indeed, if not, then there is some $k$ such that $\sum_{i=0}^{k-1}\lambda_{n-i}<-\sum_{i=0}^{k-1}\mu_{n-i}$. Now fix scalars $u_{i}\in(0,1)$ and consider the sequence $y(t)\coloneqq(u_{1}/t,\dots,u_{k}/t,u_{k+1},\dots,u_{n})\in(0,1)^{n}$ for a sequence $1\leqslant t=t_{m}\uparrow\infty$. Then
$M_{\lambda}(y(t))\sim O(t^{-\lambda_{n}-\cdots-\lambda_{n-k+1}}),\qquad M_{\mu}(y(t))\sim O(t^{-\mu_{n}-\cdots-\mu_{n-k+1}}),$ (3.6)
so $M_{\lambda}(y(t_{m}))-M_{\mu}(y(t_{m}))\to$-$\infty$ as $m\to\infty$, which is again false.
It follows that if $M_{\lambda}\geqslant M_{\mu}$ on two sequences $x(t_{m})\in(1,\infty)^{n}$ and $y(t_{m})\in(0,1)^{n}$, then $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$. But these are together equivalent to $\lambda\mathrel{\succcurlyeq}\mu$.
The same proof works if one considers power-sums, since $p_{\lambda}(x(t))$ is a polynomial in $t$ of degree $\sum_{i=1}^{k}\lambda_{i}$ as above, and the same claim as in Eq. 3.6 for $p_{\lambda}(y(t))/p_{\lambda}(\bm{1})$. Ditto for the “normalized (transposed) elementary symmetric polynomials” $e_{\lambda^{\prime}}(x)/e_{\lambda^{\prime}}(\bm{1})$. ∎
We can now prove Theorem 3.1.
Proof of Theorem 3.1.
[…]
Hence, to settle both conjectures it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1 and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ $\implies$ Eq. 3.4. We end by showing that the latter follows from the former. Indeed, by Section 1.2.3 there exists $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$. Now write (suppressing the $\tau$)
$\frac{P_{\lambda}(x+\bm{1})}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x+\bm{1})}{P_{\mu}(\bm{1})}=\left(\frac{P_{\lambda}(x+\bm{1})}{P_{\lambda}(\bm{1})}-\frac{P_{\nu}(x+\bm{1})}{P_{\nu}(\bm{1})}\right)+\left(\frac{P_{\nu}(x+\bm{1})}{P_{\nu}(\bm{1})}-\frac{P_{\mu}(x+\bm{1})}{P_{\mu}(\bm{1})}\right).$
The first difference is Jack-positive by Theorem 1.1, and in particular, in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ when evaluated at $x\in[0,\infty)^{n}$. The second difference is in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ by the assumption that Eq. 3.1 holds. ∎
Remark 3.3.
As a consequence of Theorem 3.1, it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1. Note that for $\tau_{0}=0,1,\infty$, this is already known by Theorem 1.3.
3.2. Reductions of the conjectures; proof for two variables
In this subsection, we reduce Conjecture 1 to some important special cases. We also work out an example, which shows, via the reductions, that the conjecture holds when $n\leqslant 2$:
Theorem 3.4.
Conjectures 1 and 2 hold when $\lambda$ and $\mu$ are of length at most 2.
We begin with some preliminary results and reductions, which hold for all $n$.
Proposition 3.5.
For a one-row partition $(d,0^{n-1})$, the Jack polynomials are given by
$P_{(d,0^{n-1})}(x;\tau)=\frac{d!}{(\tau)_{{d}}}\cdot\sum_{|\eta|=d}\frac{(\tau)_{{\eta}}}{\eta!}x^{\eta},$ (3.7)
where the sum runs over all $n$-tuples $\eta=(\eta_{1},\dots,\eta_{n})\in\mathbb{Z}_{\geqslant 0}^{n}$ such that $\eta_{1}+\dots+\eta_{n}=d$, and
$(\tau)_{{\eta}}\coloneqq\prod_{i}(\tau)_{{\eta_{i}}},\quad(x)_{{n}}\coloneqq x(x+1)\cdots(x+n-1)=\frac{\Gamma(x+n)}{\Gamma(x)}$ (3.8)
is the rising factorial. When evaluated at $\bm{1}$, we have $P_{(d,0^{n-1})}(\bm{1};\tau)=\frac{(n\tau)_{{d}}}{(\tau)_{{d}}}$.
Proof.
The first assertion follows from the combinatorial formula (see e.g. [4, Eq. (2.10)]). The second follows from [18, VI. (10.20)] (note, $\alpha$ is $\frac{1}{\tau}$ here), or the Chu–Vandermonde identity. ∎
Lemma 3.6.
Let $\lambda\in\mathcal{P}_{n}$ be such that $\lambda_{n}>0$. Then
$P_{\lambda}(x;\tau)=x_{1}\cdots x_{n}\cdot P_{\lambda-\bm{1}}(x;\tau).$ (3.9)
Proof.
See [18, VI. (4.17)] for the Macdonald case. ∎
Lemma 3.7 (Adding columns).
To prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1), it suffices to assume $\lambda_{n}=0$.
Proof.
Let $\lambda\mathrel{\succcurlyeq}\mu$, then $\lambda_{n}\leqslant\mu_{n}$. Clearly, $\lambda-\lambda_{n}\bm{1}\mathrel{\succcurlyeq}\mu-\lambda_{n}\bm{1}$, and by assumption, Eq. 3.1 holds for the pair $(\lambda-\lambda_{n}\bm{1},\mu-\lambda_{n}\bm{1})$. By Section 3.2, it holds for the pair $(\lambda,\mu)$ as well. ∎
Lemma 3.8 (Appending 0).
If Eq. 3.2 with n1n-1 variables holds for the pair $(\lambda,\mu)$, then Eq. 3.2 with $n$ variables holds for the pair $((\lambda,0),(\mu,0))$ as well.
See the next subsection for the proof.
Lemma 3.9 (Raising operator).
To prove that $\lambda\mathrel{\succcurlyeq}\mu$ implies Eq. 3.1, it suffices to assume that $\lambda$ and $\mu$ are adjacent in the majorization order. In that case, $\lambda=R_{ij}(\mu)$ for some $i<j$, where $R_{ij}(\mu)=(\dots,\mu_{i}+1,\dots,\mu_{j}-1,\dots)$ is the raising operator, see [18, pp. 8].
Proof.
The first claim follows by transitivity. The second claim is [18, I. (1.16)]. ∎
Note that if $\lambda$ and $\mu$ are majorization-adjacent, then they are of exactly two possible forms:
  • $\lambda=(\nu,k+l,l,\eta)$ and $\mu=(\nu,k+l-1,l+1,\eta)$, where $k\geqslant 2$, $l\geqslant 0$, and $\nu$ and $\eta$ (possibly empty) are partitions that make $\lambda$ and $\mu$ valid partitions;
  • $\lambda=(\nu,k+1,k^{m},k-1,\eta)$ and $\mu=(\nu,k^{m+2},\eta)$, where $k,m\geqslant 1$, and $\nu$ and $\eta$ (possibly empty) are partitions that make $\lambda$ and $\mu$ valid partitions.
By the reduction lemmas above, in order to prove the weaker statement $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq. 3.2, we may assume that
  • either $\lambda=(\nu,k,0)$ and $\mu=(\nu,k-1,1)$, where $\nu\in\mathcal{P}_{n-2}$ such that $\nu_{n-2}\geqslant k\geqslant 2$;
  • or $\lambda=(\nu,2,1^{m},0)$ and $\mu=(\nu,1^{m+2})$, where $1\leqslant m\leqslant n-2$, and $\nu\in\mathcal{P}_{n-m-2}$ such that $\nu_{n-m-2}\geqslant 2$.
The following toy example studies the simplest case of the first kind of these pairs $(\lambda,\mu)$.
Example 3.10 (Toy Example for CGS).
Let $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. Suppressing the argument $\tau$ from $P_{\lambda}(\cdot),P_{\mu}(\cdot)$ henceforth, we have:
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}x_{1}^{d-i}x_{2}^{i}$ (3.10)
and
$\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$ $=\frac{x_{1}x_{2}}{1}\frac{P_{\mu-\bm{1}}(x)}{P_{\mu-\bm{1}}(\bm{1})}=\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}x_{1}^{d-i-1}x_{2}^{i+1}.$ (3.11)
Setting $x=(1,s)$, $s>0$, we have
$f(s)$ $\coloneqq\frac{P_{\lambda}(1,s)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(1,s)}{P_{\mu}(\bm{1})}$
$=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}s^{i}-\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}s^{i+1}$
$=\sum_{i=0}^{d}\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right)s^{i}\eqqcolon\sum_{i}f_{i}s^{i}.$
Define for $0\leqslant i\leqslant d-2$,
$g_{i}\coloneqq\binom{d-2}{i}\frac{(\tau+d-1)}{\tau}\frac{(\tau)_{{i+1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}\in\mathbb{F}_{>0},$
and let $g_{-1}=g_{-2}=g_{d-1}=g_{d}=0$. Then one can show that
$f_{i}=g_{i-2}-2g_{i-1}+g_{i},\text{\quad or equivalently, \quad}\sum_{j=i+1}^{d}(j-i-1)f_{j}=g_{i}.$
This yields Eq. 3.1, because we have for any $s>0$:
$f(s)=(s-1)^{2}g(s)\coloneqq(s-1)^{2}\sum_{i=0}^{d-2}g_{i}s^{i}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}.$
Combining the results and the example above, Theorem 3.4 follows. ∎
3.3. Okounkov–Olshanski’s formula
As promised above, in this subsection we show Section 3.2. Set $\mathbb{R}^{n}_{\geqslant}\coloneqq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}\geqslant\cdots\geqslant x_{n}\,\}$. For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, write $y\prec x$ (interlacing) if
$x_{1}\geqslant y_{1}\geqslant x_{2}\geqslant y_{2}\geqslant\cdots\geqslant x_{n-1}\geqslant y_{n-1}\geqslant x_{n}.$ (3.12)
Let $\lambda\in\mathcal{P}_{n-1}$ be a partition of length at most n1n-1, and $(\lambda,0)\in\mathcal{P}_{n}$. Then [21, pp. 77] (where $\theta$ is our $\tau$) shows that for $x\in\mathbb{R}^{n}_{\geqslant}$,
$P_{(\lambda,0)}(x)=\frac{1}{C(\lambda,n;\tau)V(x)^{2\tau-1}}\int_{y\prec x}P_{\lambda}(y)V(y)\Pi(x,y)^{\tau-1}\mathop{}\!\mathrm{d}y,$ (3.13)
where
$C(\lambda,n;\tau)$ $\coloneqq\prod_{i=1}^{n-1}\mathop{\mathrm{B}}(\lambda_{i}+(n-i)\tau,\tau),$
$V(x)$ $\coloneqq\prod_{i<j}(x_{i}-x_{j}),$
$\Pi(x,y)$ $\coloneqq\prod_{i=1}^{n}\prod_{j=1}^{n-1}|x_{i}-y_{j}|=\prod_{i\leqslant j}(x_{i}-y_{j})\prod_{i>j}(y_{j}-x_{i}),$
and $\mathop{\mathrm{B}}(\alpha,\beta)$ is the Beta function.
We note that Okounkov–Olshanski’s formula has a better form when changed to the following normalization.
Proposition 3.11.
For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, define the integral kernel $K(x,y)$ by
$K(x,y)\coloneqq K_{n}(x,y;\tau)=\frac{V(y)\Pi(x,y)^{\tau-1}}{V(x)^{2\tau-1}}\frac{\Gamma(n\tau)}{\Gamma(\tau)^{n}}.$ (3.14)
Then
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}=\int_{y\prec x}\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}K(x,y)\mathop{}\!\mathrm{d}y,\qquad\forall\lambda\in\mathcal{P}_{n-1},\ x\in\mathbb{R}^{n}_{\geqslant}.$ (3.15)
In particular, $\int_{y\prec x}K(x,y)\mathop{}\!\mathrm{d}y=1$.
Proof.
By [18, VI. (10.20)], we have for any $\mu\in\mathcal{P}_{n}$,
$P_{\mu}(\bm{1}_{n})=\prod_{s\in\mu}\frac{a_{\mu}^{\prime}(s)+(n-l_{\mu}^{\prime}(s))\tau}{a_{\mu}(s)+(l_{\mu}(s)+1)\tau}.$
Now,
$\frac{P_{(\lambda,0)}(\bm{1}_{n})}{P_{\lambda}(\bm{1}_{n-1})}=\prod_{s\in\lambda}\frac{a_{\lambda}^{\prime}(s)+(n-l_{\lambda}^{\prime}(s))\tau}{a_{\lambda}^{\prime}(s)+(n-1-l_{\lambda}^{\prime}(s))\tau}=\prod_{i=1}^{n-1}\frac{((n-i+1)\tau)_{{\lambda_{i}}}}{((n-i)\tau)_{{\lambda_{i}}}}=\prod_{i=1}^{n-1}\frac{\frac{\Gamma(\lambda_{i}+(n-i+1)\tau)}{\Gamma((n-i+1)\tau)}}{\frac{\Gamma(\lambda_{i}+(n-i)\tau)}{\Gamma((n-i)\tau)}}$
and hence
$\frac{P_{(\lambda,0)}(\bm{1}_{n})}{P_{\lambda}(\bm{1}_{n-1})}\frac{1}{C(\lambda,n;\tau)}=\prod_{i=1}^{n-1}\frac{\Gamma(\tau)\Gamma((n-i)\tau)}{\Gamma((n-i+1)\tau)}=\frac{\Gamma(\tau)^{n}}{\Gamma(n\tau)}.$
We are now ready to prove Section 3.2.
Proof of Section 3.2.
Assume that Eq. 3.1 with n1n-1 variables $y=(y_{1},\dots,y_{n-1})$ holds for the pair $(\lambda,\mu)$. We first prove Eq. 3.2 for $x=(x_{1},\dots,x_{n})$ and the pair $((\lambda,0),(\mu,0))$. For $\tau>0$, $x\in\mathbb{R}_{\geqslant 0}^{n}$, since the kernel $K(x,y)>0$, we have
$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}-\frac{P_{(\mu,0)}(x)}{P_{(\mu,0)}(\bm{1}_{n})}=\int_{y\prec x}\left(\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}-\frac{P_{\mu}(y)}{P_{\mu}(\bm{1}_{n-1})}\right)K(x,y)\mathop{}\!\mathrm{d}y>0.$
Since this is symmetric in $x$, Eq. 3.2 holds for all $x\in[0,\infty]^{n}$. When $\tau=0$, Jack polynomials reduce to monomial symmetric polynomials and Eq. 3.2 is simply Muirhead’s inequality. ∎
3.4. Supporting evidence: Kadell’s integral
We end this section by providing supporting evidence for the assertion $(4)\implies(2)$ in Conjecture 1. We begin by recalling Kadell’s integral in [18, VI. Ex. 10.7] and [8] (with a different normalization). Namely, let $\lambda$ be a partition of length at most $n$, as usual, and let $r,s>0$ (or more generally, $r,s\in\mathbb{C}$ with $\Re r,\Re s>0$). Then
$I_{\lambda}=I_{\lambda}(\tau;r,s)$ $\coloneqq\int_{[0,1]^{n}}\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}|V(x)|^{2\tau}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}\mathop{}\!\mathrm{d}x$
$=\prod_{i=1}^{n}\frac{\Gamma((n-i)\tau+\lambda_{i}+r)\Gamma((i-1)\tau+s)\Gamma(i\tau+1)}{\Gamma((2n-i-1)\tau+\lambda_{i}+r+s)\Gamma(\tau+1)},$ (3.16)
where $x=(x_{1},\dots,x_{n})$, $\mathop{}\!\mathrm{d}x=\mathop{}\!\mathrm{d}x_{1}\cdots\mathop{}\!\mathrm{d}x_{n}$, and $V(x)=\prod_{i<j}(x_{i}-x_{j})$.
Let $\tilde{I}_{\lambda}\coloneqq I_{\lambda}/I_{(0^{n})}$ be the normalized Kadell’s integral. Then $\tilde{I}_{(0^{n})}=1$ and
$\tilde{I}_{\lambda}=\tilde{I}_{\lambda}(\tau;r,s)$ $=\prod_{i=1}^{n}\frac{\Gamma((n-i)\tau+\lambda_{i}+r)}{\Gamma((n-i)\tau+r)}\frac{\Gamma((2n-i-1)\tau+r+s)}{\Gamma((2n-i-1)\tau+\lambda_{i}+r+s)}$
$=\prod_{i=1}^{n}\frac{((n-i)\tau+r)_{{\lambda_{i}}}}{((2n-i-1)\tau+r+s)_{{\lambda_{i}}}}$ (3.17)
is a rational polynomial in $\tau$. We now show:
Proposition 3.12.
If $\lambda\mathrel{\succcurlyeq}\mu$, then $\tilde{I}_{\lambda}-\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$. If moreover $r$ and $s$ are rational numbers, then $\tilde{I}_{\lambda}-\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}$.
This supports $(4)\implies(2)$ in Conjecture 1, since for each fixed $\tau=\tau_{0}\in(0,\infty)$, if the difference $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}$ is positive over $0,1]^{n}$, then so is its integral against Kadell’s density
$\frac{1}{I_{(0^{n})}}|V(x)|^{2\tau}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}.$
Proof.
Note that $\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$. By Section 3.2, we may assume there exist $1\leqslant i<j\leqslant n$ such that $\lambda=R_{ij}(\mu)$, i.e., $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$ and $\lambda_{k}=\mu_{k}$ for $k\neq i,j$. Since $\lambda_{i}-\lambda_{j}-1\geqslant 0$,
$\tilde{I}_{\lambda}/\tilde{I}_{\mu}-1$ $=\frac{(n-i)\tau+\lambda_{i}+r-1}{(n-j)\tau+\lambda_{j}+r}\cdot\frac{(2n-j-1)\tau+\lambda_{j}+r+s}{(2n-i-1)\tau+\lambda_{i}+r+s}-1$
$=\frac{((j-i)\tau+\lambda_{i}-\lambda_{j}-1)((n-1)\tau+s)}{((n-j)\tau+\lambda_{j}+r)((2n-i-1)\tau+\lambda_{i}+r+s-1)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}.\qed$
4. Jack polynomial differences and the Muirhead semiring
In this section we prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq. 3.1) in some more cases. In fact, we prove a stronger version of the conjecture, in which positivity is replaced by sums of Muirhead differences (which are positive by Muirhead’s inequality).
4.1. Jack differences in the Muirhead semiring
To remind the reader: consider the outstanding theorem from the introduction (Theorem 1.4) and the ensuing discussion. We upgrade the final statement in that subsection to the following.
Conjecture 3.
Fix $\tau_{0}\in[0,\infty]$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$ (and conversely):
$\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (4.1)
Note that when $\tau_{0}=0$, Conjecture 3 is trivial as $P_{\lambda}(x;0)=m_{\lambda}(x)$. For $\tau_{0}=1$, $P_{\lambda}(x;1)=s_{\lambda}(x)$, and it is (informally) conjectured in [5, Section 7] that the difference in Eq. 4.1 is in the Muirhead cone $\mathcal{M}_{C}(\mathbb{Q}_{\geqslant 0})$. However, we now show that this is false.
Remark 4.1.
Conjecture 3 is false if we use the Muirhead cone instead of the semiring. For example, consider $n=3$, $\lambda=(5,2,0)$, and $\mu=(5,1,1)$. We have
$P(x)$ $\coloneqq 3(2\tau+1)(3\tau+1)(3\tau+2)(3\tau+4)\Bigg(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\Bigg)$
= $=\;2(\tau+1)(\tau+2)^{2}(2\tau+3)M_{(5,2,0)}-(\tau+1)(\tau+2)(2\tau^{2}+13\tau+12)M_{(5,1,1)}$
$\quad+6\tau(\tau+1)(\tau+2)(2\tau+3)M_{(4,3,0)}-2\tau(\tau+1)(2\tau^{2}+17\tau+17)M_{(4,2,1)}$
$\quad-12\tau(\tau+1)(\tau+2)M_{(3,3,1)}-\tau(\tau+1)(10\tau^{2}+\tau-20)M_{(3,2,2)}.$
If $10\tau^{2}+\tau-20<0$, that is, $0<\tau<\frac{3\sqrt{89}-1}{20}\approx 1.365$, the coefficient of $M_{(3,2,2)}$ is positive. Since $(3,2,2)$ is the minimum (in the majorization order) partition that appears in $P(x)$, $P(x)$ cannot lie in the Muirhead cone. In particular, this happens for the Schur case $\tau=1$. However, this remains consistent with Conjecture 3 because
$P(x)$ $=(2\tau^{4}+19\tau^{3}+43\tau^{2}+54\tau+24)(M_{(5,2,0)}-M_{(5,1,1)})$
$\mathrel{\phantom{=}}+\tau^{2}(2\tau^{2}+7\tau+7)(M_{(5,2,0)}-M_{(4,3,0)})$
$\mathrel{\phantom{=}}+\tau(14\tau^{3}+61\tau^{2}+75\tau+36)(M_{(4,3,0)}-M_{(4,2,1)})$
$\mathrel{\phantom{=}}+\tau^{3}(10\tau+23)(M_{(4,2,1)}-M_{(3,3,1)})$
$\mathrel{\phantom{=}}+\tau^{3}(10\tau+11)(M_{(3,3,1)}-M_{(3,2,2)})$
$\mathrel{\phantom{=}}+15\tau^{2}M_{(1,0,0)}(M_{(3,3,0)}-M_{(3,2,1)})$
$\mathrel{\phantom{=}}6\tau(6\tau+5)(M_{(2,0,0)}-M_{(1,1,0)})(M_{(3,2,0)}-M_{(3,1,1)})$
$\mathrel{\phantom{=}}+6\tau(M_{(2,0,0)}-M_{(1,1,0)})(M_{(3,1,1)}-M_{(2,2,1)}),$
which is in the Muirhead semiring $\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0})$ but not in the cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$.
Remark 4.2.
Here are three connections of Conjecture 3 to Muirhead’s inequality (Theorem 1.3). First, Conjecture 3, together with the inequality implies Conjecture 1. Second, the converse to this conjecture holds because of Muirhead’s inequality – see Theorem 1.3(2) – and the argument in the proof of Theorem 3.1. Thus Conjecture 3 provides yet another conjectural characterization of majorization. Third, the proof of Conjecture 3 for $\mu=(1^{|\lambda|})$ uses Muirhead’s inequality; see below.
Remark 4.3.
One of the conjectures above ($(4)\implies(1)$ in Conjecture 1) was formulated in all Lie types by McSwiggen and Novak in an interesting paper [19]; see their Conjecture 4.7 and the discussion on pp. 3997. The authors also showed the “easier half” in [19, Proposition 4.8] using asymptotics, as well as the other half for $\dim V=1$ in their Proposition 4.10. While we only work in type $A$, here are four ways in which we go beyond [19]: (a) We show the easier half (see Theorems 3.1 and 3.2) using alternate techniques, which prove the result not just for Jack polynomials but also for power sums (as was suspected in light of Theorem 1.2). These are not Jack specializations, so not covered in [19]. (b) Our method also characterizes weak majorization, again not addressed in [19]. (c) In fact we show Conjecture 3, which is stronger than Conjecture 1 (via the preceding remark) and which is not in [19]. Finally, (d) our proof of the “harder half” holds (for two variables) not just for normalized Jack differences, but also (Theorem 5.6) for Macdonald polynomials, which were “hoped for” in [19].
We now show
Theorem 4.4.
Conjectures 1, 2 and 3 hold for all $\lambda$ when $\mu=(1^{|\lambda|})$. Conjecture 3 moreover holds when $\lambda$ and $\mu$ are of length at most 2.
Proof of Theorem 4.4.
If $\lambda\mathrel{\succcurlyeq}\mu=(1^{|\lambda|})$, we prove more strongly that the difference lies in the Muirhead cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$. Note that $\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}$ is an $\mathbb{F}_{\geqslant 0}$-linear combination of $M_{\nu}$ (see the arguments for [18, VI. (10.15)] for the expansion of monic Jack into monomials; see also [12] for the expansion of integral Jack into augmented monomials), say $\sum_{\nu}u_{\lambda\nu}M_{\nu}$. Then $\sum_{\nu}u_{\lambda\nu}=1$ by our choice of normalization. Also, we have $\frac{P_{\mu}}{P_{\mu}(\bm{1})}=M_{\mu}$. Hence,
$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}}{P_{\mu}(\bm{1})}=\sum_{\nu}u_{\lambda\nu}(M_{\nu}-M_{\mu})$ (4.2)
lies in the Muirhead cone. This shows Conjecture 3, and hence Conjectures 1 and 2 for $\mu=(1^{|\lambda|})$.
We next come to the case of $\lambda,\mu\in\mathcal{P}_{2}$. Here we continue the calculations in Toy Example 3.2. Expanding the difference of the normalized Jack polynomials in Muirhead differences and suppressing the argument $\tau$ in them as usual, we obtain:
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\sum_{i=0}^{d}\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right)x_{1}^{d-i}x_{2}^{i}.$
Let
$a_{i}$ $\coloneqq 2\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right),\quad 0\leqslant i\leqslant\lfloor\frac{d-1}{2}\rfloor,$
$a_{d/2}$ $\coloneqq\binom{d}{d/2}\frac{(\tau)_{{d/2}}(\tau)_{{d/2}}}{(2\tau)_{{d}}}-\binom{d-2}{d/2-1}\frac{(\tau)_{{d/2-1}}(\tau)_{{d/2-1}}}{(2\tau)_{{d-2}}},\text{\quad if $d$ even}.$
Then by Abel’s lemma,
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\sum_{i=0}^{\lfloor\frac{d}{2}\rfloor}a_{i}M_{(d-i,i)}(x)=\sum_{i=0}^{\lfloor\frac{d}{2}\rfloor-1}\left(\sum_{j\leqslant i}a_{j}\right)(M_{(d-i,i)}(x)-M_{(d-i-1,i+1)}(x))$
where the summand $M_{(d/2,d/2)}$ is only present when $d$ is even. It is not hard to prove by telescoping that
$\sum_{j\leqslant i}a_{j}=2\binom{d-1}{i}\frac{d-1-2i}{d-1}\frac{(\tau)_{{i}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}(\tau+d-1)\in\mathbb{F}_{\geqslant 0},\quad i=0,\dots,\lfloor\frac{d}{2}\rfloor-1.$
Note that if $d$ is even, then $\sum_{j\leqslant d/2}a_{j}=0$. ∎
We end this part with a two-variable consequence (in fact reformulation) of the final part of Theorem 4.4.
Corollary 4.5.
Fix integer partitions $\lambda,\mu\in\mathcal{P}_{2}$ with $\lambda\succcurlyeq\mu$. For $x=(x_{1},x_{2})$, we have
$\frac{1}{V(x)^{2}}\left(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\right)\in\bigoplus_{\nu}\mathbb{F}_{\geqslant 0}\cdot s_{\nu}(x).$
Remark 4.6.
Notice that Section 4.1 does not hold for higher $n$ – not even for $n=3$, even at $\tau=1$, i.e. for Schur polynomials. (This is in parallel to Conjecture 3 also not holding for the Muirhead cone even with $n=3$ and $\tau=1$, as verified in Section 4.1.) For example, working with $(3)\succcurlyeq(1,1,1)$ yields
$s_{(3)}(x_{1},x_{2},x_{3})-s_{(3)}(1,1,1)s_{(1,1,1)}(x_{1},x_{2},x_{3})=\sum_{i}x_{i}^{3}+\sum_{i<j}(x_{i}^{2}x_{j}+x_{j}^{2}x_{i})-9x_{1}x_{2}x_{3},$
and this is not divisible by $x_{i}-x_{j}$ for any $i\neq j$. (To see why, set $x_{i}=x_{j}$ in the above expression, and note that the expression does not vanish.) Nor is it monomial-positive.
Proof of Section 4.1.
In fact for two variables, not only does the conjecture imply this assertion, but the converse is also true. This is because – we claim that – every normalized Muirhead-difference $M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})$ (for $\lambda\mathrel{\succcurlyeq}\mu$), when divided by $V(x)^{2}=(x_{1}-x_{2})^{2}$, is a Schur-positive polynomial with half-integer coefficients. This claim reduces by transitivity to majorization-adjacent partitions $\lambda=(a+b+1,a),\mu=(a+b,a+1)$ (for $a\geqslant 0,b\geqslant 1$), for which it holds by an explicit computation:
$\frac{2}{(x_{1}-x_{2})^{2}}\left(M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})\right)=s_{(a+b-1,a)}(x_{1},x_{2}).\qed$
4.2. Jack polynomial differences with varying parameter
Since $M_{\lambda}(x)=P_{\lambda}(x;0)/P_{\lambda}(\bm{1};0)$, Eq. 4.1 can be viewed as a positivity property involving Jack polynomial differences with different parameters. Let us generalize this idea further.
Define the Jack cone and Jack semiring over the cone $\mathbb{F}_{\geqslant 0}$ as follows: $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$ consists of $\mathbb{R}_{\geqslant 0}$-linear combinations of Jack differences $\{\,\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}-\frac{P_{\mu}(x;\sigma)}{P_{\mu}(\bm{1};\sigma)}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{J}_{S}^{\sigma}(\mathbb{R}_{\geqslant 0})$ consists of $\mathbb{R}_{\geqslant 0}$-linear combinations of products of functions in $\{\,\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}-\frac{P_{\mu}(x;\sigma)}{P_{\mu}(\bm{1};\sigma)}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$.
Conjecture 4.
Fix $0\leqslant\sigma\leqslant\tau\leqslant\infty$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference with parameter $\tau$ is positive in the Jack semiring $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$:
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0}).$ (4.3)
Note that Conjecture 4 is false when $\tau<\sigma$. For example, let $n=3$, $\tau=0$, $\sigma=1$, then $M_{(3)}-M_{(21)}=\frac{10}{3}S_{(3)}-4S_{(21)}+\frac{2}{3}S_{(111)}$, where $S_{\lambda}=s_{\lambda}/s_{\lambda}(\bm{1})$.
Also note that the expansion of $P_{\lambda}(x;\tau)$ into a sum of polynomials $P_{\mu}(x;\sigma)$ with $\sigma\leqslant\tau$ may not be positive. For example, when $n=3$, we have (suppressing the $x$)
$P_{(33)}(\tau)=P_{(33)}(\sigma)+\frac{6(\tau-\sigma)}{(2+\sigma)(2+\tau)}P_{(321)}(\sigma)+\frac{6(\tau-2\sigma)(\tau-\sigma)}{(1+\sigma)(1+2\sigma)(1+\tau)(2+\tau)}P_{(222)}(\sigma),$
and so the last coefficient is negative if $\sigma<\tau<2\sigma$.
When $\sigma=1$, these coefficients are falling factorials in $\tau$. There is a combinatorial conjecture in this case in [2]. (It should be noted that the parameter $\alpha$ in [2] is not the same as Macdonald’s $\alpha$, instead, it is the same as our $\tau$.)
Now, we again look at the case $n=2$. This time, the computation is more complicated.
Example 4.7 (Toy Example for Conjecture 4).
Let $n=2$ and $d\geqslant 2$. In this example, we mostly suppress the argument $x$, and also assume the indices $i,j=0,\dots,\lfloor d/2\rfloor$ unless otherwise stated.
We prove the stronger statement the difference in Eq. 4.3 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. By Sections 3.2 and 3.2, we have
$P_{(d-i,i)}(x;\tau)=\sum_{j}\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\cdot m_{(d-j,j)}(x),$
hence the transition matrix from $\left(P_{\lambda}(x;\tau)\right)$ to $\left(m_{\lambda}(x)\right)$ is
$M_{d}(P(\tau),m)\coloneqq\left(\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\right).$
For example, when $d=2,3$, we have
$M_{2}(P(\tau),m)=\begin{pmatrix}1&\frac{2\tau}{1+\tau}\\ &1\end{pmatrix},\quad M_{3}(P(\tau),m)=\begin{pmatrix}1&\frac{3\tau}{1+\tau}\\ &1\end{pmatrix}$,
that is, $P_{(2,0)}(\tau)=m_{(2,0)}+\frac{2\tau}{1+\tau}m_{(1,1)}$ and $P_{(3,0)}(\tau)=m_{(3,0)}+\frac{3\tau}{1+\tau}m_{(2,1)}$.
The transition matrix from $\left(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $(P_{\lambda}(x;\tau))$ is a diagonal matrix
$M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)\coloneqq\left(\frac{1}{P_{(d-i,i)}(\bm{1};\tau)}\delta_{ij}\right)=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-i}}}\delta_{ij}\right).$
We claim that its inverse is given by
$M_{d}\left(P(\sigma),m\right)^{-1}=\left((-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}\frac{(\sigma+i-j+1)_{{d-i-j}}(\sigma+d-i-j+1)_{{j-i}}}{(\sigma+1)_{{d-2i}}}\right).$
(When $i=j=d/2$, the factor $\frac{d-2i}{d-i-j}$ is understood to be 1.) For example, if $d=2,3$, then
$M_{2}(P(\tau),m)=\begin{pmatrix}1&-\frac{2\tau}{1+\tau}\\ &1\end{pmatrix},\quad M_{3}(P(\tau),m)=\begin{pmatrix}1&-\frac{3\tau}{1+\tau}\\ &1\end{pmatrix}$,
Note that $(-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}$ is related to the Cardan polynomial $2T_{n}(x/2)$ (where $T_{n}(x)$ is the Chebyshev polynomial), see [OEIS A034807].
We claim that the transition matrix $M_{d}\left(P(\tau),P(\sigma)\right)$ between Jack polynomials with different parameters, which is given by the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$, can be computed by
$M_{d}\left(P(\tau),P(\sigma)\right)$ $=M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$
$=\left(\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\right).$ (4.4)
(When $i>j$, the entries are understood to be 0.) Note that the first claim follows from the second, since upon setting $\tau=\sigma$, Section 4.2 yields the identity matrix.
To prove Section 4.2, it suffices to check the strictly upper diagonal entries. Let $i<j$, and compute the $(i,j)$-entry of the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$:
$\mathrel{\phantom{=}}\sum_{k=i}^{j}\binom{d-2i}{k-i}\frac{(\tau)_{{k-i}}(\tau)_{{d-k-i}}}{(\tau)_{{d-2i}}}(-1)^{k-j}\times$
$\mathrel{\phantom{=}}\qquad\binom{d-k-j}{j-k}\frac{d-2k}{d-k-j}\frac{(\sigma+k-j+1)_{{d-k-j}}(\sigma+d-k-j+1)_{{j-k}}}{(\sigma+1)_{{d-2k}}}$
$=\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{1}{(\tau)_{{d-2i}}}\cdot\sum_{k=i}^{j}(-1)^{k-j}\binom{j-i}{k-i}\frac{(d-k-j)!}{(d-k-i)!}\frac{d-2k}{d-k-j}\times$
$\mathrel{\phantom{=}}\qquad(\tau)_{{k-i}}(\tau)_{{d-k-i}}\frac{(\sigma+k-j+1)_{{d-k-j}}(\sigma+d-k-j+1)_{{j-k}}}{(\sigma+1)_{{d-2k}}}$
$=\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{1}{(\tau)_{{d-2i}}}(-1)^{i-j}\cdot\sum_{l=0}^{j-i}(-1)^{l}\binom{j-i}{l}\frac{(d-l-i-j)!}{(d-l-2i)!}\frac{d-2l-2i}{d-l-i-j}\times$
$\mathrel{\phantom{=}}\qquad(\tau)_{{l}}(\tau)_{{d-l-2i}}\frac{(\sigma+l+i-j+1)_{{d-l-i-j}}(\sigma+d-l-i-j+1)_{{j-i-l}}}{(\sigma+1)_{{d-2l-2i}}}$
where in the last identity, we set l=kil=k-i.
In other words, it suffices to show that
$\sum_{l=0}^{j-i}(-1)^{l}\binom{j-i}{l}\frac{(d-l-i-j)!}{(d-l-2i)!}\frac{d-2l-2i}{d-l-i-j}\times$
$\qquad(\tau)_{{l}}(\tau)_{{d-l-2i}}\frac{(\sigma+l+i-j+1)_{{d-l-i-j}}(\sigma+d-l-i-j+1)_{{j-i-l}}}{(\sigma+1)_{{d-2l-2i}}}$ (4.5)
$=$ $(-1)^{i-j}\frac{(\tau)_{{d-2i}}(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}.$
Denote by $f(i,j,d;l)$ the summand on the left, and $r(i,j,d)$ the RHS. We shall prove Section 4.2 using the so-called Wilf–Zeilberger algorithm [23, Chapter 7]. Let
$\mathrel{\phantom{=}}F(i,j,d;l)\coloneqq f(i,j,d;l)/r(i,j,d)$
$=(-1)^{j-i-l}\binom{j-i}{l}\frac{(d-l-i-j)!}{(d-l-2i)!}\frac{d-2l-2i}{d-l-i-j}\cdot\frac{1}{(\tau-\sigma)_{{j-i}}}\frac{(\tau)_{{l}}(\tau)_{{d-l-2i}}(\tau+d-i-j)_{{j-i}}}{(\tau)_{{d-2i}}}\times$
$\mathrel{\phantom{=}}\qquad\frac{(\sigma+l+i-j+1)_{{d-l-i-j}}(\sigma+d-l-i-j+1)_{{j-i-l}}(\sigma+d-2j+1)_{{j-i}}}{(\sigma+1)_{{d-2l-2i}}}$
$=(-1)^{j-i-l}\binom{j-i}{l}\frac{(d-l-i-j)!}{(d-l-2i)!}\frac{d-2l-2i}{d-l-i-j}\cdot\frac{1}{(\tau-\sigma)_{{j-i}}}\frac{(\tau)_{{l}}(\tau)_{{d-l-2i}}(\tau+d-i-j)_{{j-i}}}{(\tau)_{{d-2i}}}\times$
$\mathrel{\phantom{=}}\qquad\frac{(\sigma+l+i-j+1)_{{d-l-2i}}(\sigma+d-l-i-j+1)_{{j-i-l}}}{(\sigma+1)_{{d-2l-2i}}},$
and let
$G(i,j,d;l)\coloneqq F(i,j,d;l)\cdot\frac{l(\tau+d-l-2i)(\sigma-2j-1+d)(\sigma+l+i-j)}{(d-2l-2i)(-j+i+l-1)(\sigma-j+d-i)(-\tau+\sigma-j+i)}.$
It can be checked that
$F(i,j+1,d;l)-F(i,j,d;l)=G(i,j,d;l+1)-G(i,j,d;l).$
Hence,
$\sum_{l\geqslant 0}(F(i,j+1,d;l)-F(i,j,d;l))$ $=\sum_{l\geqslant 0}(G(i,j,d;l+1)-G(i,j,d;l))$
$=(G(i,j,d;j-i+1)-G(i,j,d;0))=0.$
In other words, we have
$\sum_{l\geqslant 0}F(i,j+1,d;l)=\sum_{l\geqslant 0}F(i,j,d;l).$
Note that, when $j=i+1$, we have
$\sum_{l\geqslant 0}F(i,i+1,d;l)=-\frac{(d-2i+\tau-1)\sigma}{(d-1-2i)(\tau-\sigma)}+\frac{\tau(d-1-2i+\sigma)}{(d-1-2i)(\tau-\sigma)}=1,$
and so for any $j>i$, $\sum_{l=0}^{j-i}F(i,j,d;l)=\sum_{l\geqslant 0}F(i,j,d;l)=1$. We have thus proved Section 4.2. The transition matrix from $\left(\frac{P_{\lambda}(\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $\left(\frac{P_{\lambda}(\sigma)}{P_{\lambda}(\bm{1};\sigma)}\right)$ is
$\mathrel{\phantom{=}}M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},\frac{P(\sigma)}{P(\bm{1};\sigma)}\right)$
$=M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)M_{d}\left(P(\tau),P(\sigma)\right)M_{d}\left(\frac{P(\sigma)}{P(\bm{1};\sigma)},P(\sigma)\right)^{-1}$
$=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-2i}}}\cdot\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\cdot\frac{(2\sigma)_{{d-2j}}}{(\sigma)_{{d-2j}}}\right).$
In particular, we are interested in the difference of the first two rows, that is, the coefficient of $\frac{P_{(d-j,j)}(\sigma)}{P_{(d-j,j)}(\bm{1};\sigma)}$ in $\frac{P_{(d,0)}(\tau)}{P_{(d,0)}(\bm{1};\tau)}-\frac{P_{(d-1,1)}(\tau)}{P_{(d-1,1)}(\bm{1};\tau)}$. For $j=0$, this coefficient is evidently positive. For $1\leqslant j\leqslant\lfloor d/2\rfloor$, this coefficient is
$\mathrel{\phantom{=}}\frac{(\tau)_{{d}}}{(2\tau)_{{d}}}\cdot\frac{d!}{(d-2j)!j!}\frac{(\tau-\sigma)_{{j}}}{(\tau+d-j)_{{j}}(\sigma+d-2j+1)_{{j}}}\cdot\frac{(2\sigma)_{{d-2j}}}{(\sigma)_{{d-2j}}}$
$\mathrel{\phantom{=}}-\frac{(\tau)_{{d-2}}}{(2\tau)_{{d-2}}}\cdot\frac{(d-2)!}{(d-2j)!(j-1)!}\frac{(\tau-\sigma)_{{j-1}}}{(\tau+d-1-j)_{{j-1}}(\sigma+d-2j+1)_{{j-1}}}\cdot\frac{(2\sigma)_{{d-2j}}}{(\sigma)_{{d-2j}}}$
$=\frac{(\tau)_{{d-2}}}{(2\tau)_{{d-2}}}\cdot\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-1}}}{(\tau+d-1-j)_{{j-1}}(\sigma+d-2j+1)_{{j-1}}}\cdot\frac{(2\sigma)_{{d-2j}}}{(\sigma)_{{d-2j}}}$
$\mathrel{\phantom{=}}\cdot\left(\frac{(\tau+d-2)(\tau+d-1)}{(2\tau+d-2)(2\tau+d-1)}\frac{d(d-1)}{j}\frac{\tau-\sigma+j-1}{\frac{(\tau+d-2)(\tau+d-1)}{\tau+d-j-1}(\sigma+d-j)}-1\right)$
$\propto\frac{d(d-1)}{j}\frac{\tau-\sigma+j-1}{(2\tau+d-2)(2\tau+d-1)}\frac{\tau+d-j-1}{\sigma+d-j}-1$
$\propto 2(2\tau-1)j^{2}-2(2\tau-1)(d+\sigma)j+d(d-1)(\tau-\sigma-1)\eqqcolon a_{j},$
where $\propto$ means that a positive factor is omitted. It suffices to show that $a_{j}<0$ for $j=j_{0},\dots,\lfloor d/2\rfloor$ for some $j_{0}\geqslant 1$.
Note that the last entry, $a_{\lfloor d/2\rfloor}$, is always negative: for $d=2k$ even, we have $a_{k}=-2k(\tau+k-1)(2\sigma+1)<0$; for $d=2k+1$ odd, we have $a_{k}=-2k(\tau+k)(2\sigma+1)<0$.
Also, we have $a_{j+1}-a_{j}=-2(2\tau-1)(\sigma+d-2j-1)$. If $\tau\geqslant\frac{1}{2}$, $(a_{j})$ is decreasing, in particular, $a_{j}$ satisfies the desired condition. If $\tau<\frac{1}{2}$, $(a_{j})$ is increasing, hence $a_{j}\leqslant a_{\lfloor d/2\rfloor}<0$.
This calculation leads to:
Theorem 4.8.
Conjecture 4 holds when $n=2$. It also holds for the pair $\lambda=(2,1^{n-2},0)\mathrel{\succcurlyeq}\mu=(1^{n})$, when $n\geqslant 3$.
Proof.
Again, we prove the stronger statement the difference in Eq. 4.3 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. First, let $n=2$. The example above shows that Conjecture 4 holds for the pair $\lambda=(d,0)\mathrel{\succcurlyeq}\mu=(d-1,1)$. It suffices to show that it holds for any adjacent pair $\lambda=(d-i,i)\mathrel{\succcurlyeq}\mu=(d-i-1,i+1)$, as the general case would follow by telescoping. But this holds because
$\frac{P_{(d-i,i)}(x;\tau)}{P_{(d-i,i)}(\bm{1};\tau)}-\frac{P_{(d-i-1,i+1)}(x;\tau)}{P_{(d-i-1,i+1;\tau)}(\bm{1};\tau)}$ $=(x_{1}x_{2})^{i}\cdot\left(\frac{P_{(d-2i,0)}(x;\tau)}{P_{(d-2i,0)}(\bm{1};\tau)}-\frac{P_{(d-2i-1,1)}(x;\tau)}{P_{(d-2i-1,1)}(\bm{1};\tau)}\right)$
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu}(x;\tau)}{P_{\nu}(\bm{1};\tau)}-\frac{P_{\xi}(x;\tau)}{P_{\xi}(\bm{1};\tau)}\right)$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu+(i,i)}(x;\tau)}{P_{\nu+(i,i)}(\bm{1};\tau)}-\frac{P_{\xi+(i,i)}(x;\tau)}{P_{\xi+(i,i)}(\bm{1};\tau)}\right).$
Now, let $n\geqslant 3$, $\lambda=(2,1^{n-2},0)$, and $\mu=(1^{n})$. We have
$P_{\lambda}(x;\tau)=m_{\lambda}(x)+\frac{n(n-1)\tau}{1+(n-1)\tau}m_{\mu}(x),\quad P_{\mu}(x;\tau)$ $=m_{\mu}(x).$
Note that $m_{\lambda}(\bm{1})=n(n-1)$ and $m_{\mu}(\bm{1})=1$, and so
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}=\frac{1+(n-1)\tau}{1+n\tau}M_{\lambda}(x)+\frac{\tau}{1+n\tau}M_{\mu}(x),\quad\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}=M_{\mu}(x).$
Thus,
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}=\frac{1+(n-1)\tau}{1+n\tau}(M_{\lambda}(x)-M_{\mu}(x)).$ (4.6)
Inverting Eq. 4.6, we have
$M_{\lambda}(x)-M_{\mu}(x)=\frac{1+n\sigma}{1+(n-1)\sigma}\left(\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}-\frac{P_{\mu}(x;\sigma)}{P_{\mu}(\bm{1};\sigma)}\right).$
Hence,
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}=\frac{1+(n-1)\tau}{1+n\tau}\frac{1+n\sigma}{1+(n-1)\sigma}\left(\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}-\frac{P_{\mu}(x;\sigma)}{P_{\mu}(\bm{1};\sigma)}\right).$
Hence in this case, $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}$ is a positive multiple of $\frac{P_{\lambda}(x;\sigma)}{P_{\lambda}(\bm{1};\sigma)}-\frac{P_{\mu}(x;\sigma)}{P_{\mu}(\bm{1};\sigma)}$ if $\tau,\sigma\in[0,\infty]$. ∎
4.3. A conjecture of Alexandersson–Haglund–Wang
For completeness, we end this section by recalling a related conjecture in [2]. Following the convention in [4, Section 6.2], we define the integral Jack polynomials as
$J_{\lambda}(x;\tau)=c_{\lambda}(\tau)P_{\lambda}(x;\tau),\quad\text{where}\quad c_{\lambda}(\tau)\coloneqq\prod_{s\in\lambda}(a_{\lambda}(s)+\tau(l_{\lambda}(s)+1)).$ (4.7)
For $s=(i,j)\in\lambda$, we have $a_{\lambda}(s)\coloneqq\lambda_{i}-j$ and $l_{\lambda}(s)=\lambda_{j}^{\prime}-i$ are the number of boxes direct to the East and South of $s$ (drawn in the English manner).
Note that our $J_{\lambda}(x;\tau)$ is the same as the $\tilde{J}_{\lambda}^{(\alpha)}(x)$ in [2], with $\alpha$ replaced by $\tau$ (note: not $\frac{1}{\tau}$); and is related to the $J_{\lambda}^{(\alpha)}(x)$ in [18, VI. (10.22)] by $J_{\lambda}(x;\tau)=\tau^{-|\lambda|}J^{(1/\tau)}_{\lambda}(x)$.
As shown in [12], the expansion coefficients of integral Jack into the (augmented) monomial basis are positive and integral and can be combinatorially interpreted as admissible (also known as non-attacking) tableaux.
It was recently conjectured in [2] that the expansion coefficients of the integral Jack $J_{\lambda}(x;\tau)$ into the Schur basis $s_{\mu}$ have certain positivity and integrality. Define
$v_{\lambda\mu}(\tau)\coloneqq\langle J_{\lambda}(\tau),s_{\mu}\rangle,\qquad\text{that is,}\qquad J_{\lambda}(\tau)=\sum_{\mu}v_{\lambda\mu}(\tau)s_{\mu}.$ (4.8)
Conjecture 5 ([2]).
Let $\lambda,\mu$ be partitions of $d$, and define the expansion coefficients $a_{k}(\lambda,\mu)$ and $b_{k}(\lambda,\mu)$ as follows:
$v_{\lambda\mu}(\tau)=\sum_{k=0}^{d-1}a_{k}(\lambda,\mu)\binom{\tau+k}{d}=\sum_{k=1}^{d}b_{d-k}(\lambda,\mu)\binom{\tau}{k}k!.$ (4.9)
Then $a_{k}(\lambda,\mu)$ and $b_{d-k}(\lambda,\mu)$ are in $\mathbb{Z}_{\geqslant 0}$. Furthermore, the polynomials $\sum_{k=0}^{d}a_{k}(\lambda,\mu)z^{k}$ and $\sum_{k=0}^{d}b_{d-k}(\lambda,\mu)z^{k}$ have only real zeros.
5. Macdonald polynomial differences: conjectures and results
Let $P_{\lambda}(x)=P_{\lambda}(x;q,t)$ denote the monic Macdonald polynomials. Our goal in this section is to explore if Conjecture 3, characterizing majorization in terms of (normalized) Jack differences belonging to the Muirhead semiring, can be strengthened to Macdonald polynomials – leading to a strengthening of Conjecture 1 as well. And indeed, we achieve both of these goals, in the cases where we have shown the conjectures above.
We begin by setting notation. Recall that the $q$-Pochhammer, $q$-integer, $q$-factorial, and $q$-binomial are defined respectively as follows:
$(a;q)_{{k}}\coloneqq\prod_{i=0}^{k-1}(1-aq^{i}),\quad[k]_{q}\coloneqq\sum_{i=0}^{k-1}q^{i}=\begin{dcases}\frac{q^{k}-1}{q-1},&q\neq 1;\\ k,&q=1,\end{dcases}\quad[k]_{q}!\coloneqq\prod_{i=1}^{k}[i]_{q},$ (5.1)
[…]
We now study Macdonald differences for $\lambda\succcurlyeq\mu$, starting with the case of two variables. The first question in extending the aforementioned conjectures is: how should one normalize the Macdonald polynomials. We begin by exploring the natural choice [18] of $x=t^{\delta}=(t^{n-1},\dots,t,1)$.
Example 5.1.
Let $n=2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. We have, using the combinatorial formula [18, VI. (7.13)],
$P_{\lambda}(x)=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}x_{1}^{i}x_{2}^{d-i}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}.$ (5.3)
By [18, VI. (6.11)], where $n=2$ and $t^{\delta}=(t,1)$, $P_{\lambda}(t^{\delta})=\frac{(t^{2};q)_{{d}}}{(t;q)_{{d}}}$. Hence,
$\frac{P_{\lambda}(x)}{P_{\lambda}(t^{\delta})}=\frac{(q;q)_{{d}}}{(t^{2};q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}.$
For $\mu=(d-1,1)$, we have $P_{\mu}(x)=x_{1}x_{2}P_{(d-2,0)}(x)$, hence
$\frac{P_{\mu}(x)}{P_{\mu}(t^{\delta})}$ $=\frac{x_{1}x_{2}}{t}\frac{(q;q)_{{d-2}}}{(t^{2};q)_{{d-2}}}\sum_{i=0}^{d-2}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-2-i}}}{(q;q)_{{d-2-i}}}x_{1}^{i}x_{2}^{d-2-i}$
$=\frac{(q;q)_{{d-2}}}{t(t^{2};q)_{{d-2}}}\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}.$
For example, let $d=2$. Then $\frac{P_{(1,1)}(x)}{P_{(1,1)}(t^{\delta})}=\frac{1}{t}x_{1}x_{2}$, and
$\frac{P_{(2,0)}(x)}{P_{(2,0)}(t^{\delta})}=\frac{1-qt}{(1+t)(1-qt^{2})}(x_{1}^{2}+x_{2}^{2})+\frac{(1+q)(1-t)}{(1+t)(1-qt^{2})}x_{1}x_{2},$
so
$\frac{P_{(2,0)}(x)}{P_{(2,0)}(t^{\delta})}-\frac{P_{(1,1)}(x)}{P_{(1,1)}(t^{\delta})}=\frac{(1-qt)}{t(t+1)(1-qt^{2})}(tx_{1}-x_{2})(x_{1}-tx_{2}).$
Assuming $q,t\in(0,1)$ and $x_{1},x_{2}\in(0,\infty)$, then the difference is negative if $t<\frac{x_{1}}{x_{2}}<\frac{1}{t}$ and non-negative otherwise.
Numerical experiments suggest that for $d\geqslant 3$, the difference $\frac{P_{(d,0)}(x)}{P_{(d,0)}(t^{\delta})}-\frac{P_{(d-1,1)}(x)}{P_{(d-1,1)}(t^{\delta})}$ behaves in the same way: assuming $q,t\in(0,1)$ and $x_{1},x_{2}\in(0,\infty)$, then the difference is negative if $t<\frac{x_{1}}{x_{2}}<\frac{1}{t}$ and non-negative otherwise.
Given the above example, we see that the natural choice of normalization, $x=t^{\delta}$, is not good for positivity. Instead, the previous choice for Jack polynomials, $x=\bm{1}_{n}$, will lead to a certain positivity, as demonstrated below.
Example 5.2.
Continue with $n=2$ and consider the difference $\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$. Now
$P_{\lambda}(\bm{1})=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}.$
Hence we have
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}\text{\quad and \quad}\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}.$
Their difference is
$\mathrel{\phantom{=}}\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$
$=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}-\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}$
$\propto\sum_{j=1}^{d-1}\frac{(t;q)_{{j-1}}}{(q;q)_{{j-1}}}\frac{(t;q)_{{d-1-j}}}{(q;q)_{{d-1-j}}}\cdot\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}$
$\mathrel{\phantom{=}}-\sum_{j=0}^{d}\frac{(t;q)_{{j}}}{(q;q)_{{j}}}\frac{(t;q)_{{d-j}}}{(q;q)_{{d-j}}}\cdot\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}$
$=\sum_{j=1}^{d-1}\frac{(t;q)_{{j-1}}}{(q;q)_{{j-1}}}\frac{(t;q)_{{d-1-j}}}{(q;q)_{{d-1-j}}}\cdot\frac{(t;q)_{{d}}}{(q;q)_{{d}}}(x_{1}^{d}+x_{2}^{d})+\sum_{i=1}^{d-1}\left(\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}\cdot\sum_{j=1}^{d-1}\frac{(t;q)_{{j-1}}}{(q;q)_{{j-1}}}\frac{(t;q)_{{d-1-j}}}{(q;q)_{{d-1-j}}}\right$.
$\mathrel{\phantom{=}}\left.-\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}\cdot\sum_{j=0}^{d}\frac{(t;q)_{{j}}}{(q;q)_{{j}}}\frac{(t;q)_{{d-j}}}{(q;q)_{{d-j}}}\right)x_{1}^{i}x_{2}^{d-i}.$
Define $a_{i}$ to be the coefficient such that (via Abel’s lemma)
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\sum_{i=0}^{\lfloor\frac{d}{2}\rfloor}a_{i}M_{(d-i,i)}(x)=\sum_{l=0}^{\lfloor\frac{d}{2}\rfloor-1}\left(\sum_{i=0}^{l}a_{i}\right)\left(M_{(d-l,l)}(x)-M_{(d-l-1,l+1)}(x)\right).$
Note that the sum $a_{0}+\cdots+a_{\lfloor\frac{d}{2}\rfloor}=0$, by evaluating at $x=(1,1)$.
For $k\geqslant 1$, let
$b_{k}\coloneqq\dfrac{1-tq^{k-1}}{1-q^{k}},\quad c_{k}\coloneqq\dfrac{(t;q)_{{k}}}{(q;q)_{{k}}}=\prod_{i=1}^{k}b_{i}.$
Then we have
$a_{0}$ $\coloneqq 2c_{d}\sum_{j=1}^{d-1}c_{j-1}c_{d-j-1},$
$a_{i}$ $\coloneqq 2c_{i}c_{d-i}\sum_{j=1}^{d-1}c_{j-1}c_{d-j-1}-2c_{i-1}c_{d-i-1}\sum_{j=0}^{d}c_{j}c_{d-j},\quad 1\leqslant i\leqslant\lfloor\frac{d-1}{2}\rfloor,$
$a_{d/2}$ $\coloneqq\left(c_{d/2}\right)^{2}\sum_{j=1}^{d-1}c_{j-1}c_{d-j-1}-\left(c_{d/2-1}\right)^{2}\sum_{j=0}^{d}c_{j}c_{d-j},\quad\text{if $d$ is even}.$
In order to “upgrade” Conjecture 3 from Jack to Macdonald polynomials, it suffices to show that $\sum_{i=0}^{l}a_{i}\geqslant 0$, for $1\leqslant l\leqslant\lfloor d/2\rfloor-1$.
Lemma 5.3.
Let $q,t>1$.
  1. (1)
    The sequence $(b_{i})_{i\geqslant 1}$ is positive and monotone.
  2. (2)
    For $1\leqslant l\leqslant\lfloor(d-1)/2\rfloor$, and $l<s_{i}<d-l$ for $i=1\dots,l$, the following $(2l+1)$-term polynomial is positive.
    $\mathrel{\phantom{=}}\sum_{i=1}^{l}\left(\prod_{m=1}^{i-1}b_{m}\cdot(b_{i}b_{d-i}-b_{s_{i}}b_{d-s_{i}})\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right)+\prod_{m=d-l}^{d}b_{m}$
    $=\prod_{m=1}^{l}b_{m}\cdot b_{d-l}+\sum_{i=1}^{l}\left(\prod_{m=1}^{i-1}b_{m}\cdot(b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}})\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right).$ (5.4)
Proof.
(1) It is clear that $b_{i}>0$. Note that
$b_{i+1}-b_{i}=\dfrac{1-tq^{i}}{1-q^{i+1}}-\dfrac{1-tq^{i-1}}{1-q^{i}}=\frac{q^{i-1}(q-t)(q-1)}{(1-q^{i})(1-q^{i+1})}.$
Hence $(b_{i})_{i}$ is increasing if $q>t$ and decreasing if $q<t$.
(2) We first show that the two expressions are equal. Separate the positive term for $i=l$, and collect the negative term for $i$ and the positive term for i1i-1; now the last term above becomes the positive term for $i=1$. If $q=t$, the $(2l+1)$-term polynomial becomes $1$. If $q>t$, then by part (1), $b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}}>0$. If $q<t$, then
$b_{i}b_{d-i}-b_{s_{i}}b_{d-s_{i}}$ $=\frac{1-tq^{i-1}}{1-q^{i}}\frac{1-tq^{d-i-1}}{1-q^{d-i}}-\frac{1-tq^{s_{i}-1}}{1-q^{s_{i}}}\frac{1-tq^{d-s_{i}-1}}{1-q^{d-s_{i}}}$
$=\frac{-q^{i-1}(t-q)(1-tq^{d-1})(1-q^{s_{i}-i})(1-q^{d-i-s_{i}})}{(1-q^{i})(1-q^{d-i-1})(1-q^{s_{i}})(1-q^{d-s_{i}-1})}>0.\qed$
Now, we prove $\sum_{i=0}^{l}a_{i}\geqslant 0$, for $1\leqslant l\leqslant\lfloor d/2\rfloor-1$.
$\bullet$ Let $d=2k$, $k\geqslant 1$. We have
$a_{0}$ $\coloneqq 4c_{d}\sum_{j=1}^{k-1}c_{j-1}c_{d-j-1}+2c_{k-1}^{2}c_{d},$
$a_{i}$ $\coloneqq 4c_{i}c_{d-i}\sum_{j=1}^{k-1}c_{j-1}c_{d-j-1}-4c_{i-1}c_{d-i-1}\sum_{j=1}^{k-1}c_{j}c_{d-j}-4c_{i-1}c_{d-i-1}c_{d}$
$\mathrel{\phantom{=}}+2c_{i}c_{k-1}^{2}c_{d-i}-2c_{i-1}c_{k}^{2}c_{d-i-1},\quad 1\leqslant i\leqslant k-1,$
$a_{k}$ $\coloneqq 2c_{k}^{2}\sum_{j=1}^{k-1}c_{j-1}c_{d-j-1}-2c_{k-1}^{2}\sum_{j=1}^{k-1}c_{j}c_{d-j}-2c_{k-1}^{2}c_{d}.$
For $1\leqslant l\leqslant k-1$, we have
$\mathrel{\phantom{=}}a_{0}+\cdots+a_{l}$
$=4c_{d}\sum_{j=1}^{k-1}c_{j-1}c_{d-j-1}+2c_{k-1}^{2}c_{d}$
$\mathrel{\phantom{=}}+\sum_{i=1}^{l}4c_{i}c_{d-i}\sum_{j=1}^{k-1}c_{j-1}c_{d-j-1}-\sum_{i=1}^{l}4c_{i-1}c_{d-i-1}\sum_{j=1}^{k-1}c_{j}c_{d-j}-\sum_{i=1}^{l}4c_{i-1}c_{d-i-1}c_{d}$
$\mathrel{\phantom{=}}+\sum_{i=1}^{l}2c_{i}c_{k-1}^{2}c_{d-i}-\sum_{i=1}^{l}2c_{i-1}c_{k}^{2}c_{d-i-1}$
$=4c_{d}\sum_{j=l+1}^{k-1}c_{j-1}c_{d-j-1}+4\sum_{i=1}^{l}c_{i}c_{d-i}\sum_{j=l+1}^{k-1}c_{j-1}c_{d-j-1}-4\sum_{i=1}^{l}c_{i-1}c_{d-i-1}\sum_{j=l+1}^{k-1}c_{j}c_{d-j}$
$\mathrel{\phantom{=}}+2c_{k-1}^{2}c_{d}+2\sum_{i=1}^{l}c_{i}c_{k-1}^{2}c_{d-i}-2\sum_{i=1}^{l}c_{i-1}c_{k}^{2}c_{d-i-1},\quad 1\leqslant i\leqslant k-1,$
$=4\sum_{j=l+1}^{k-1}\left(c_{j-1}c_{d-j-1}c_{d}+\sum_{i=1}^{l}\left(c_{i}c_{d-i}c_{j-1}c_{d-j-1}-c_{i-1}c_{d-i-1}c_{j}c_{d-j}\right)\right)$
$\mathrel{\phantom{=}}+2\left(c_{k-1}^{2}c_{d}+\sum_{i=1}^{l}\left(c_{i}c_{k-1}^{2}c_{d-i}-c_{i-1}c_{k}^{2}c_{d-i-1}\right)\right)$
$=4\sum_{j=l+1}^{k-1}c_{j-1}c_{d-j-1}c_{d-l-1}\left(\prod_{m=d-l}^{d}b_{m}+\sum_{i=1}^{l}\prod_{m=1}^{i-1}b_{m}\cdot\left(b_{i}b_{d-i}-b_{j}b_{d-j}\right)\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right)$
$\mathrel{\phantom{=}}+2c_{k-1}^{2}c_{d-l-1}\left(\prod_{m=d-l}^{d}b_{m}+\sum_{i=1}^{l}\prod_{m=1}^{i-1}b_{m}\cdot\left(b_{i}b_{d-i}-b_{k}^{2}\right)\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right).$
Each expression enclosed in parentheses is of the form in Eq. 5.4, hence is positive.
$\bullet$ Let $d=2k+1$, $k\geqslant 1$. We have
$a_{0}$ $\coloneqq 4c_{d}\sum_{j=1}^{k}c_{j-1}c_{d-j-1},$
$a_{i}$ $\coloneqq 4c_{i}c_{d-i}\sum_{j=1}^{k}c_{j-1}c_{d-j-1}-4c_{i-1}c_{d-i-1}\sum_{j=1}^{k}c_{j}c_{d-j}-4c_{i-1}c_{d-i-1}c_{d},\quad 1\leqslant i\leqslant k.$
For $1\leqslant l\leqslant k$, we have
$\mathrel{\phantom{=}}\frac{1}{4}(a_{0}+\cdots+a_{l})$
$=\sum_{j=l+1}^{k}c_{j-1}c_{d-j-1}c_{d}+\sum_{i=1}^{l}c_{i}c_{d-i}\sum_{j=l+1}^{k}c_{j-1}c_{d-j-1}-\sum_{i=1}^{l}c_{i-1}c_{d-i-1}\sum_{j=l+1}^{k}c_{j}c_{d-j}$
$=\sum_{j=l+1}^{k}\left(c_{j-1}c_{d-j-1}c_{d}+\sum_{i=1}^{l}\left(c_{i}c_{d-i}c_{j-1}c_{d-j-1}-c_{i-1}c_{d-i-1}c_{j}c_{d-j}\right)\right)$
$=\sum_{j=l+1}^{k}c_{j-1}c_{d-j-1}c_{d-l-1}\left(\prod_{m=d-l}^{d}b_{m}+\sum_{i=1}^{l}\prod_{m=1}^{i-1}b_{m}\cdot\left(b_{i}b_{d-i}-b_{j}b_{d-j}\right)\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right).$
As the expression within each parentheses is of the form in Eq. 5.4, it is positive. ∎
Thus, we arrive at the following conjecture for normalized Macdonald differences in the Muirhead semiring – and we now merge this with the Macdonald analogue of the Cuttler–Greene–Skandera conjecture too. Recall here that $\mathbb{F}_{\geqslant 0}$ and $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ for the Macdonald case are defined in Eq. 2.12.
Conjecture 6 (CGS and Muirhead conjectures for Macdonald polynomials).
Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
  1. (1)
    The Macdonald difference lies in the Muirhead semiring over $\mathbb{F}_{\geqslant 0}$:
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0}).$ (5.5)
  2. (2)
    We have
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.6)
  3. (3)
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
    $\frac{P_{\lambda}(x;q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x;q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (5.7)
  4. (4)
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
    $\frac{P_{\lambda}(x;q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x;q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in(0,1)^{n}\cup(1,\infty)^{n}.$ (5.8)
  5. (5)
    $\lambda$ majorizes $\mu$.
Remark 5.4.
By the well-known duality $P_{\lambda}(x;q,t)=P_{\lambda}(x;1/q,1/t)$, one can define instead $\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }0<q,t<1\,\}$ and get an equivalent statement.
Remark 5.5.
For completeness, we record two additional equivalent conditions in Conjecture 6:
  • $(1^{\prime})$
    For all real $q,t>1$, the Macdonald difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (5.9)
  • $(2^{\prime})$
    For all real $q,t>1$, we have
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in[0,\infty),\quad\forall x\in[0,\infty)^{n}.$ (5.10)
These come from specializing assertions (1) and (2) respectively, to scalars $q,t>1$. It is also clear that $(1)\implies(1^{\prime})\implies(2^{\prime})$ and $(1)\implies(2)\implies(2^{\prime})$; and $(2^{\prime})\implies(3)$. Moreover, the counterpart of $(2^{\prime})$ can also be added in Conjectures 1 and 2 (and in Conjecture 7); as also the counterpart of $(1^{\prime})$ in Conjecture 3.
As promised, we now prove Conjecture 6 in the cases where we have shown it for Jack polynomials above.
Theorem 5.6.
Each assertion in Conjecture 6 implies the next. Moreover, the implication $(5)\implies(1)$ – and hence Conjecture 6 – moreover holds for all $\lambda$, with $\mu=(1^{|\lambda|})$. It also holds for $n=2$ and $\lambda,\mu$ arbitrary.
Proof.
That $(1)\implies(2)$ follows from Muirhead’s inequality, and $(2)\implies(3)\implies(4)$ are obvious. The proof of $(4)\implies(5)$ is via the argument used to show Theorem 3.2: if (4) holds for all $x\in(1,\infty)^{n}$ then $\lambda$ weakly majorizes $\mu$; if it holds for $x\in(0,1)^{n}$ then $\lambda\succcurlyeq_{w}-\mu$. Together, these are equivalent to $\lambda\succcurlyeq\mu$.
Next, we assume $\mu={\bf 1}_{|\lambda|}$, with $n,\lambda$ arbitrary and show $(5)\implies(1)$. Haglund–Haiman–Loehr have shown [7] that the integral Macdonald polynomials are monomial-positive over $\mathbb{F}_{\geqslant 0}$ for $q,t\in(0,1)$, whence so are the $P_{\lambda}$. Thus by the duality in Section 5, $P_{\lambda}(x)$ is monomial-positive for $q_{0},t_{0}\in(1,\infty)$. Now the calculation in Eq. 4.2 implies (1).
Finally, let $n=2$ and $\lambda,\mu$ be arbitrary. To show $(5)\implies(1)$, it suffices via telescoping to do so for majorization-adjacent partitions $\lambda=(d-i,i)$ and $\mu=(d-i-1,i+1)$. We showed this in Section 5 for $i=0$; and we use this to compute for $i>0$:
$\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}$ $=(x_{1}x_{2})^{i}\left(\frac{P_{(d-2i,0)}(x;q,t)}{P_{(d-2i,0)}(\bm{1};q,t)}-\frac{P_{(d-2i-1,1)}(x;q,t)}{P_{(d-2i-1,1)}(\bm{1};q,t)}\right)$
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu}(x)-M_{\xi}(x))$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu+(i,i)}(x)-M_{\xi+(i,i)}(x)).\qed$
Remark 5.7.
For two variables, the above proof reveals that Conjecture 6(5) implies the stronger (than (1) a priori, and hence equivalent) statement that the normalized Macdonald difference for $\lambda,\mu$ lies in the Muirhead cone, just like for Jack polynomials. As Section 4.1 shows, this is false in general for three variables.
We conclude with two other conjectures, which “upgrade” weak majorization and containment from Jack to Macdonald polynomials.
Conjecture 7 (KT Conjecture for Macdonald polynomials).
The following are equivalent:
  1. (1)
    We have
    $\frac{P_{\lambda}(x+\bm{1};q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x+\bm{1};q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.11)
  2. (2)
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
    $\frac{P_{\lambda}(x+\bm{1};q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x+\bm{1};q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (5.12)
  3. (3)
    $\lambda$ weakly majorizes $\mu$.
As above, we now provide some positive evidence for this conjecture.
Theorem 5.8.
Each assertion in Conjecture 7 implies the next. Moreover, Conjecture 7 holds for all $n,\lambda$ when $\mu=(1^{m})$ for some $m\geqslant 1$.
Proof.
That $(1)\implies(2)$ is immediate, and that $(2)\implies(3)$ again follows via the argument used to show Theorem 3.2. Next, let $\lambda\succcurlyeq_{w}\mu$, and let $\mu^{\prime}={\bf 1}_{|\lambda|}$; then $\lambda\succcurlyeq\mu^{\prime}\supseteq\mu$. Now directly compute (suppressing $q_{0},t_{0}$ from the arguments):
$\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}=\left(\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}\right)+\left(\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}\right).$
At each $x\in[0,\infty)^{n}$, the first difference on the right is in $\mathbb{F}_{\geqslant 0}$ by Theorem 5.6, while the second is the scalar $M_{\mu^{\prime}}(x+{\bf 1})-M_{\mu}(x+{\bf 1})$, where both terms are positive reals. Now compute their ratio:
$\frac{M_{\mu^{\prime}}(x+{\bf 1})}{M_{\mu}(x+{\bf 1})}=\prod_{i=1}^{n}(x_{i}+1)^{\mu^{\prime}_{i}-\mu_{i}},$
and this is at least $1$ since all $x_{i}\in[0,\infty)$ and $\mu^{\prime}\supseteq\mu$. ∎
Conjecture 7 is not contingent upon our next and final conjecture, but the proof of its counterpart for Jack polynomials used a characterization of containment via Jack polynomials. Thus, we end with:
Conjecture 8 (Containment via Macdonald differences).
The Macdonald-counterpart over $\mathbb{F}_{\geqslant 0}$ of Theorem 1.1 holds.
Acknowledgments
S.S. and A.K. would like to thank IISc Bangalore and IAS Princeton, respectively, for their hospitality during summer 2022 and summer 2023, when this work was initiated and then continued. H.C. was partially supported by the Lebowitz Summer Research Fellowship and the SAS Fellowship at Rutgers University. A.K. was partially supported by SwarnaJayanti Fellowship grants SB/SJF/2019-20/14 and DST/SJF/MS/2019/3 from SERB and DST (Govt. of India) and by a Shanti Swarup Bhatnagar Award from CSIR (Govt. of India). The research of S.S. during this project has been supported by NSF grant DMS-2001537 and Simons Foundation grants 509766 and 00006698.

### v2

@@
 Here we extend the inequalities by these authors to Jack and Macdonald polynomials, and obtain conjectural characterizations of majorization and of weak majorization of the underlying partitions.
-We prove these characterizations for several cases of partitions, including all partitions with two parts.
-In fact, we upgrade – and prove in the above cases – the characterization of majorization, to containment of Jack and Macdonald differences lying in the Muirhead semiring.
+We prove these characterizations for two variables.
+In fact, we upgrade—and prove in the above cases—the characterization of majorization, to containment of Jack and Macdonald differences lying in the Muirhead semiring.
 Key words and phrases:
@@
 By the work of Green and Macdonald, these functions are related to the representation theory of $GL_{n}$ over finite and $p$-adic fields.
-In the late 1960s, Jack discovered Jack polynomials, unifying Schur polynomials and zonal polynomials – the latter being related to the representation theory of $GL_{n}(\mathbb{R})$ and multivariate statistics.
+In the late 1960s, Jack discovered Jack polynomials, unifying Schur polynomials and zonal polynomials—the latter being related to the representation theory of $GL_{n}(\mathbb{R})$ and multivariate statistics.
 Hall–Littlewood polynomials and Jack polynomials are quite different generalizations of Schur polynomials.
@@
 1.2. Symmetric function inequalities
-The notion of positivity has an even longer history: it begin with mathematics itself – counting and measuring objects. Let us restrict our attention to inequalities about symmetric functions.
+The notion of positivity has an even longer history: it begin with mathematics itself—counting and measuring objects. Let us restrict our attention to inequalities about symmetric functions.
 Most families of symmetric polynomials are indexed by integer partitions.
-An integer partition (of length at most $n$) is an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of integers such that $\lambda_{1}\geqslant\cdots\geqslant\lambda_{n}\geqslant 0$.
-Denote the set of these partitions by $\mathcal{P}_{n}$.
+A partition (of length at most $n$) is an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of integers such that $\lambda_{1}\geqslant\cdots\geqslant\lambda_{n}\geqslant 0$.
+Denote the set of such partitions by $\mathcal{P}_{n}$.
+Let $|\lambda|\coloneqq\lambda_{1}+\cdots+\lambda_{n}$ be the size of $\lambda$.
 Recall that for partitions $\lambda$ and $\mu$ in $\mathcal{P}_{n}$, one has the following partial orders:
@@
 Of great importance in representation theory and algebraic combinatorics is the notion of expansion positivity.
-We say a symmetric polynomial is Schur positive if it can be written as a sum of Schur polynomials with non-negative (integer) coefficients.
+We say a symmetric polynomial is Schur positive if it can be written as a sum of Schur polynomials with non-negative coefficients.
 For example, by the celebrated Littlewood–Richardson rule, the product of two Schur polynomials is Schur positive: $s_{\lambda}\cdot s_{\mu}=\sum_{\nu}c_{\lambda\mu}^{\nu}s_{\nu}$, where the Littlewood–Richardson coefficients $c_{\lambda\mu}^{\nu}\geqslant 0$.
@@
 Expansion positivity naturally suggests underlying combinatorial objects or structure.
-For example, Littlewood–Richardson coefficients have various combinatorial interpretations: Littlewood–Richardson tableaux [27, 18], Berenstein–Zelevinsky patterns [3] and Knutson–Tao honeycombs [13]; and Kostka numbers count the number of semi-standard Young tableaux (SSYT) [27, 18].
-For Jack polynomials $P_{\lambda}^{(\alpha)}(x)$, it was conjectured in Macdonald’s book [18, VI.(10.26?)] and proved by Knop and Sahi [12] that their expansions in terms of the monomial basis $(m_{\lambda})$ are positive (and integral, under certain normalizations) in the parameter $\alpha$.
+For example, Littlewood–Richardson coefficients have various combinatorial interpretations: Littlewood–Richardson tableaux [26, 17], Berenstein–Zelevinsky patterns [2] and Knutson–Tao honeycombs [12]; and Kostka numbers count the number of semi-standard Young tableaux [26, 17].
+For Jack polynomials $P_{\lambda}^{(\alpha)}(x)$, it was conjectured in Macdonald’s book [17, VI.(10.26?)] and proved by Knop and Sahi [11] that their expansions in terms of the monomial basis $(m_{\lambda})$ are positive (and integral, under certain normalizations) in the parameter $\alpha$.
 The following binomial formula for Jack polynomials is due to Okounkov and Olshanski [20]:
@@
 Recently, Chen and Sahi proved [3] that the coefficient $\binom{\lambda}{\nu}_{\alpha}$ is monotone in $\lambda$:
-$\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\in\mathbb{F}_{\geqslant 0},\quad\lambda\supseteq\mu$ (1.2)
+$\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\in\mathbb{F}_{\geqslant 0},\quad\lambda\supseteq\mu,$ (1.2)
 which, in turn, leads to the Jack positivity of $\frac{P_{\lambda}^{(\alpha)}(x+\bm{1})}{P_{\lambda}^{(\alpha)}(\bm{1})}-\frac{P_{\mu}^{(\alpha)}(x+\bm{1})}{P_{\mu}^{(\alpha)}(\bm{1})}$ for $\lambda\supseteq\mu$:
@@
 The study of generalized binomial coefficients originated in two distinct contexts.
-In statistics, their zonal case ($\alpha=2$) was first examined by Constantine in the 1960s, while Lascoux [15] explored their Schur case ($\alpha=1$) to compute Chern classes for exterior and symmetric squares of vector bundles (see also [18, p.
+In statistics, their zonal case ($\alpha=2$) was first examined by Constantine in the 1960s, while Lascoux [14] explored their Schur case ($\alpha=1$) to compute Chern classes for exterior and symmetric squares of vector bundles (see also [17, p.
 47, Example 10]).
@@
 In what follows, we write Jack polynomials with parameter $\tau=1/\alpha$, that is, our $P_{\lambda}(x;\tau)$ is equal to Macdonald’s $P_{\lambda}^{(1/\tau)}(x)$.
-Theorem 1.1 ([4, Theorem 6.1]).
-The following statements are equivalent:
+Theorem 1.1 ([3, Theorem 6.1]).
+The following are equivalent for partitions $\lambda,\mu$:
+  1. (1)
+    The difference $\frac{P_{\lambda}(x+\bm{1};\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x+\bm{1};\tau)}{P_{\mu}(\bm{1};\tau)}$ is Jack positive over $\mathbb{F}_{\geqslant 0}$, namely, can be written as an $\mathbb{F}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau)$.
+  2. (2)
+    For every $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
+  3. (3)
+    For some $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
   4. (4)
     $\lambda$ contains $\mu$, $\lambda\supseteq\mu$.
-  2. (2)
-    For any fixed $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
-  3. (3)
-    The difference $\frac{P_{\lambda}(x+\bm{1};\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x+\bm{1};\tau)}{P_{\mu}(\bm{1};\tau)}$ is Jack positive over $\mathbb{F}_{\geqslant 0}$, namely, can be written as an $\mathbb{F}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau)$.
-It is well-known, see [18, Chapters I, VI, VII], that Jack polynomials $P_{\lambda}(x;\tau)$ specialize to many other symmetric polynomials: monomial symmetric polynomials $m_{\lambda}$ when $\tau=0$, zonal polynomials $Z_{\lambda}$ when $\tau=1/2$ (spherical functions for $GL_{n}(\mathbb{R})/O(n)$) and $\tau=2$ (for $GL_{n}(\mathbb{H})/Sp(n)$), Schur polynomials $s_{\lambda}$ when $\tau=1$, and elementary symmetric polynomials $e_{\lambda^{\prime}}$ when $\tau=\infty$ (where $\lambda^{\prime}$ is the conjugate of $\lambda$).
-Hence, we also have the following characterization, which somewhat surprisingly also holds for power-sums:
-Theorem 1.2 ([4, Theorems 6.2, 6.4]).
-The following statements are equivalent:
+It is well-known, see [17, Chapters I, VI, VII], that Jack polynomials $P_{\lambda}(x;\tau)$ specialize to many other symmetric polynomials: monomial symmetric polynomials $m_{\lambda}$ when $\tau=0$, zonal polynomials $Z_{\lambda}$ when $\tau=1/2$ (spherical functions for $GL_{n}(\mathbb{R})/O_{n}$) and $\tau=2$ (for $GL_{n}(\mathbb{H})/Sp_{n}$), Schur polynomials $s_{\lambda}$ when $\tau=1$, and elementary symmetric polynomials $e_{\lambda^{\prime}}$ when $\tau=\infty$ (where $\lambda^{\prime}$ is the conjugate of $\lambda$).
+Hence, we also have the following characterization, which somewhat surprisingly also holds for power sums:
+Theorem 1.2 ([3, Theorems 6.2, 6.3]).
+The following are equivalent for partitions $\lambda,\mu$:
   1. (1)
@@
 For partitions $\lambda$ and $\mu$ with $|\lambda|=|\mu|$, Muirhead showed in [19] that for the monomial symmetric polynomial $m_{\lambda}$
-$\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}\geqslant\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})},\quad x\in(0,\infty)^{n}$ (1.6)
+$\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}\geqslant\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})},\quad x\in[0,\infty)^{n}$ (1.6)
 if and only if $\lambda\mathrel{\succcurlyeq}\mu$.
 (Note that (1.6) holds even when $\lambda\in\mathbb{R}^{n}$, in which case, $m_{\lambda}$ is defined as the orbit-sum of $x^{\eta}=x_{1}^{\eta_{1}}\cdots x_{n}^{\eta_{n}}$, where $\eta$ runs over the orbit $S_{n}\cdot\lambda$, acting by permutation.)
-Inspired by the inequalities above (and some others), Cuttler, Greene, and Skandera [5] generalized the inequalities of Newton and Gantmacher to elementary symmetric polynomials and power sums indexed by partitions, and conjectured an inequality for Schur polynomials, which was later proved by Sra [26] (again, for $\lambda,\mu$ real $n$-tuples).
+Cuttler, Greene, and Skandera [4] generalized the inequalities of Newton and Gantmacher to elementary symmetric polynomials and power sums indexed by partitions, and conjectured an inequality for Schur polynomials, which was later proved by Sra [25] (again, for $\lambda,\mu$ real $n$-tuples).
 We recall all these results.
@@
   3. (3)
-    ([5, Theorem 3.2], generalized Newton’s inequality) The following difference is positive:
+    ([4, Theorem 3.2], generalized Newton’s inequality) The following difference is positive:
     $\frac{e_{\lambda^{\prime}}(x)}{e_{\lambda^{\prime}}(\bm{1})}-\frac{e_{\mu^{\prime}}(x)}{e_{\mu^{\prime}}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.8)
   4. (4)
-    ([5, Theorem 4.2], generalized Gantmacher’s inequality) The following difference is positive:
+    ([4, Theorem 4.2], generalized Gantmacher’s inequality) The following difference is positive:
     $\frac{p_{\lambda}(x)}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}(x)}{p_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.9)
   5. (5)
-    ([5, Conjecture 7.4, Theorem 7.5] and [26], Cuttler–Greene–Skandera and Sra’s inequality) The following difference is positive:
+    ([4, Conjecture 7.4, Theorem 7.5] and [25], Cuttler–Greene–Skandera and Sra’s inequality) The following difference is positive:
     $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.10)
@@
 In this paper, we extend these conjectures and prove them in some cases.
-As we will see below – as also in [5, 10, 19, 26] – the more challenging implication is to show that if $\lambda\mathrel{\succcurlyeq}\mu$ then the normalized differences of various symmetric polynomials are non-negative on $0,\infty)^{n}$.
-An alternate, stronger approach would be to show that these differences “decompose” non-negatively into non-negative quantities, like monomials and differences of monomials as in part (2).
+As we will see below—as also in [4, 9, 18, 25]—the more challenging implication is to show that if $\lambda\mathrel{\succcurlyeq}\mu$ then the normalized differences of various symmetric polynomials are non-negative on $0,\infty)^{n}$.
+An alternate, stronger approach would be to show that these differences “decompose” non-negatively into non-negative quantities, like monomials and differences of monomials as in Muirhead’s inequality.
 This is known for $e_{\lambda^{\prime}}$ and for $p_{\lambda}$, as we now state.
 Let $M_{\lambda}\coloneqq m_{\lambda}/m_{\lambda}(\bm{1})$ be the normalized monomial.
-Recall [5, Section 6] that the Muirhead cone $\mathcal{M}_{C}(\mathcal{C})$ and the Muirhead semiring $\mathcal{M}_{S}(\mathcal{C})$ over a cone $\mathcal{C}$ are defined as follows: $\mathcal{M}_{C}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of Muirhead differences $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{M}_{S}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of products of functions in $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,M_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$.
+Recall [4, Section 6] that the Muirhead cone $\mathcal{M}_{C}(\mathcal{C})$ and the Muirhead semiring $\mathcal{M}_{S}(\mathcal{C})$ over a cone $\mathcal{C}$ are defined as follows: $\mathcal{M}_{C}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of Muirhead differences $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{M}_{S}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of products of functions in $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,M_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$.
 In particular, when evaluated at $x\in[0,\infty)^{n}$, functions in $\mathcal{M}_{C}(\mathcal{C})$ and $\mathcal{M}_{S}(\mathcal{C})$ take values in the cone $\mathcal{C}$.
-Theorem 1.4 ([5, Theorem 6.1]).
+Theorem 1.4 ([4, Theorem 6.1, Corollary 6.2]).
 Given partitions $\lambda\mathrel{\succcurlyeq}\mu$, the following differences
@@
 lie in the Muirhead semiring $\mathcal{M}_{S}$ over $\mathbb{Q}_{\geqslant 0}$.
-It should be noted that the Muirhead semiring $\mathcal{M}_{S}$ is strictly larger than the Muirhead cone $\mathcal{M}_{C}$.
-Some power sum differences belong to $\mathcal{M}_{S}\setminus\mathcal{M}_{C}$.
-For example, when $n=3$, $\lambda=(33)$ and $\mu=(321)$, we have
-$\frac{p_{\lambda}}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}}{p_{\mu}(\bm{1})}=\frac{2}{3}M_{(3)}\left(M_{(3)}-M_{(21)}\right)=\frac{2}{9}\left(M_{(6)}-M_{(51)}-M_{(42)}+2M_{(33)}-M_{(321)}\right),$
-which is in $\mathcal{M}_{S}$ but not in $\mathcal{M}_{C}$. However, it seems that the differences of elementary symmetric polynomials lie in the Muirhead cone $\mathcal{M}_{C}$.
+It should be noted that the Muirhead semiring $\mathcal{M}_{S}$ is strictly larger than the Muirhead cone $\mathcal{M}_{C}$; see Section 4.1 below.
 1.2.3. Weak majorization-type and evaluation positivity
-These inequalities (specifically, the final one in the next result) were first recorded in [10, Theorem 1.14] for Schur polynomials, followed by the other variants that were shown in [4].
-Theorem 1.5 ([4, Theorem 6.5]).
-The following statements are equivalent:
+These inequalities (specifically, the final one in the next result) were first recorded in [9, Theorem 1.14] for Schur polynomials, followed by the other variants that were shown in [3].
+Theorem 1.5 ([3, Theorem 6.7]).
+The following are equivalent for partitions $\lambda,\mu$:
   1. (1)
@@
     $\frac{s_{\lambda}(x+\bm{1})}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x+\bm{1})}{s_{\mu}(\bm{1})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (1.15)
-As noted in [4, Theorem 6.5], these inequalities follow easily from their containment- and majorization- analogues, via the following easy lemma that connects the three partial orders.
-Lemma 1.6 (see [4]).
-If $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, then there exists some $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
+As noted in [3], these inequalities follow easily from their containment- and majorization- analogues, via the following easy lemma that connects the three partial orders.
+Lemma 1.6 ([3, Lemma 6.6]).
+$\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ if and only if there exists some $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
 As mentioned above, [9] provides a different proof of Theorem 1.5 for Schur polynomials.
 1.3. Our work
-This work was motivated by the goal of extending the above characterizations to Jack and even Macdonald polynomials. We refer the reader to:
-  • Theorems 3.1, 3.2, 3.4 and 4.8, connecting (weak) majorization and normalized Jack differences;
+As Jack polynomials unify Schur, monomial, and elementary symmetric polynomials, it is a pleasing fact that Theorem 1.1 holds, unifying several related but disparate positive-coefficient expansion-characterizations of containment in Theorem 1.2.
+In the same vein, it is natural to ask if Theorems 1.3 and 1.5 can also be extended to Jack polynomials for all parameters, or more strongly, to Jack-positivity in the sense of Theorem 1.1(3).
+The goal of this work is to achieve the above unifications in several cases. We also go beyond Jack polynomials, to Macdonald polynomials with parameters $q,t\in(0,1)$. We refer the reader to:
+  • Theorems 3.4, 3.2 and 4.9, connecting (weak) majorization and normalized Jack differences;
   • Theorem 4.4 for the Muirhead semiring and Jack polynomial differences;
-  • Theorems 5.6 and 5.8 for (weak) majorization via Macdonald polynomials.
+  • Theorem 5.5 for majorization via Macdonald polynomials.
+Our results also fit into—and contribute to—the framework of duality discussed in [3]: given a partial order $\mathrel{\succcurlyeq}_{1}$ on symmetric polynomials and a basis $\{b_{\lambda}(x)\}$ of them, as well as a partial order $\mathrel{\succcurlyeq}_{2}$ on partitions, we say that $\mathrel{\succcurlyeq}_{1}$ and $\mathrel{\succcurlyeq}_{2}$ are dual for the family $b_{\lambda}$ if the following holds:
+$b_{\lambda}\mathrel{\succcurlyeq}_{1}b_{\mu}\ \text{if and only if}\ \lambda\mathrel{\succcurlyeq}_{2}\mu.$ (1.16)
+In [3], two of us showed that the containment order $\mathrel{\succcurlyeq}_{2}$ is dual to “Jack expansion positivity” (wherein $p(x)\mathrel{\succcurlyeq}_{1}q(x)$ means pqp-q is a positive combination of Jack polynomials—for a specific choice of parameter or for an indeterminate parameter).
+In this spirit, Theorem 1.3 asserts that majorization, or the dominance order $\mathrel{\succcurlyeq}$ on partitions (which we take to be $\mathrel{\succcurlyeq}_{2}$) is dual to “evaluation positivity” on the positive orthant: $p\mathrel{\succcurlyeq}_{1}q$ if $p(x)\geqslant q(x)$ for all $x\in(0,\infty)^{n}$—this is for normalized Schur, or monomial, or other specific families of symmetric functions.
+Similarly, Theorem 1.5 shows the duality between the weak dominance order $\mathrel{\succcurlyeq}_{2}$ (which is now $\mathrel{\succcurlyeq_{\mathrm{w}}}$) and evaluation positivity for the shifted normalized bases $f_{\lambda}(x+{\bf 1})$ (with the same $f$ as previously) on the positive orthant.
+It is our goal in this work to extend both of these dualities to Jack polynomials with arbitrary parameter—and even more strongly, to Macdonald polynomials.
 2. Preliminaries
@@
 2.1. Partitions
-An (integer) partition is an infinite tuple $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of weakly-decreasing non-negative integers with finitely many nonzero entries.
+A partition is an infinite tuple $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of weakly-decreasing non-negative integers with finitely many nonzero entries.
 The non-zero $\lambda_{i}$’s are called the parts of $\lambda$.
@@
 and that the numerator (and hence $s_{\lambda}$) is a well-defined function for any real $n$-tuple $\lambda$.
-Each of $\{\,m_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,e_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,p_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ and $\{\,s_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ is a $\mathbb{Q}$-basis of $\Lambda_{n}$ [18, Chapter I].
+Each of $\{\,m_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,e_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,p_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ and $\{\,s_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ is a $\mathbb{Q}$-basis of $\Lambda_{n}$ [17, Chapter I].
 2.3. Jack polynomials and Macdonald polynomials
@@
 $\langle p_{\lambda},p_{\mu}\rangle_{\tau}\coloneqq\delta_{\lambda}z_{\lambda}\tau^{-\ell(\lambda)},\quad\langle p_{\lambda},p_{\mu}\rangle_{q,t}\coloneqq\delta_{\lambda}z_{\lambda}\prod_{i=1}^{\ell(\lambda)}\frac{1-q^{\lambda_{i}}}{1-t^{\lambda_{i}}},$ (2.8)
-where $\lambda,\mu\in\mathcal{P}_{n}^{\prime}$, and $z_{\lambda}$ is a constant given in [18, p.
+where $\lambda,\mu\in\mathcal{P}_{n}^{\prime}$, and $z_{\lambda}$ is a constant given in [17, p.
 24].
@@
 $\mathbb{F}_{\geqslant 0}\coloneqq\left\{\,\frac{f}{g}\mathclose{}\nonscript\;\middle|\penalty 500\nonscript\;\mathopen{}f,g\in\mathbb{Z}_{\geqslant 0}[\tau],\ g\neq 0\,\right\},\quad\mathbb{F}_{\geqslant 0}^{\mathbb{R}}\coloneqq\left\{\,\frac{f}{g}\mathclose{}\nonscript\;\middle|\penalty 500\nonscript\;\mathopen{}f,g\in\mathbb{R}_{\geqslant 0}[\tau],\ g\neq 0\,\right\}.$ (2.11)
-In the Macdonald case, we define the cone of positivity $\mathbb{F}_{\geqslant 0}$ and its real extension $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ as
-$\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\},\quad\mathbb{F}_{\geqslant 0}^{\mathbb{R}}\coloneqq\{\,f\in\mathbb{R}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t>1\,\}.$ (2.12)
+In the Macdonald case, we define the cone of positivity $\mathbb{K}_{\geqslant 0}$ and its real extension $\mathbb{K}_{\geqslant 0}^{\mathbb{R}}$ as
+$\mathbb{K}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(0,1)\,\},\quad\mathbb{K}_{\geqslant 0}^{\mathbb{R}}\coloneqq\{\,f\in\mathbb{R}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(0,1)\,\}.$ (2.12)
 3. Jack polynomial conjectures and their partial resolutions
@@
 Given similar manifestations of (weak) majorization in Theorems 1.3 and 1.5, it is natural to expect the analogues of Theorem 1.1 to also hold for (weak) majorization.
-At the same time, one would expect Theorem 3.2 to also generalize.
-Thus, we begin with the following upgradation of [4, Conjecture 1(1)] to majorization.
+Thus, we begin with the following upgradation of [3, Conjecture 1(1)] to majorization.
 For reasons of exposition, we write the assertion that $\lambda\mathrel{\succcurlyeq}\mu$ at the end:
 Conjecture 1 (CGS Conjecture for Jack polynomials).
-Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
+The following are equivalent for partitions $\lambda$ and $\mu$ :
   1. (1)
@@
   2. (2)
-    For some fixed $\tau_{0}\in[0,\infty]$, we have
+    For every $\tau_{0}\in[0,\infty]$, we have
     $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (3.2)
   3. (3)
-    For some fixed $\tau_{0}\in[0,\infty]$, we have
+    For some $\tau_{0}\in[0,\infty]$, we have
     $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in(0,1)^{n}\cup(1,\infty)^{n}.$ (3.3)
@@
     $\lambda$ majorizes $\mu$.
-In exact parallel, we next upgrade [4, Conjecture 1(2)] to weak majorization.
+Remark 3.1.
+We note two ways in which our conjecture (and results below) work with weaker hypotheses than [4] for instance.
+Firstly, [4] imposed the extra hypothesis $|\lambda|=|\mu|$—see e.g.
+Theorem 1.3.
+We do not do so, instead working with arbitrary partitions $\lambda,\mu$.
+Second, the assertions in Theorem 1.3 all claim that an inequality of normalized symmetric differences implies majorization, when evaluated at all test vectors $x$ in $0,\infty)^{n}$.
+In contrast, Conjecture 1 (and results below) work only with $x\in(0,1)^{n}\cup(1,\infty)^{n}$.
+In Section 3.1 below, we will further shrink the test set, to a countable set of test vectors $x$.
+This strengthens even the previously known characterizations of majorization, by Muirhead, Cuttler–Greene–Skandera, and others.
+In exact parallel, we next upgrade [3, Conjecture 1(2)] to weak majorization.
 Conjecture 2 (KT Conjecture for Jack polynomials).
@@
   2. (2)
-    For some fixed $\tau_{0}\in[0,\infty]$, we have
+    For every $\tau_{0}\in[0,\infty]$, we have
+    $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (3.5)
+  3. (3)
+    For some $\tau_{0}\in[0,\infty]$, we have
     $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (3.6)
@@
     $\lambda$ weakly majorizes $\mu$.
-The next result shows “most of” the implications in these conjectures, leaving one unproved implication.
+In the remaining parts of this section, we prove the following partial results:
 Theorem 3.2.
-[…]
-3.1. The proof, using majorization via log-positive and log-negative sequences
-Before showing Theorem 3.1, we take a closer look at Theorem 1.3.
-If $\lambda\mathrel{\succcurlyeq}\mu$, then the inequalities in the final four parts of the theorem hold over the entire orthant.
-However, to deduce $\lambda\mathrel{\succcurlyeq}\mu$, it suffices to work with just two sequences $x_{n}$ – one in the “log-positive” orthant $(1,\infty)^{n}$ and the other in the log-negative orthant $(0,1)^{n}$ – as opposed to all $x\in[0,\infty)^{n}$.
-This provides additional characterizations of (weak) majorization, and is the content of (the proof of) the next result.
+Conjectures 1 and 2 hold for Jack polynomials in $n=2$ variables, as well as for any $n$ and $\mu={\bf 1}_{|\lambda|}$.
+Moreover, the assertions (2)–(4) are equivalent in either conjecture whenever $\lambda$ and $\mu$ have at most two parts (with any number of variables).
+Remark 3.3.
+An analogue of the implication $(4)\Rightarrow(2)$ in Conjecture 1 was formulated by McSwiggen and Novak for Heckman–Opdam polynomials for arbitrary root systems, and proved in semisimple rank 1, subsuming the case of Jack polynomials in two variables.
+See [18, Conjecture 4.7, Proposition 4.10].
+Although we only work in type $A$, our results for two variables, as well as our Conjectures 1 and 2 and later Conjectures 3, 5 and 4, go considerably beyond [18].
+3.1. The reduction to $(4)\Rightarrow(1)$
+We first show “most of” the implications in Conjectures 1 and 2, reducing them to only one unproved implication.
 Theorem 3.4.
-If any of the final four inequalities in Theorem 1.3 holds for all $x$ in $(0,1)^{n}\cup(1,\infty)^{n}$, then $\lambda\mathrel{\succcurlyeq}\mu$. (The converse follows from Theorem 1.3.)
+In both Conjectures 1 and 2, each part implies the next. Moreover, the remaining cyclic implication $(4)\Rightarrow(1)$ in Conjecture 2 follows from $(4)\Rightarrow(1)$ in Conjecture 1.
 Proof.
-The assertion for normalized Schur differences was shown in [10].
-In a similar vein, we consider these inequalities for the monomial symmetric polynomials.
-Since $M_{\lambda}\geqslant M_{\mu}$ on $(1,\infty)^{n}$ (and these functions are continuous), it follows from Theorem 1.5 that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$.
-Specifically, if $\lambda$ does not weakly majorize $\mu$, then there is some $k$ such that $\sum_{i=1}^{k}\lambda_{i}<\sum_{i=1}^{k}\mu_{i}$.
-Now fix scalars $u_{i}>1$, and consider the sequence of vectors
-$x(t)\coloneqq(tu_{1},\dots,tu_{k},u_{k+1},\dots,u_{n})\in(1,\infty)^{n}$
-for a sequence $1\leqslant t=t_{m}\uparrow\infty$.
-Then as polynomials in $t$, $M_{\lambda}(x(t))$ has degree $\sum_{i=1}^{k}\lambda_{i}$, while $M_{\mu}(x(t))$ has degree $\sum_{i=1}^{k}\mu_{i}$, so $M_{\lambda}(x(t_{m}))-M_{\mu}(x(t_{m}))\to$-$\infty$ as $m\to\infty$, which is false.
-Similarly, since $M_{\lambda}\geqslant M_{\mu}$ on $(0,1)^{n}$, we claim that $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$.
-Indeed, if not, then there is some $k$ such that $\sum_{i=0}^{k-1}\lambda_{n-i}<-\sum_{i=0}^{k-1}\mu_{n-i}$.
-Now fix scalars $u_{i}\in(0,1)$ and consider the sequence $y(t)\coloneqq(u_{1}/t,\dots,u_{k}/t,u_{k+1},\dots,u_{n})\in(0,1)^{n}$ for a sequence $1\leqslant t=t_{m}\uparrow\infty$.
-Then
-$M_{\lambda}(y(t))\sim O(t^{-\lambda_{n}-\cdots-\lambda_{n-k+1}}),\qquad M_{\mu}(y(t))\sim O(t^{-\mu_{n}-\cdots-\mu_{n-k+1}}),$ (3.6)
-so $M_{\lambda}(y(t_{m}))-M_{\mu}(y(t_{m}))\to$-$\infty$ as $m\to\infty$, which is again false.
-It follows that if $M_{\lambda}\geqslant M_{\mu}$ on two sequences $x(t_{m})\in(1,\infty)^{n}$ and $y(t_{m})\in(0,1)^{n}$, then $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$.
-But these are together equivalent to $\lambda\mathrel{\succcurlyeq}\mu$.
-The same proof works if one considers power-sums, since $p_{\lambda}(x(t))$ is a polynomial in $t$ of degree $\sum_{i=1}^{k}\lambda_{i}$ as above, and the same claim as in Eq.
-3.6 for $p_{\lambda}(y(t))/p_{\lambda}(\bm{1})$.
-Ditto for the “normalized (transposed) elementary symmetric polynomials” $e_{\lambda^{\prime}}(x)/e_{\lambda^{\prime}}(\bm{1})$.
-∎
-We can now prove Theorem 3.1.
-Proof of Theorem 3.1.
-[…]
-Hence, to settle both conjectures it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq.
-3.1 and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ $\implies$ Eq.
-3.4.
-We end by showing that the latter follows from the former.
-Indeed, by Section 1.2.3 there exists $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
-Now write (suppressing the $\tau$)
+$\bullet$ In both Conjectures 1 and 2, the implications $(1)\Rightarrow(2)\Rightarrow(3)$ are evident.
+$\bullet$ We now prove $(3)\Rightarrow(4)$ in both conjectures.
+First, assume that $\lambda$ does not weakly majorize $\mu$, namely, $d_{k}(\lambda)\coloneqq\lambda_{1}+\cdots+\lambda_{k}<\mu_{1}+\cdots+\mu_{k}=d_{k}(\mu)$ for some $k$. Let
+$x(t)=(\underbrace{2t,\dots,2t}_{\text{$k$ times}},\underbrace{2,\dots,2}_{\text{$n-k$ times}})\in(1,\infty)^{n},\quad t>1.$
+Then as polynomials in $t$, $\frac{P_{\lambda}(x(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}$ and $\frac{P_{\mu}(x(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ have degree $d_{k}(\lambda)$ and $d_{k}(\mu)$, respectively, since the leading term (in the majorization order) of $P_{\lambda}$ is $m_{\lambda}$.
+Now, let $t\to\infty$, we see that
+$\frac{P_{\lambda}(x(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\to-\infty,$
+contradicting Eqs. 3.3 and 3.6 (in Eq. 3.6, we view the argument $x(t)$ as $(x(t)-\bm{1})+\bm{1}$, with $x(t)-\bm{1}\in[0,\infty)^{n}$).
+Second, assume that $\lambda\coloneqq(-\lambda_{n},\dots,-\lambda_{1})$ does not weakly majorize $\mu=(-\mu_{n},\dots,-\mu_{1})$, namely, $s_{k}(\lambda)\coloneqq\lambda_{n}+\dots+\lambda_{n-k+1}>\mu_{n}+\dots+\mu_{n-k+1}=s_{k}(\mu)$.
+Let
+$y(t)=\Big(\underbrace{\tfrac{1}{2t},\dots,\tfrac{1}{2t}}_{k\text{ times}},\underbrace{\tfrac{1}{2},\dots,\tfrac{1}{2}}_{n-k\text{ times}}\Big)\in(0,1)^{n},\qquad t>1.$
+Let $\nu$ be a partition. Then
+$m_{\nu}(y(t))=\Big(\tfrac{1}{2}\Big)^{|\nu|}\sum_{\beta\in\mathfrak{S}_{n}\cdot\nu}t^{-(\beta_{1}+\cdots+\beta_{k})}.$
+The minimum of $\beta_{1}+\cdots+\beta_{k}$ over all permutations $\beta$ of $\nu$ is attained by placing the $k$ smallest parts of $\nu$ into the first $k$ coordinates, so this minimum equals $s_{k}(\nu)=\nu_{n-k+1}+\cdots+\nu_{n}$.
+Therefore there exists a constant $A_{\nu,k}>0$ such that
+$m_{\nu}(y(t))=A_{\nu,k}\,t^{-s_{k}(\nu)}+O(t^{-s_{k}(\nu)-1}),\qquad t\to\infty.$ (3.7)
+Note that Jack polynomials expand positively into monomials:
+$P_{\lambda}(x;\tau_{0})=m_{\lambda}(x)+\sum_{\nu\prec\lambda}c_{\lambda\nu}m_{\nu}(x),\qquad c_{\lambda\nu}\geqslant 0.$
+For $\nu\prec\lambda$, we have
+$s_{k}(\nu)=\nu_{n}+\dots+\nu_{n-k+1}\geqslant\lambda_{n}+\dots+\lambda_{n-k+1}=s_{k}(\lambda).$
+Hence there exists a constant $A_{\lambda,k}^{\prime}>0$ such that
+$P_{\lambda}(y(t);\tau_{0})=A_{\lambda,k}^{\prime}t^{-s_{k}(\lambda)}+O\big(t^{-s_{k}(\lambda-1}\big),\qquad t\to\infty,$ (3.8)
+and similarly for $P_{\lambda}(y(t);\tau_{0})$. Then we have
+$\frac{P_{\lambda}(y(t);\tau_{0})/P_{\lambda}(\bm{1};\tau_{0})}{P_{\mu}(y(t);\tau_{0})/P_{\mu}(\bm{1};\tau_{0})}\sim\frac{A_{\lambda,k}^{\prime}/P_{\lambda}(\bm{1};\tau_{0})}{A_{\mu,k}^{\prime}/P_{\mu}(\bm{1};\tau_{0})}\,t^{-(s_{k}(\lambda)-s_{k}(\mu))}\longrightarrow 0,\qquad t\to\infty.$
+Thus for all sufficiently large $t$,
+$\frac{P_{\lambda}(y(t);\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(y(t);\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}<0,$
+contradicting Eq. 3.3, hence $\lambda$ weakly dominates $\mu$.
+To summarize, we see that Eq.
+3.6 implies $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, and that Eq.
+3.3 implies $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ and $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}-\mu$, forcing $|\lambda|=|\mu|$ and hence $\lambda\mathrel{\succcurlyeq}\mu$.
+These are exactly the implications $(3)\Rightarrow(4)$ in the conjectures.
+$\bullet$ Finally, we prove that Conjecture 2 follows from Conjecture 1.
+Assume Conjecture 2(4), namely, $\lambda$ weakly majorizes $\mu$. Then by Section 1.2.3, there exists $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$. Now write (suppressing the $\tau$)
 $\frac{P_{\lambda}(x+\bm{1})}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x+\bm{1})}{P_{\mu}(\bm{1})}=\left(\frac{P_{\lambda}(x+\bm{1})}{P_{\lambda}(\bm{1})}-\frac{P_{\nu}(x+\bm{1})}{P_{\nu}(\bm{1})}\right)+\left(\frac{P_{\nu}(x+\bm{1})}{P_{\nu}(\bm{1})}-\frac{P_{\mu}(x+\bm{1})}{P_{\mu}(\bm{1})}\right).$
 The first difference is Jack-positive by Theorem 1.1, and in particular, in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ when evaluated at $x\in[0,\infty)^{n}$.
-The second difference is in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ by the assumption that Eq.
-3.1 holds.
+The second difference is in $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ by Eq.
+3.1 (for the pair $\nu\mathrel{\succcurlyeq}\mu$).
 ∎
 Remark 3.5.
-As a consequence of Theorem 3.1, it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1. Note that for $\tau_{0}=0,1,\infty$, this is already known by Theorem 1.3.
-3.2. Reductions of the conjectures; proof for two variables
-In this subsection, we reduce Conjecture 1 to some important special cases. We also work out an example, which shows, via the reductions, that the conjecture holds when $n\leqslant 2$:
-Theorem 3.4.
-Conjectures 1 and 2 hold when $\lambda$ and $\mu$ are of length at most 2.
-We begin with some preliminary results and reductions, which hold for all $n$.
+As a consequence of Theorem 3.4, to completely settle Conjectures 1 and 2, it suffices to show that $\lambda\mathrel{\succcurlyeq}\mu$ $\Rightarrow$ Eq.
+3.1.
+Note that for $\tau_{0}=0,1,\infty$, this is already known by Theorem 1.3.
+Remark 3.6.
+As mentioned in Section 3, we now reduce the test set $(0,1)^{n}\cup(1,\infty)^{n}$ to a potentially “minimum”, countable test set.
+Let $S\subseteq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}>\dots>x_{n}>0\,\}$ be any subset.
+We say that $S$ has asymptotes in direction $i$ ($1\leqslant i\leqslant n-1$), if there exists a constant $M>0$ and a sequence $(x^{(1)},x^{(2)},\dots)$ in $S$, such that
+$\begin{dcases}\lim_{m\to\infty}\frac{x_{i}^{(m)}}{x_{i+1}^{(m)}}\to\infty,\\ \lim_{m\to\infty}\frac{x_{j}^{(m)}}{x_{j+1}^{(m)}}<M,\quad j=1,\dots,i-1,i+1,\dots,n-1,\\ \lim_{m\to\infty}x_{n}^{(m)}<M.\end{dcases}$
+We say that $S$ has asymptotes in direction $n$, if there exists a constant $M>0$ and a sequence $(x^{(1)},x^{(2)},\dots)$ in $S$, such that
+[…]
+To prove Eq. 3.2 for some fixed $\tau_{0}$, one may assume without lost of generality that $x_{1}>\dots>x_{n}>0$ by the symmetry of Jack polynomials. We now have:
 Proposition 3.7.
-For a one-row partition $(d,0^{n-1})$, the Jack polynomials are given by
+Let $S\subseteq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}>\dots>x_{n}>0\,\}$ be a subset that has asymptotes in all directions $1,\dots,n$. If
+$\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\geqslant 0,\quad\forall x\in S,$ (3.9)
+then $\lambda\mathrel{\succcurlyeq}\mu$.
+Proof.
+For direction $k=1,\dots,n-1$ and for $k=n$ with $\lim_{m\to\infty}x_{n}^{(m)}=\infty$, we see that $x_{1}^{(m)},\dots,x_{k}^{(m)}$ approach infinity at the same rate and $x_{k+1}^{(m)},\dots,x_{n}^{(m)}$ are bounded.
+Then the “order” of $P_{\lambda}(x^{(m)};\tau_{0})$ is $\lambda_{1}+\cdots+\lambda_{k}$, and similarly for $P_{\mu}(x^{(m)};\tau_{0})$.
+Since Eq.
+3.9 holds for all $x\in S$, letting $m\to\infty$, we see that $\lambda_{1}+\cdots+\lambda_{k}\geqslant\mu_{1}+\cdots+\mu_{k}$.
+As this holds for all $k$, we have $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$.
+For direction $k=n$ with $\lim_{m\to\infty}x_{n}^{(m)}=0$, we see that $x_{1}^{(m)},\dots,x_{n}^{(m)}$ approach 0 at the same rate, and the “order” of $P_{\lambda}(x;\tau_{0})$ is $|\lambda|$.
+Since Eq.
+3.9 holds for all $x\in S$, we see that $|\lambda|\leqslant|\mu|$.
+Hence we conclude that $|\lambda|=|\mu|$ and $\lambda\mathrel{\succcurlyeq}\mu$.
+∎
+3.2. Reductions of the conjectures
+In this subsection, we further reduce Conjecture 1 to some important special cases.
+3.2.1. Adding columns
+Proposition 3.8.
+For a one-row partition $(d)=(d,0^{n-1})$, the Jack polynomials are given by
 $P_{(d,0^{n-1})}(x;\tau)=\frac{d!}{(\tau)_{{d}}}\cdot\sum_{|\eta|=d}\frac{(\tau)_{{\eta}}}{\eta!}x^{\eta},$ (3.10)
@@
 The first assertion follows from the combinatorial formula (see e.g.
-[4, Eq.
+[3, Eq.
 (2.10)]).
-The second follows from [18, VI.
+The second follows from [17, VI.
 (10.20)] (note, $\alpha$ is $\frac{1}{\tau}$ here), or the Chu–Vandermonde identity.
 ∎
-Lemma 3.6.
-Let $\lambda\in\mathcal{P}_{n}$ be such that $\lambda_{n}>0$. Then
+Proposition 3.9.
+Let $\lambda\in\mathcal{P}_{n}$ be such that $\lambda_{n}>0$. Write $\lambda-\bm{1}=(\lambda_{1}-1,\dots,\lambda_{n}-1)$. Then
 $P_{\lambda}(x;\tau)=x_{1}\cdots x_{n}\cdot P_{\lambda-\bm{1}}(x;\tau).$ (3.12)
 Proof.
-See [18, VI. (4.17)] for the Macdonald case. ∎
+See [17, VI. (4.17)] for the Macdonald case. ∎
 Lemma 3.10 (Adding columns).
-To prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu$ $\implies$ Eq. 3.1), it suffices to assume $\lambda_{n}=0$.
+To prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu$ $\Rightarrow$ Eq. 3.1), it suffices to assume $\lambda_{n}=0$.
 Proof.
@@
 3.1 holds for the pair $(\lambda-\lambda_{n}\bm{1},\mu-\lambda_{n}\bm{1})$.
-By Section 3.2, it holds for the pair $(\lambda,\mu)$ as well.
+By Section 3.2.1, Eq.
+3.1 holds for the pair $(\lambda,\mu)$ as well.
 ∎
-Lemma 3.8 (Appending 0).
-If Eq. 3.2 with n1n-1 variables holds for the pair $(\lambda,\mu)$, then Eq. 3.2 with $n$ variables holds for the pair $((\lambda,0),(\mu,0))$ as well.
-See the next subsection for the proof.
+3.2.2. Raising operator
 Lemma 3.11 (Raising operator).
-To prove that $\lambda\mathrel{\succcurlyeq}\mu$ implies Eq.
-3.1, it suffices to assume that $\lambda$ and $\mu$ are adjacent in the majorization order.
-In that case, $\lambda=R_{ij}(\mu)$ for some $i<j$, where $R_{ij}(\mu)=(\dots,\mu_{i}+1,\dots,\mu_{j}-1,\dots)$ is the raising operator, see [18, pp.
+To prove Conjecture 1 ($\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq.
+3.1), it suffices to assume that $\lambda$ and $\mu$ are adjacent in the majorization order.
+In that case, $\lambda=R_{ij}(\mu)$ for some $i<j$, where $R_{ij}(\mu)=(\dots,\mu_{i}+1,\dots,\mu_{j}-1,\dots)$ is the raising operator, see [17, pp.
 8].
 Proof.
-The first claim follows by transitivity. The second claim is [18, I. (1.16)]. ∎
+The first claim follows by transitivity. The second claim is [17, I. (1.16)]. ∎
 Note that if $\lambda$ and $\mu$ are majorization-adjacent, then they are of exactly two possible forms:
@@
   • $\lambda=(\nu,k+1,k^{m},k-1,\eta)$ and $\mu=(\nu,k^{m+2},\eta)$, where $k,m\geqslant 1$, and $\nu$ and $\eta$ (possibly empty) are partitions that make $\lambda$ and $\mu$ valid partitions.
-By the reduction lemmas above, in order to prove the weaker statement $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq. 3.2, we may assume that
-  • either $\lambda=(\nu,k,0)$ and $\mu=(\nu,k-1,1)$, where $\nu\in\mathcal{P}_{n-2}$ such that $\nu_{n-2}\geqslant k\geqslant 2$;
-  • or $\lambda=(\nu,2,1^{m},0)$ and $\mu=(\nu,1^{m+2})$, where $1\leqslant m\leqslant n-2$, and $\nu\in\mathcal{P}_{n-m-2}$ such that $\nu_{n-m-2}\geqslant 2$.
-The following toy example studies the simplest case of the first kind of these pairs $(\lambda,\mu)$.
-Example 3.10 (Toy Example for CGS).
-Let $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. Suppressing the argument $\tau$ from $P_{\lambda}(\cdot),P_{\mu}(\cdot)$ henceforth, we have:
-$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}x_{1}^{d-i}x_{2}^{i}$ (3.10)
-and
-$\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$ $=\frac{x_{1}x_{2}}{1}\frac{P_{\mu-\bm{1}}(x)}{P_{\mu-\bm{1}}(\bm{1})}=\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}x_{1}^{d-i-1}x_{2}^{i+1}.$ (3.11)
-Setting $x=(1,s)$, $s>0$, we have
-$f(s)$ $\coloneqq\frac{P_{\lambda}(1,s)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(1,s)}{P_{\mu}(\bm{1})}$
-$=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}s^{i}-\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}s^{i+1}$
-$=\sum_{i=0}^{d}\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right)s^{i}\eqqcolon\sum_{i}f_{i}s^{i}.$
-Define for $0\leqslant i\leqslant d-2$,
-$g_{i}\coloneqq\binom{d-2}{i}\frac{(\tau+d-1)}{\tau}\frac{(\tau)_{{i+1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}\in\mathbb{F}_{>0},$
-and let $g_{-1}=g_{-2}=g_{d-1}=g_{d}=0$. Then one can show that
-$f_{i}=g_{i-2}-2g_{i-1}+g_{i},\text{\quad or equivalently, \quad}\sum_{j=i+1}^{d}(j-i-1)f_{j}=g_{i}.$
-This yields Eq. 3.1, because we have for any $s>0$:
-$f(s)=(s-1)^{2}g(s)\coloneqq(s-1)^{2}\sum_{i=0}^{d-2}g_{i}s^{i}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}.$
-Combining the results and the example above, Theorem 3.4 follows. ∎
-3.3. Okounkov–Olshanski’s formula
-As promised above, in this subsection we show Section 3.2.
+3.2.3. Appending 0
+We first recall the integration formula due to Okounkov–Olshanski [20].
 Set $\mathbb{R}^{n}_{\geqslant}\coloneqq\{\,x\in\mathbb{R}^{n}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}x_{1}\geqslant\cdots\geqslant x_{n}\,\}$.
-For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, write $y\prec x$ (interlacing) if
+For $x\in\mathbb{R}^{n}_{\geqslant}$ and $y\in\mathbb{R}^{n-1}_{\geqslant}$, write $y\prec_{\mathrm{int}}x$ (interlacing) if
 $x_{1}\geqslant y_{1}\geqslant x_{2}\geqslant y_{2}\geqslant\cdots\geqslant x_{n-1}\geqslant y_{n-1}\geqslant x_{n}.$ (3.13)
 Let $\lambda\in\mathcal{P}_{n-1}$ be a partition of length at most n1n-1, and $(\lambda,0)\in\mathcal{P}_{n}$.
-Then [21, pp.
+Then [20, pp.
 77] (where $\theta$ is our $\tau$) shows that for $x\in\mathbb{R}^{n}_{\geqslant}$,
-$P_{(\lambda,0)}(x)=\frac{1}{C(\lambda,n;\tau)V(x)^{2\tau-1}}\int_{y\prec x}P_{\lambda}(y)V(y)\Pi(x,y)^{\tau-1}\mathop{}\!\mathrm{d}y,$ (3.13)
+$P_{(\lambda,0)}(x)=\frac{1}{C(\lambda,n;\tau)V(x)^{2\tau-1}}\int_{y\prec_{\mathrm{int}}x}P_{\lambda}(y)V(y)\Pi(x,y)^{\tau-1}\mathop{}\!\mathrm{d}y,$ (3.14)
 where
@@
 and $\mathop{\mathrm{B}}(\alpha,\beta)$ is the Beta function.
-We note that Okounkov–Olshanski’s formula has a better form when changed to the following normalization.
+Okounkov–Olshanski’s formula has a better form when changed to the unital normalization.
 Proposition 3.12.
@@
 Then
-$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}=\int_{y\prec x}\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}K(x,y)\mathop{}\!\mathrm{d}y,\qquad\forall\lambda\in\mathcal{P}_{n-1},\ x\in\mathbb{R}^{n}_{\geqslant}.$ (3.15)
-In particular, $\int_{y\prec x}K(x,y)\mathop{}\!\mathrm{d}y=1$.
+$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}=\int_{y\prec_{\mathrm{int}}x}\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}K(x,y)\mathop{}\!\mathrm{d}y,\qquad\forall\lambda\in\mathcal{P}_{n-1},\ x\in\mathbb{R}^{n}_{\geqslant}.$ (3.16)
+In particular, $\int_{y\prec_{\mathrm{int}}x}K(x,y)\mathop{}\!\mathrm{d}y=1$.
 Proof.
-By [18, VI. (10.20)], we have for any $\mu\in\mathcal{P}_{n}$,
+By [17, VI. (10.20)], we have for any $\mu\in\mathcal{P}_{n}$,
 $P_{\mu}(\bm{1}_{n})=\prod_{s\in\mu}\frac{a_{\mu}^{\prime}(s)+(n-l_{\mu}^{\prime}(s))\tau}{a_{\mu}(s)+(l_{\mu}(s)+1)\tau}.$
@@
 $\frac{P_{(\lambda,0)}(\bm{1}_{n})}{P_{\lambda}(\bm{1}_{n-1})}\frac{1}{C(\lambda,n;\tau)}=\prod_{i=1}^{n-1}\frac{\Gamma(\tau)\Gamma((n-i)\tau)}{\Gamma((n-i+1)\tau)}=\frac{\Gamma(\tau)^{n}}{\Gamma(n\tau)}.$
-We are now ready to prove Section 3.2.
-Proof of Section 3.2.
+Lemma 3.13 (Appending 0).
+If Eq. 3.2 holds for the pair $(\lambda,\mu)$ in n1n-1 variables, then Eq. 3.2 holds for the pair $((\lambda,0),(\mu,0))$ in $n$ variables as well.
+Proof.
 Assume that Eq.
-3.1 with n1n-1 variables $y=(y_{1},\dots,y_{n-1})$ holds for the pair $(\lambda,\mu)$.
-We first prove Eq.
-3.2 for $x=(x_{1},\dots,x_{n})$ and the pair $((\lambda,0),(\mu,0))$.
-For $\tau>0$, $x\in\mathbb{R}_{\geqslant 0}^{n}$, since the kernel $K(x,y)>0$, we have
-$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}-\frac{P_{(\mu,0)}(x)}{P_{(\mu,0)}(\bm{1}_{n})}=\int_{y\prec x}\left(\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}-\frac{P_{\mu}(y)}{P_{\mu}(\bm{1}_{n-1})}\right)K(x,y)\mathop{}\!\mathrm{d}y>0.$
-Since this is symmetric in $x$, Eq. 3.2 holds for all $x\in[0,\infty]^{n}$. When $\tau=0$, Jack polynomials reduce to monomial symmetric polynomials and Eq. 3.2 is simply Muirhead’s inequality. ∎
+3.1 holds for the pair $(\lambda,\mu)$ in n1n-1 variables $y=(y_{1},\dots,y_{n-1})$.
+When $\tau=0$, Jack polynomials reduce to monomial symmetric polynomials and Eq.
+3.2 is simply Muirhead’s inequality.
+For every $\tau>0$, $x\in\mathbb{R}_{\geqslant 0}^{n}$, since the kernel $K(x,y)>0$, we have
+$\frac{P_{(\lambda,0)}(x)}{P_{(\lambda,0)}(\bm{1}_{n})}-\frac{P_{(\mu,0)}(x)}{P_{(\mu,0)}(\bm{1}_{n})}=\int_{y\prec_{\mathrm{int}}x}\left(\frac{P_{\lambda}(y)}{P_{\lambda}(\bm{1}_{n-1})}-\frac{P_{\mu}(y)}{P_{\mu}(\bm{1}_{n-1})}\right)K(x,y)\mathop{}\!\mathrm{d}y>0.$
+Since the LHS is symmetric in $x$, Eq. 3.2 holds for all $x\in[0,\infty]^{n}$. ∎
+3.3. An example and the proof
+Example 3.14.
+Consider $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. Suppressing the parameter $\tau$, we have
+$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}x_{1}^{d-i}x_{2}^{i}$ (3.17)
+and
+$\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$ $=\frac{x_{1}x_{2}}{1}\frac{P_{(d-2,0)}(x_{1},x_{2})}{P_{(d-2,0)}(1,1)}=\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}x_{1}^{d-i-1}x_{2}^{i+1}.$ (3.18)
+Setting $x=(1,s)$, $s>0$, we have
+$f(s)$ $\coloneqq\frac{P_{\lambda}(1,s)}{P_{\lambda}(1,1)}-\frac{P_{\mu}(1,s)}{P_{\mu}(1,1)}$
+$=\sum_{i=0}^{d}\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}s^{i}-\sum_{i=0}^{d-2}\binom{d-2}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i-2}}}{(2\tau)_{{d-2}}}s^{i+1}$
+$=\sum_{i=0}^{d}\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right)s^{i}\eqqcolon\sum_{i}f_{i}s^{i}.$
+Define $g(s)\coloneqq\sum_{i=0}^{d-2}g_{i}s^{i}$, where
+$g_{i}\coloneqq\binom{d-2}{i}\frac{(\tau+d-1)}{\tau}\frac{(\tau)_{{i+1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d}}}\in\mathbb{F}_{\geqslant 0},\quad 0\leqslant i\leqslant d-2.$
+Let $g_{-1}=g_{-2}=g_{d-1}=g_{d}=0$. Then one can show that
+$f_{i}=g_{i-2}-2g_{i-1}+g_{i},$
+hence
+$f(s)=(s-1)^{2}g(s)\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad s>0.$
+Now, Eq. 3.1 follows since
+$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}=x_{2}^{d}\cdot f(x_{1}/x_{2})\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad(x_{1},x_{2})\in(0,\infty)^{n}.$
+Proof of Theorem 3.2.
+In order to prove Conjectures 1 and 2 in the case of two variables, it suffices to prove $\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq.
+3.1 by Theorem 3.4.
+By Section 3.2.2, it suffices to show this implication for the majorization-adjacent pair: $\lambda=(d+l,l)\mathrel{\succcurlyeq}\mu=(d+l-1,l+1)$, $l\geqslant 0$, $d\geqslant 2$.
+By Section 3.2.1, it can be further reduced to $\lambda=(d,0)\mathrel{\succcurlyeq}\mu=(d-1,1)$, and this case was proved in Section 3.3.
+The next assertion – in which $\mu={\bf 1}_{|\lambda|}$ – is proved in Theorems 4.2 and 4.4 and the remark in between them.
+We next show the final assertion.
+By Section 3.2.3, the weaker implication $(4)\Rightarrow(2)$ in Conjecture 1 holds for partitions $\lambda$ and $\mu$ of length at most 2 and any number of variables.
+We are now done by Theorem 3.4.
+∎
 3.4. Supporting evidence: Kadell’s integral
-We end this section by providing supporting evidence for the assertion $(4)\implies(2)$ in Conjecture 1.
-We begin by recalling Kadell’s integral in [18, VI.
+We end this section by providing supporting evidence for the assertion $(4)\Rightarrow(2)$ in Conjecture 1.
+We begin by recalling Kadell’s integral in [17, VI.
 Ex.
@@
 If moreover $r$ and $s$ are rational numbers, then $\tilde{I}_{\lambda}-\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}$.
-This supports $(4)\implies(2)$ in Conjecture 1, since for each fixed $\tau=\tau_{0}\in(0,\infty)$, if the difference $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}$ is positive over $0,1]^{n}$, then so is its integral against Kadell’s density
-$\frac{1}{I_{(0^{n})}}|V(x)|^{2\tau}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}.$
 Proof.
-Note that $\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$.
-By Section 3.2, we may assume there exist $1\leqslant i<j\leqslant n$ such that $\lambda=R_{ij}(\mu)$, i.e., $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$ and $\lambda_{k}=\mu_{k}$ for $k\neq i,j$.
+Note that $\tilde{I}_{\mu}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ for $r,s>0$.
+By Section 3.2.2, we may assume there exist $1\leqslant i<j\leqslant n$ such that $\lambda=R_{ij}(\mu)$, i.e., $\lambda_{i}=\mu_{i}+1$, $\lambda_{j}=\mu_{j}-1$ and $\lambda_{k}=\mu_{k}$ for $k\neq i,j$.
 Since $\lambda_{i}-\lambda_{j}-1\geqslant 0$,
@@
 $=\frac{((j-i)\tau+\lambda_{i}-\lambda_{j}-1)((n-1)\tau+s)}{((n-j)\tau+\lambda_{j}+r)((2n-i-1)\tau+\lambda_{i}+r+s-1)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}}.\qed$
+This supports $(4)\Rightarrow(2)$ in Conjecture 1, since for each fixed $\tau_{0}\in(0,\infty)$, if the difference $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is positive over $0,1]^{n}$, then so is its integral against Kadell’s density
+$\frac{1}{I_{(0^{n})}}|V(x)|^{2\tau_{0}}\prod_{i=1}^{n}x_{i}^{r-1}(1-x_{i})^{s-1}>0.$
 4. Jack polynomial differences and the Muirhead semiring
-In this section we prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu\implies$ Eq.
+In this section we prove Conjecture 1 (i.e., that $\lambda\mathrel{\succcurlyeq}\mu\Rightarrow$ Eq.
 3.1) in some more cases.
-In fact, we prove a stronger version of the conjecture, in which positivity is replaced by sums of Muirhead differences (which are positive by Muirhead’s inequality).
+In fact, we prove a stronger version of the conjecture, in which evaluation positivity is replaced by sums and products of Muirhead differences (which are positive by Muirhead’s inequality).
 4.1. Jack differences in the Muirhead semiring
-To remind the reader: consider the outstanding theorem from the introduction (Theorem 1.4) and the ensuing discussion. We upgrade the final statement in that subsection to the following.
-Conjecture 3.
-Fix $\tau_{0}\in[0,\infty]$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$ (and conversely):
-$\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (4.1)
-Note that when $\tau_{0}=0$, Conjecture 3 is trivial as $P_{\lambda}(x;0)=m_{\lambda}(x)$.
-For $\tau_{0}=1$, $P_{\lambda}(x;1)=s_{\lambda}(x)$, and it is (informally) conjectured in [5, Section 7] that the difference in Eq.
-4.1 is in the Muirhead cone $\mathcal{M}_{C}(\mathbb{Q}_{\geqslant 0})$.
-However, we now show that this is false.
-Remark 4.1.
-Conjecture 3 is false if we use the Muirhead cone instead of the semiring. For example, consider $n=3$, $\lambda=(5,2,0)$, and $\mu=(5,1,1)$. We have
-$P(x)$ $\coloneqq 3(2\tau+1)(3\tau+1)(3\tau+2)(3\tau+4)\Bigg(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\Bigg)$
-= $=\;2(\tau+1)(\tau+2)^{2}(2\tau+3)M_{(5,2,0)}-(\tau+1)(\tau+2)(2\tau^{2}+13\tau+12)M_{(5,1,1)}$
-$\quad+6\tau(\tau+1)(\tau+2)(2\tau+3)M_{(4,3,0)}-2\tau(\tau+1)(2\tau^{2}+17\tau+17)M_{(4,2,1)}$
-$\quad-12\tau(\tau+1)(\tau+2)M_{(3,3,1)}-\tau(\tau+1)(10\tau^{2}+\tau-20)M_{(3,2,2)}.$
-If $10\tau^{2}+\tau-20<0$, that is, $0<\tau<\frac{3\sqrt{89}-1}{20}\approx 1.365$, the coefficient of $M_{(3,2,2)}$ is positive.
-Since $(3,2,2)$ is the minimum (in the majorization order) partition that appears in $P(x)$, $P(x)$ cannot lie in the Muirhead cone.
-In particular, this happens for the Schur case $\tau=1$.
-However, this remains consistent with Conjecture 3 because
-$P(x)$ $=(2\tau^{4}+19\tau^{3}+43\tau^{2}+54\tau+24)(M_{(5,2,0)}-M_{(5,1,1)})$
+Recall the Muirhead cone and semiring, defined above Theorem 1.4.
+In [4, Section 7], the authors wondered if normalized Schur differences $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}$ lie in the Muirhead cone, for $\lambda\mathrel{\succcurlyeq}\mu$.
+However, this is not the case as the example below shows.
+It seems that the result would hold if one replaces the Muirhead cone by the semiring—which is strictly larger, as the example also confirms.
+Example 4.1.
+Consider $n=3$, $\lambda=(5,2,0)$, and $\mu=(5,1,1)$. We have
+$70\cdot\Bigg(\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}\Bigg)=10M_{(5,2,0)}-9M_{(5,1,1)}+10M_{(4,3,0)}-8M_{(4,2,1)}-4M_{(3,3,1)}+M_{(3,2,2)}.$
+Note that the coefficient of $M_{(3,2,2)}$ is positive and $(3,2,2)$ is the minimum (in the majorization order) partition that appears in the expansion of $\frac{s_{\lambda}(x)}{s_{\lambda}(\bm{1})}-\frac{s_{\mu}(x)}{s_{\mu}(\bm{1})}$, and hence the difference cannot lie in the Muirhead cone.
+However, this difference indeed lies in the Muirhead semiring, in fact for Jack polynomials with arbitrary parameter $\tau=\tau_{0}$.
+Even more strongly, this holds for the cone $\mathcal{C}=\mathbb{F}_{\geqslant 0}$:
+$\mathrel{\phantom{=}}3(2\tau+1)(3\tau+1)(3\tau+2)(3\tau+4)\Bigg(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\Bigg)$
+$=(2\tau^{4}+19\tau^{3}+43\tau^{2}+54\tau+24)(M_{(5,2,0)}-M_{(5,1,1)})$
 $\mathrel{\phantom{=}}+\tau^{2}(2\tau^{2}+7\tau+7)(M_{(5,2,0)}-M_{(4,3,0)})$
@@
 $\mathrel{\phantom{=}}+6\tau(M_{(2,0,0)}-M_{(1,1,0)})(M_{(3,1,1)}-M_{(2,2,1)}),$
-which is in the Muirhead semiring $\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0})$ but not in the cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$.
+which is in the Muirhead semiring $\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0})$.
+Motivated by this calculation, we make the following conjecture.
+Conjecture 3.
+The following are equivalent for partitions $\lambda$ and $\mu$:
+  1. (1)
+    The Jack difference lies in the Muirhead semiring over $\mathbb{F}_{\geqslant 0}$:
+    $\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0}).$ (4.1)
+  2. (2)
+    For every $\tau_{0}\in[0,\infty]$, the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
+    $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (4.2)
+  3. (3)
+    For some $\tau_{0}\in[0,\infty]$, the Jack difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
+    $\frac{P_{\lambda}(x;\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x;\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (4.3)
+  4. (4)
+    $\lambda$ majorizes $\mu$.
+Theorem 4.2.
+In Conjecture 3, clearly $(1)\Rightarrow(2)\Rightarrow(3)$, and (3) implies Conjecture 1(3), which gives $\lambda\mathrel{\succcurlyeq}\mu$. So Conjecture 3 reduces to showing $(4)\Rightarrow(1)$. ∎
 Remark 4.3.
-Here are three connections of Conjecture 3 to Muirhead’s inequality (Theorem 1.3).
-First, Conjecture 3, together with the inequality implies Conjecture 1.
-Second, the converse to this conjecture holds because of Muirhead’s inequality – see Theorem 1.3(2) – and the argument in the proof of Theorem 3.1.
-Thus Conjecture 3 provides yet another conjectural characterization of majorization.
-Third, the proof of Conjecture 3 for $\mu=(1^{|\lambda|})$ uses Muirhead’s inequality; see below.
-Remark 4.3.
-One of the conjectures above ($(4)\implies(1)$ in Conjecture 1) was formulated in all Lie types by McSwiggen and Novak in an interesting paper [19]; see their Conjecture 4.7 and the discussion on pp.
-3997.
-The authors also showed the “easier half” in [19, Proposition 4.8] using asymptotics, as well as the other half for $\dim V=1$ in their Proposition 4.10.
-While we only work in type $A$, here are four ways in which we go beyond [19]: (a) We show the easier half (see Theorems 3.1 and 3.2) using alternate techniques, which prove the result not just for Jack polynomials but also for power sums (as was suspected in light of Theorem 1.2).
-These are not Jack specializations, so not covered in [19].
-(b) Our method also characterizes weak majorization, again not addressed in [19].
-(c) In fact we show Conjecture 3, which is stronger than Conjecture 1 (via the preceding remark) and which is not in [19].
-Finally, (d) our proof of the “harder half” holds (for two variables) not just for normalized Jack differences, but also (Theorem 5.6) for Macdonald polynomials, which were “hoped for” in [19].
+By Muirhead’s inequality, Conjecture 3(1) implies Conjecture 1(1). Thus Conjecture 3 implies Conjecture 1.
 We now show
 Theorem 4.4.
-Conjectures 1, 2 and 3 hold for all $\lambda$ when $\mu=(1^{|\lambda|})$. Conjecture 3 moreover holds when $\lambda$ and $\mu$ are of length at most 2.
-Proof of Theorem 4.4.
-If $\lambda\mathrel{\succcurlyeq}\mu=(1^{|\lambda|})$, we prove more strongly that the difference lies in the Muirhead cone $\mathcal{M}_{C}(\mathbb{R}_{\geqslant 0})$.
-Note that $\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}$ is an $\mathbb{F}_{\geqslant 0}$-linear combination of $M_{\nu}$ (see the arguments for [18, VI.
-(10.15)] for the expansion of monic Jack into monomials; see also [12] for the expansion of integral Jack into augmented monomials), say $\sum_{\nu}u_{\lambda\nu}M_{\nu}$.
-Then $\sum_{\nu}u_{\lambda\nu}=1$ by our choice of normalization.
-Also, we have $\frac{P_{\mu}}{P_{\mu}(\bm{1})}=M_{\mu}$.
-Hence,
-$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}}{P_{\mu}(\bm{1})}=\sum_{\nu}u_{\lambda\nu}(M_{\nu}-M_{\mu})$ (4.2)
-lies in the Muirhead cone. This shows Conjecture 3, and hence Conjectures 1 and 2 for $\mu=(1^{|\lambda|})$.
-We next come to the case of $\lambda,\mu\in\mathcal{P}_{2}$.
-Here we continue the calculations in Toy Example 3.2.
-Expanding the difference of the normalized Jack polynomials in Muirhead differences and suppressing the argument $\tau$ in them as usual, we obtain:
+Conjectures 1, 2 and 3 hold for all $\lambda$ when $\mu=(1^{|\lambda|})$. Conjecture 3 moreover holds in two variables.
+Proof.
+If $\lambda\mathrel{\succcurlyeq}\mu=(1^{|\lambda|})$, we prove more strongly that the difference lies in the Muirhead cone $\mathcal{M}_{C}(\mathbb{F}_{\geqslant 0})$. Write
+$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}=\sum_{\nu\preceq\lambda}u_{\lambda\nu}M_{\nu},\quad u_{\lambda\nu}\in\mathbb{F}_{\geqslant 0}.$
+Then $\sum_{\nu}u_{\lambda\nu}=1$ by our choice of normalization. Also, we have $\frac{P_{\mu}}{P_{\mu}(\bm{1})}=M_{\mu}$. Hence,
+$\frac{P_{\lambda}}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}}{P_{\mu}(\bm{1})}=\sum_{\nu}u_{\lambda\nu}(M_{\nu}-M_{\mu})\in\mathcal{M}_{C}(\mathbb{F}_{\geqslant 0}).$ (4.4)
+This shows Conjecture 3, and hence Conjectures 1 and 2 for $\mu=(1^{|\lambda|})$.
+We next come to the case of two variables.
+Here we continue the calculations in Section 3.3.
+Expanding the difference of the normalized Jack polynomials in Muirhead differences and suppressing the parameter $\tau$ as usual, we obtain:
 $\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\sum_{i=0}^{d}\left(\binom{d}{i}\frac{(\tau)_{{i}}(\tau)_{{d-i}}}{(2\tau)_{{d}}}-\binom{d-2}{i-1}\frac{(\tau)_{{i-1}}(\tau)_{{d-i-1}}}{(2\tau)_{{d-2}}}\right)x_{1}^{d-i}x_{2}^{i}.$
@@
 Corollary 4.5.
-Fix integer partitions $\lambda,\mu\in\mathcal{P}_{2}$ with $\lambda\succcurlyeq\mu$. For $x=(x_{1},x_{2})$, we have
+Let $x=(x_{1},x_{2})$ and fix $\lambda\mathrel{\succcurlyeq}\mu$ partitions with at most two parts. We have
 $\frac{1}{V(x)^{2}}\left(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\right)\in\bigoplus_{\nu}\mathbb{F}_{\geqslant 0}\cdot s_{\nu}(x).$
 Remark 4.6.
-Notice that Section 4.1 does not hold for higher $n$ – not even for $n=3$, even at $\tau=1$, i.e.
+Notice that Section 4.1 does not hold for higher nn—not even for $n=3$, even at $\tau=1$, i.e.
 for Schur polynomials.
-(This is in parallel to Conjecture 3 also not holding for the Muirhead cone even with $n=3$ and $\tau=1$, as verified in Section 4.1.) For example, working with $(3)\succcurlyeq(1,1,1)$ yields
+(This is in parallel to Conjecture 3 also not holding for the Muirhead cone even with $n=3$ and $\tau=1$, as verified in Section 4.1.) For example, working with $(3)\mathrel{\succcurlyeq}(1,1,1)$ yields
 $s_{(3)}(x_{1},x_{2},x_{3})-s_{(3)}(1,1,1)s_{(1,1,1)}(x_{1},x_{2},x_{3})=\sum_{i}x_{i}^{3}+\sum_{i<j}(x_{i}^{2}x_{j}+x_{j}^{2}x_{i})-9x_{1}x_{2}x_{3},$
@@
 In fact for two variables, not only does the conjecture imply this assertion, but the converse is also true.
-This is because – we claim that – every normalized Muirhead-difference $M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})$ (for $\lambda\mathrel{\succcurlyeq}\mu$), when divided by $V(x)^{2}=(x_{1}-x_{2})^{2}$, is a Schur-positive polynomial with half-integer coefficients.
+This is because—we claim that—every normalized Muirhead-difference $M_{\lambda}(x_{1},x_{2})-M_{\mu}(x_{1},x_{2})$ (for $\lambda\mathrel{\succcurlyeq}\mu$), when divided by $V(x)^{2}=(x_{1}-x_{2})^{2}$, is a Schur-positive polynomial with half-integer coefficients.
 This claim reduces by transitivity to majorization-adjacent partitions $\lambda=(a+b+1,a),\mu=(a+b,a+1)$ (for $a\geqslant 0,b\geqslant 1$), for which it holds by an explicit computation:
@@
 4.2. Jack polynomial differences with varying parameter
-Since $M_{\lambda}(x)=P_{\lambda}(x;0)/P_{\lambda}(\bm{1};0)$, Eq.
-4.1 can be viewed as a positivity property involving Jack polynomial differences with different parameters.
+Since $m_{\lambda}(x)=P_{\lambda}(x;\tau=0)$, Eq.
+4.3 can be viewed as a positivity property involving Jack polynomial differences in two different parameters $\tau$ and 0.
 Let us generalize this idea further.
@@
 Conjecture 4.
-Fix $0\leqslant\sigma\leqslant\tau\leqslant\infty$.
-If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference with parameter $\tau$ is positive in the Jack semiring $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$:
-$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0}).$ (4.3)
-Note that Conjecture 4 is false when $\tau<\sigma$.
-For example, let $n=3$, $\tau=0$, $\sigma=1$, then $M_{(3)}-M_{(21)}=\frac{10}{3}S_{(3)}-4S_{(21)}+\frac{2}{3}S_{(111)}$, where $S_{\lambda}=s_{\lambda}/s_{\lambda}(\bm{1})$.
-Also note that the expansion of $P_{\lambda}(x;\tau)$ into a sum of polynomials $P_{\mu}(x;\sigma)$ with $\sigma\leqslant\tau$ may not be positive.
-For example, when $n=3$, we have (suppressing the $x$)
-$P_{(33)}(\tau)=P_{(33)}(\sigma)+\frac{6(\tau-\sigma)}{(2+\sigma)(2+\tau)}P_{(321)}(\sigma)+\frac{6(\tau-2\sigma)(\tau-\sigma)}{(1+\sigma)(1+2\sigma)(1+\tau)(2+\tau)}P_{(222)}(\sigma),$
-and so the last coefficient is negative if $\sigma<\tau<2\sigma$.
-When $\sigma=1$, these coefficients are falling factorials in $\tau$.
-There is a combinatorial conjecture in this case in [2].
-(It should be noted that the parameter $\alpha$ in [2] is not the same as Macdonald’s $\alpha$, instead, it is the same as our $\tau$.)
-Now, we again look at the case $n=2$. This time, the computation is more complicated.
-Example 4.7 (Toy Example for Conjecture 4).
-Let $n=2$ and $d\geqslant 2$. In this example, we mostly suppress the argument $x$, and also assume the indices $i,j=0,\dots,\lfloor d/2\rfloor$ unless otherwise stated.
-We prove the stronger statement the difference in Eq. 4.3 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. By Sections 3.2 and 3.2, we have
+Fix $0\leqslant\sigma<\tau\leqslant\infty$.
+If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference with parameter $\tau$ is positive in the Jack semiring $\mathcal{J}_{S}^{\sigma}(\mathbb{R}_{\geqslant 0})$:
+$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{J}_{S}^{\sigma}(\mathbb{R}_{\geqslant 0}).$ (4.5)
+The following example shows that Conjecture 4 is false when $\tau<\sigma$.
+Example 4.7.
+Consider $n=3$, $\tau=0$, $\sigma=1$, and let
+$F\coloneqq 3(M_{(3)}-M_{(21)})=10S_{(3)}-12S_{(21)}+2S_{(111)},$
+where $S_{\lambda}=s_{\lambda}/s_{\lambda}(\bm{1})$. Clearly, it does not belong to the Schur cone. We claim that $F$ does not belong to the Schur semiring.
+Note that in degree 3, the Schur semiring is generated by
+$f_{1}$ $\coloneqq S_{(1)}^{3},\quad f_{2}\coloneqq S_{(1)}S_{(11)},\quad f_{3}\coloneqq S_{(111)},$
+$f_{4}$ $\coloneqq S_{(1)}\bigl(S_{(2)}-S_{(11)}\bigr),\quad f_{5}\coloneqq S_{(21)}-S_{(111)},\quad f_{6}\coloneqq S_{(3)}-S_{(21)}.$
+It is not hard to see that $f_{3},f_{5},f_{6}$ are enough to generate the semiring since
+$f_{1}=f_{3}+\frac{26}{27}\,f_{5}+\frac{10}{27}\,f_{6},\quad f_{2}=f_{3}+\frac{8}{9}\,f_{5},\quad f_{4}=\frac{1}{9}\,f_{5}+\frac{5}{9}\,f_{6}.$
+However, $F=10f_{6}-2f_{5}$, hence $F$ does not lie in the Schur semiring.
+Now, we again look at the two-variable case in Section 3.3. This time, the computation is more complicated.
+Example 4.8.
+Consider $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$.
+In this example, we mostly suppress the argument $x$, and also assume the indices $i,j=0,\dots,\lfloor d/2\rfloor$ unless otherwise stated.
+We prove the stronger statement the difference in Eq. 4.5 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$.
+The transition matrix from $\left(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $(P_{\lambda}(x;\tau))$ is a diagonal matrix
+$M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)\coloneqq\left(\frac{1}{P_{(d-i,i)}(\bm{1};\tau)}\delta_{ij}\right)_{i,j}=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-i}}}\delta_{ij}\right)_{i,j}.$ (4.6)
+By Sections 3.2.1 and 3.2.1, we have
 $P_{(d-i,i)}(x;\tau)=\sum_{j}\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\cdot m_{(d-j,j)}(x),$
 hence the transition matrix from $\left(P_{\lambda}(x;\tau)\right)$ to $\left(m_{\lambda}(x)\right)$ is
-$M_{d}(P(\tau),m)\coloneqq\left(\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\right).$
+$M_{d}(P(\tau),m)\coloneqq\left(\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\right)_{i,j}.$
 For example, when $d=2,3$, we have
@@
 that is, $P_{(2,0)}(\tau)=m_{(2,0)}+\frac{2\tau}{1+\tau}m_{(1,1)}$ and $P_{(3,0)}(\tau)=m_{(3,0)}+\frac{3\tau}{1+\tau}m_{(2,1)}$.
-The transition matrix from $\left(\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $(P_{\lambda}(x;\tau))$ is a diagonal matrix
-$M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)\coloneqq\left(\frac{1}{P_{(d-i,i)}(\bm{1};\tau)}\delta_{ij}\right)=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-i}}}\delta_{ij}\right).$
-We claim that its inverse is given by
-$M_{d}\left(P(\sigma),m\right)^{-1}=\left((-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}\frac{(\sigma+i-j+1)_{{d-i-j}}(\sigma+d-i-j+1)_{{j-i}}}{(\sigma+1)_{{d-2i}}}\right).$
-(When $i=j=d/2$, the factor $\frac{d-2i}{d-i-j}$ is understood to be 1.) For example, if $d=2,3$, then
-$M_{2}(P(\tau),m)=\begin{pmatrix}1&-\frac{2\tau}{1+\tau}\\ &1\end{pmatrix},\quad M_{3}(P(\tau),m)=\begin{pmatrix}1&-\frac{3\tau}{1+\tau}\\ &1\end{pmatrix}$,
-Note that $(-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}$ is related to the Cardan polynomial $2T_{n}(x/2)$ (where $T_{n}(x)$ is the Chebyshev polynomial), see [OEIS A034807].
-We claim that the transition matrix $M_{d}\left(P(\tau),P(\sigma)\right)$ between Jack polynomials with different parameters, which is given by the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$, can be computed by
-$M_{d}\left(P(\tau),P(\sigma)\right)$ $=M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$
-$=\left(\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\right).$ (4.4)
-(When $i>j$, the entries are understood to be 0.) Note that the first claim follows from the second, since upon setting $\tau=\sigma$, Section 4.2 yields the identity matrix.
-To prove Section 4.2, it suffices to check the strictly upper diagonal entries. Let $i<j$, and compute the $(i,j)$-entry of the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$:
+Define a matrix
+$N_{d}(\sigma)=\left((-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}\frac{(\sigma+i-j+1)_{{d-i-j}}(\sigma+d-i-j+1)_{{j-i}}}{(\sigma+1)_{{d-2i}}}\right)_{i,j}.$ (4.7)
+(When $i=j=d/2$, the factor $\frac{d-2i}{d-i-j}$ is understood to be 1.) For example, when $d=2,3$, we have
+$N_{d}(\sigma)=\begin{pmatrix}1&-\frac{2\sigma}{1+\sigma}\\ &1\end{pmatrix},\quad N_{d}(\sigma)=\begin{pmatrix}1&-\frac{3\sigma}{1+\sigma}\\ &1\end{pmatrix}$.
+We claim that the product matrix $M_{d}\left(P(\tau),m\right)N_{d}(\sigma)$ is given by
+$M_{d}\left(P(\tau),m\right)N_{d}(\sigma)$ $=\left(\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\right)_{i,j}.$ (4.8)
+(When $i>j$, the entries are understood to be 0.) In particular, when setting $\sigma=\tau$, we see that the product $M_{d}\left(P(\tau),m\right)N_{d}(\tau)$ is the identity matrix, and hence $N_{d}(\tau)=M_{d}\left(P(\tau),m\right)^{-1}$ is the inverse transition matrix, and $M_{d}\left(P(\tau),m\right)N_{d}(\sigma)=M_{d}\left(P(\tau),P(\sigma)\right)$ is the transition matrix between Jack polynomials with different parameters.
+Given Eq.
+4.7, to prove Eq.
+4.8, it suffices to check the strictly upper diagonal entries.
+Let $i<j$, and compute the $(i,j)$-entry of the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$:
 $\mathrel{\phantom{=}}\sum_{k=i}^{j}\binom{d-2i}{k-i}\frac{(\tau)_{{k-i}}(\tau)_{{d-k-i}}}{(\tau)_{{d-2i}}}(-1)^{k-j}\times$
@@
 $=$ $(-1)^{i-j}\frac{(\tau)_{{d-2i}}(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}.$
-Denote by $f(i,j,d;l)$ the summand on the left, and $r(i,j,d)$ the RHS. We shall prove Section 4.2 using the so-called Wilf–Zeilberger algorithm [23, Chapter 7]. Let
+Denote by $f(i,j,d;l)$ the summand on the left, and $r(i,j,d)$ the RHS. We prove Example 4.8 using the Wilf–Zeilberger algorithm [22, Chapter 7]. Let
 $\mathrel{\phantom{=}}F(i,j,d;l)\coloneqq f(i,j,d;l)/r(i,j,d)$
@@
 $\sum_{l\geqslant 0}F(i,i+1,d;l)=-\frac{(d-2i+\tau-1)\sigma}{(d-1-2i)(\tau-\sigma)}+\frac{\tau(d-1-2i+\sigma)}{(d-1-2i)(\tau-\sigma)}=1,$
-and so for any $j>i$, $\sum_{l=0}^{j-i}F(i,j,d;l)=\sum_{l\geqslant 0}F(i,j,d;l)=1$.
-We have thus proved Section 4.2.
+and so for any $j>i$, $\sum_{l=0}^{j-i}F(i,j,d;l)=\sum_{l\geqslant 0}F(i,j,d;l)=1$. We have thus proved Eq. 4.8.
+Next, we show (as asserted just before Eq.
+4.6 that the difference in Eq.
+4.5 belongs to the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$.
 The transition matrix from $\left(\frac{P_{\lambda}(\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $\left(\frac{P_{\lambda}(\sigma)}{P_{\lambda}(\bm{1};\sigma)}\right)$ is
@@
 $\propto 2(2\tau-1)j^{2}-2(2\tau-1)(d+\sigma)j+d(d-1)(\tau-\sigma-1)\eqqcolon a_{j},$
-where $\propto$ means that a positive factor is omitted. It suffices to show that $a_{j}<0$ for $j=j_{0},\dots,\lfloor d/2\rfloor$ for some $j_{0}\geqslant 1$.
+where $\propto$ means that a factor in $\mathbb{F}_{\geqslant 0}$ has been omitted.
 Note that the last entry, $a_{\lfloor d/2\rfloor}$, is always negative: for $d=2k$ even, we have $a_{k}=-2k(\tau+k-1)(2\sigma+1)<0$; for $d=2k+1$ odd, we have $a_{k}=-2k(\tau+k)(2\sigma+1)<0$.
-Also, we have $a_{j+1}-a_{j}=-2(2\tau-1)(\sigma+d-2j-1)$.
-If $\tau\geqslant\frac{1}{2}$, $(a_{j})$ is decreasing, in particular, $a_{j}$ satisfies the desired condition.
-If $\tau<\frac{1}{2}$, $(a_{j})$ is increasing, hence $a_{j}\leqslant a_{\lfloor d/2\rfloor}<0$.
+For $j=1,\dots,\lfloor d/2\rfloor-1$, we have $a_{j+1}-a_{j}=-2(2\tau-1)(\sigma+d-2j-1)$, hence $(a_{j})$ is monotone.
+Now we have the expansion
+$\frac{P_{(d,0)}(\tau)}{P_{(d,0)}(\bm{1};\tau)}-\frac{P_{(d-1,1)}(\tau)}{P_{(d-1,1)}(\bm{1};\tau)}=\frac{P_{(d,0)}(\bm{1};\sigma)}{P_{(d,0)}(\bm{1};\tau)}\frac{P_{(d-j,j)}(\sigma)}{P_{(d-j,j)}(\bm{1};\sigma)}+\sum_{j=1}^{\lfloor d/2\rfloor}c_{j}a_{j}\frac{P_{(d-j,j)}(\sigma)}{P_{(d-j,j)}(\bm{1};\sigma)},$
+where $c_{j}>0$ denotes the factor omitted above, and $(a_{j})$ is monotone and ends at a negative number.
+In particular, there exists some $j_{0}$ such that $a_{j}\geqslant 0$ for $j<j_{0}$ and $a_{j}\leqslant 0$ for $j\geqslant j_{0}$.
+Such an expansion lies in the Jack cone by Abel’s lemma.
 This calculation leads to:
@@
 It suffices to show that it holds for any adjacent pair $\lambda=(d-i,i)\mathrel{\succcurlyeq}\mu=(d-i-1,i+1)$, as the general case would follow by telescoping.
-But this holds because
+This holds because
 $\frac{P_{(d-i,i)}(x;\tau)}{P_{(d-i,i)}(\bm{1};\tau)}-\frac{P_{(d-i-1,i+1)}(x;\tau)}{P_{(d-i-1,i+1;\tau)}(\bm{1};\tau)}$ $=(x_{1}x_{2})^{i}\cdot\left(\frac{P_{(d-2i,0)}(x;\tau)}{P_{(d-2i,0)}(\bm{1};\tau)}-\frac{P_{(d-2i-1,1)}(x;\tau)}{P_{(d-2i-1,1)}(\bm{1};\tau)}\right)$
-$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu}(x;\tau)}{P_{\nu}(\bm{1};\tau)}-\frac{P_{\xi}(x;\tau)}{P_{\xi}(\bm{1};\tau)}\right)$
-$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu+(i,i)}(x;\tau)}{P_{\nu+(i,i)}(\bm{1};\tau)}-\frac{P_{\xi+(i,i)}(x;\tau)}{P_{\xi+(i,i)}(\bm{1};\tau)}\right).$
+$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu}(x;\sigma)}{P_{\nu}(\bm{1};\sigma)}-\frac{P_{\xi}(x;\sigma)}{P_{\xi}(\bm{1};\sigma)}\right)$
+$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu+(i,i)}(x;\sigma)}{P_{\nu+(i,i)}(\bm{1};\sigma)}-\frac{P_{\xi+(i,i)}(x;\sigma)}{P_{\xi+(i,i)}(\bm{1};\sigma)}\right)\subseteq\mathcal{J}_{C}(\mathbb{R}_{\geqslant 0}).$
 Now, let $n\geqslant 3$, $\lambda=(2,1^{n-2},0)$, and $\mu=(1^{n})$. We have
@@
 ∎
-4.3. A conjecture of Alexandersson–Haglund–Wang
-For completeness, we end this section by recalling a related conjecture in [2]. Following the convention in [4, Section 6.2], we define the integral Jack polynomials as
-$J_{\lambda}(x;\tau)=c_{\lambda}(\tau)P_{\lambda}(x;\tau),\quad\text{where}\quad c_{\lambda}(\tau)\coloneqq\prod_{s\in\lambda}(a_{\lambda}(s)+\tau(l_{\lambda}(s)+1)).$ (4.7)
-For $s=(i,j)\in\lambda$, we have $a_{\lambda}(s)\coloneqq\lambda_{i}-j$ and $l_{\lambda}(s)=\lambda_{j}^{\prime}-i$ are the number of boxes direct to the East and South of $s$ (drawn in the English manner).
-Note that our $J_{\lambda}(x;\tau)$ is the same as the $\tilde{J}_{\lambda}^{(\alpha)}(x)$ in [2], with $\alpha$ replaced by $\tau$ (note: not $\frac{1}{\tau}$); and is related to the $J_{\lambda}^{(\alpha)}(x)$ in [18, VI.
-(10.22)] by $J_{\lambda}(x;\tau)=\tau^{-|\lambda|}J^{(1/\tau)}_{\lambda}(x)$.
-As shown in [12], the expansion coefficients of integral Jack into the (augmented) monomial basis are positive and integral and can be combinatorially interpreted as admissible (also known as non-attacking) tableaux.
-It was recently conjectured in [2] that the expansion coefficients of the integral Jack $J_{\lambda}(x;\tau)$ into the Schur basis $s_{\mu}$ have certain positivity and integrality. Define
-$v_{\lambda\mu}(\tau)\coloneqq\langle J_{\lambda}(\tau),s_{\mu}\rangle,\qquad\text{that is,}\qquad J_{\lambda}(\tau)=\sum_{\mu}v_{\lambda\mu}(\tau)s_{\mu}.$ (4.8)
-Conjecture 5 ([2]).
-Let $\lambda,\mu$ be partitions of $d$, and define the expansion coefficients $a_{k}(\lambda,\mu)$ and $b_{k}(\lambda,\mu)$ as follows:
-$v_{\lambda\mu}(\tau)=\sum_{k=0}^{d-1}a_{k}(\lambda,\mu)\binom{\tau+k}{d}=\sum_{k=1}^{d}b_{d-k}(\lambda,\mu)\binom{\tau}{k}k!.$ (4.9)
-Then $a_{k}(\lambda,\mu)$ and $b_{d-k}(\lambda,\mu)$ are in $\mathbb{Z}_{\geqslant 0}$.
-Furthermore, the polynomials $\sum_{k=0}^{d}a_{k}(\lambda,\mu)z^{k}$ and $\sum_{k=0}^{d}b_{d-k}(\lambda,\mu)z^{k}$ have only real zeros.
 5. Macdonald polynomial differences: conjectures and results
 Let $P_{\lambda}(x)=P_{\lambda}(x;q,t)$ denote the monic Macdonald polynomials.
-Our goal in this section is to explore if Conjecture 3, characterizing majorization in terms of (normalized) Jack differences belonging to the Muirhead semiring, can be strengthened to Macdonald polynomials – leading to a strengthening of Conjecture 1 as well.
+Our goal in this section is to explore if Conjecture 3, characterizing majorization in terms of (normalized) Jack differences belonging to the Muirhead semiring, can be strengthened to Macdonald polynomials—leading to a strengthening of Conjecture 1 as well.
 And indeed, we achieve both of these goals, in the cases where we have shown the conjectures above.
@@
 […]
-We now study Macdonald differences for $\lambda\succcurlyeq\mu$, starting with the case of two variables.
+We now study Macdonald differences for $\lambda\mathrel{\succcurlyeq}\mu$, starting with the case of two variables.
 The first question in extending the aforementioned conjectures is: how should one normalize the Macdonald polynomials.
@@
 Example 5.1.
-Let $n=2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. We have, using the combinatorial formula [18, VI. (7.13)],
+Let $n=2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. We have, using the combinatorial formula [17, VI. (7.13)],
 $P_{\lambda}(x)=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}x_{1}^{i}x_{2}^{d-i}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}.$ (5.3)
-By [18, VI. (6.11)], where $n=2$ and $t^{\delta}=(t,1)$, $P_{\lambda}(t^{\delta})=\frac{(t^{2};q)_{{d}}}{(t;q)_{{d}}}$. Hence,
+By [17, VI. (6.11)], where $n=2$ and $t^{\delta}=(t,1)$, $P_{\lambda}(t^{\delta})=\frac{(t^{2};q)_{{d}}}{(t;q)_{{d}}}$. Hence,
 $\frac{P_{\lambda}(x)}{P_{\lambda}(t^{\delta})}=\frac{(q;q)_{{d}}}{(t^{2};q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}.$
@@
 Continue with $n=2$ and consider the difference $\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$. Now
-$P_{\lambda}(\bm{1})=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}.$
+$P_{\lambda}(1,1)=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}.$
 Hence we have
-$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}\text{\quad and \quad}\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}.$
+$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}\text{\quad and \quad}\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}=\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}.$
 Their difference is
-$\mathrel{\phantom{=}}\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$
+$\mathrel{\phantom{=}}\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$
 $=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}-\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}$
@@
 Lemma 5.3.
-Let $q,t>1$.
+Let $q,t\in(0,1)$.
   1. (1)
@@
 (1) It is clear that $b_{i}>0$. Note that
-$b_{i+1}-b_{i}=\dfrac{1-tq^{i}}{1-q^{i+1}}-\dfrac{1-tq^{i-1}}{1-q^{i}}=\frac{q^{i-1}(q-t)(q-1)}{(1-q^{i})(1-q^{i+1})}.$
-Hence $(b_{i})_{i}$ is increasing if $q>t$ and decreasing if $q<t$.
+$b_{i+1}-b_{i}=\dfrac{1-tq^{i}}{1-q^{i+1}}-\dfrac{1-tq^{i-1}}{1-q^{i}}=\frac{q^{i-1}(t-q)(1-q)}{(1-q^{i})(1-q^{i+1})}.$
+Hence $(b_{i})$ is increasing if $q<t$ and decreasing if $q>t$.
 (2) We first show that the two expressions are equal.
@@
 If $q=t$, the $(2l+1)$-term polynomial becomes $1$.
-If $q>t$, then by part (1), $b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}}>0$.
-If $q<t$, then
+If $q<t$, then by part (1), $b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}}>0$.
+If $q>t$, then
 $b_{i}b_{d-i}-b_{s_{i}}b_{d-s_{i}}$ $=\frac{1-tq^{i-1}}{1-q^{i}}\frac{1-tq^{d-i-1}}{1-q^{d-i}}-\frac{1-tq^{s_{i}-1}}{1-q^{s_{i}}}\frac{1-tq^{d-s_{i}-1}}{1-q^{d-s_{i}}}$
-$=\frac{-q^{i-1}(t-q)(1-tq^{d-1})(1-q^{s_{i}-i})(1-q^{d-i-s_{i}})}{(1-q^{i})(1-q^{d-i-1})(1-q^{s_{i}})(1-q^{d-s_{i}-1})}>0.\qed$
+$=\frac{q^{i-1}(q-t)(1-tq^{d-1})(1-q^{s_{i}-i})(1-q^{d-i-s_{i}})}{(1-q^{i})(1-q^{d-i-1})(1-q^{s_{i}})(1-q^{d-s_{i}-1})}>0.\qed$
 Now, we prove $\sum_{i=0}^{l}a_{i}\geqslant 0$, for $1\leqslant l\leqslant\lfloor d/2\rfloor-1$.
@@
 $=\sum_{j=l+1}^{k}c_{j-1}c_{d-j-1}c_{d-l-1}\left(\prod_{m=d-l}^{d}b_{m}+\sum_{i=1}^{l}\prod_{m=1}^{i-1}b_{m}\cdot\left(b_{i}b_{d-i}-b_{j}b_{d-j}\right)\cdot\prod_{m=d-l}^{d-i-1}b_{m}\right).$
-As the expression within each parentheses is of the form in Eq. 5.4, it is positive. ∎
-Thus, we arrive at the following conjecture for normalized Macdonald differences in the Muirhead semiring – and we now merge this with the Macdonald analogue of the Cuttler–Greene–Skandera conjecture too.
-Recall here that $\mathbb{F}_{\geqslant 0}$ and $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ for the Macdonald case are defined in Eq.
+As the expression within each parentheses is of the form in Eq. 5.4, it is positive.
+Thus, we arrive at the following conjecture for normalized Macdonald differences in the Muirhead semiring—and we now merge this with the Macdonald analogue of the Cuttler–Greene–Skandera conjecture too.
+Let $\mathbb{K}_{\geqslant 0}$ and $\mathbb{K}_{\geqslant 0}^{\mathbb{R}}$ be the cones for the Macdonald case, as defined in Eq.
 2.12.
 Conjecture 5 (CGS and Muirhead conjectures for Macdonald polynomials).
-Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
+The following are equivalent for partitions $\lambda$ and $\mu$ :
   1. (1)
-    The Macdonald difference lies in the Muirhead semiring over $\mathbb{F}_{\geqslant 0}$:
-    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0}).$ (5.5)
+    The Macdonald difference lies in the Muirhead semiring over $\mathbb{K}_{\geqslant 0}$:
+    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{K}_{\geqslant 0}).$ (5.5)
   2. (2)
     We have
-    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.6)
+    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{K}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.6)
+  3. $(2^{\prime})$
+    For every $q_{0},t_{0}\in(0,1)$, the Macdonald difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
+    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (5.7)
   4. (3)
-    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
+    For every $q_{0},t_{0}\in(0,1)$, we have
     $\frac{P_{\lambda}(x;q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x;q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (5.8)
   5. (4)
-    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
+    For some $q_{0},t_{0}\in(0,1)$, we have
     $\frac{P_{\lambda}(x;q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x;q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in(0,1)^{n}\cup(1,\infty)^{n}.$ (5.9)
@@
 Remark 5.4.
-By the well-known duality $P_{\lambda}(x;q,t)=P_{\lambda}(x;1/q,1/t)$, one can define instead $\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }0<q,t<1\,\}$ and get an equivalent statement.
-Remark 5.5.
-For completeness, we record two additional equivalent conditions in Conjecture 6:
-  • $(1^{\prime})$
-    For all real $q,t>1$, the Macdonald difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
-    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{R}_{\geqslant 0}).$ (5.9)
-  • $(2^{\prime})$
-    For all real $q,t>1$, we have
-    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in[0,\infty),\quad\forall x\in[0,\infty)^{n}.$ (5.10)
-These come from specializing assertions (1) and (2) respectively, to scalars $q,t>1$.
-It is also clear that $(1)\implies(1^{\prime})\implies(2^{\prime})$ and $(1)\implies(2)\implies(2^{\prime})$; and $(2^{\prime})\implies(3)$.
-Moreover, the counterpart of $(2^{\prime})$ can also be added in Conjectures 1 and 2 (and in Conjecture 7); as also the counterpart of $(1^{\prime})$ in Conjecture 3.
+By the well-known duality $P_{\lambda}(x;q,t)=P_{\lambda}(x;1/q,1/t)$, one can define instead $\mathbb{K}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(1,\infty)\,\}$ and get an equivalent statement.
 As promised, we now prove Conjecture 5 in the cases where we have shown it for Jack polynomials above.
 Theorem 5.5.
-Each assertion in Conjecture 6 implies the next.
-Moreover, the implication $(5)\implies(1)$ – and hence Conjecture 6 – moreover holds for all $\lambda$, with $\mu=(1^{|\lambda|})$.
-It also holds for $n=2$ and $\lambda,\mu$ arbitrary.
+In Conjecture 5, it is clear that $(1)\Rightarrow(2)\&(2^{\prime})$, $(2)\Rightarrow(3)$, $(2^{\prime})\Rightarrow(3)$, and $(3)\Rightarrow(4)\Rightarrow(5)$.
+Hence the conjecture reduces to the implication $(5)\Rightarrow(1)$.
+Moreover, this implication holds for $\mu=(1^{|\lambda|})$; and it also holds for arbitrary $\lambda$ and $\mu$ in the case of two variables.
 Proof.
-That $(1)\implies(2)$ follows from Muirhead’s inequality, and $(2)\implies(3)\implies(4)$ are obvious.
-The proof of $(4)\implies(5)$ is via the argument used to show Theorem 3.2: if (4) holds for all $x\in(1,\infty)^{n}$ then $\lambda$ weakly majorizes $\mu$; if it holds for $x\in(0,1)^{n}$ then $\lambda\succcurlyeq_{w}-\mu$.
-Together, these are equivalent to $\lambda\succcurlyeq\mu$.
-Next, we assume $\mu={\bf 1}_{|\lambda|}$, with $n,\lambda$ arbitrary and show $(5)\implies(1)$.
-Haglund–Haiman–Loehr have shown [7] that the integral Macdonald polynomials are monomial-positive over $\mathbb{F}_{\geqslant 0}$ for $q,t\in(0,1)$, whence so are the $P_{\lambda}$.
-Thus by the duality in Section 5, $P_{\lambda}(x)$ is monomial-positive for $q_{0},t_{0}\in(1,\infty)$.
+Note that $(1)\Rightarrow(2)$ and $(2^{\prime})\Rightarrow(3)$ follows from Muirhead’s inequality; $(1)\Rightarrow(2^{\prime})$ and $(2)\Rightarrow(3)$ follow by specializing $q,t$; $(3)\Rightarrow(4)$ is obvious; $(4)\Rightarrow(5)$ is similar to Theorem 3.4 and is omitted.
+Next, we assume $\mu=(1^{|\lambda|})$, with $n,\lambda$ arbitrary and show $(5)\Rightarrow(1)$.
+Haglund–Haiman–Loehr have shown [6] that the integral Macdonald polynomials are monomial-positive over $\mathbb{K}_{\geqslant 0}$ for $q,t\in(0,1)$, whence so are the $P_{\lambda}$.
 Now the calculation in Eq.
 4.4 implies (1).
-Finally, let $n=2$ and $\lambda,\mu$ be arbitrary.
-To show $(5)\implies(1)$, it suffices via telescoping to do so for majorization-adjacent partitions $\lambda=(d-i,i)$ and $\mu=(d-i-1,i+1)$.
+Finally, let $n=2$ and $\lambda,\mu$ have length at most 2.
+To show $(5)\Rightarrow(1)$, it suffices via telescoping to do so for majorization-adjacent partitions $\lambda=(d-i,i)$ and $\mu=(d-i-1,i+1)$.
 We showed this in Section 5 for $i=0$; and we use this to compute for $i>0$:
 $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}$ $=(x_{1}x_{2})^{i}\left(\frac{P_{(d-2i,0)}(x;q,t)}{P_{(d-2i,0)}(\bm{1};q,t)}-\frac{P_{(d-2i-1,1)}(x;q,t)}{P_{(d-2i-1,1)}(\bm{1};q,t)}\right)$
-$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu}(x)-M_{\xi}(x))$
-$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu+(i,i)}(x)-M_{\xi+(i,i)}(x)).\qed$
+$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{K}_{\geqslant 0}\cdot(M_{\nu}(x)-M_{\xi}(x))$
+$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{K}_{\geqslant 0}\cdot(M_{\nu+(i,i)}(x)-M_{\xi+(i,i)}(x))\subseteq\mathcal{M}_{C}(\mathbb{K}_{\geqslant 0}).\qed$
 Remark 5.6.
@@
 As Section 4.1 shows, this is false in general for three variables.
-We conclude with two other conjectures, which “upgrade” weak majorization and containment from Jack to Macdonald polynomials.
-Conjecture 7 (KT Conjecture for Macdonald polynomials).
-The following are equivalent:
-  1. (1)
-    We have
-    $\frac{P_{\lambda}(x+\bm{1};q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x+\bm{1};q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.11)
-  2. (2)
-    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
-    $\frac{P_{\lambda}(x+\bm{1};q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x+\bm{1};q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (5.12)
-  3. (3)
-    $\lambda$ weakly majorizes $\mu$.
-As above, we now provide some positive evidence for this conjecture.
-Theorem 5.8.
-Each assertion in Conjecture 7 implies the next. Moreover, Conjecture 7 holds for all $n,\lambda$ when $\mu=(1^{m})$ for some $m\geqslant 1$.
-Proof.
-That $(1)\implies(2)$ is immediate, and that $(2)\implies(3)$ again follows via the argument used to show Theorem 3.2.
-Next, let $\lambda\succcurlyeq_{w}\mu$, and let $\mu^{\prime}={\bf 1}_{|\lambda|}$; then $\lambda\succcurlyeq\mu^{\prime}\supseteq\mu$.
-Now directly compute (suppressing $q_{0},t_{0}$ from the arguments):
-$\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}=\left(\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}\right)+\left(\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}\right).$
-At each $x\in[0,\infty)^{n}$, the first difference on the right is in $\mathbb{F}_{\geqslant 0}$ by Theorem 5.6, while the second is the scalar $M_{\mu^{\prime}}(x+{\bf 1})-M_{\mu}(x+{\bf 1})$, where both terms are positive reals.
-Now compute their ratio:
-$\frac{M_{\mu^{\prime}}(x+{\bf 1})}{M_{\mu}(x+{\bf 1})}=\prod_{i=1}^{n}(x_{i}+1)^{\mu^{\prime}_{i}-\mu_{i}},$
-and this is at least $1$ since all $x_{i}\in[0,\infty)$ and $\mu^{\prime}\supseteq\mu$. ∎
-Conjecture 7 is not contingent upon our next and final conjecture, but the proof of its counterpart for Jack polynomials used a characterization of containment via Jack polynomials. Thus, we end with:
-Conjecture 8 (Containment via Macdonald differences).
-The Macdonald-counterpart over $\mathbb{F}_{\geqslant 0}$ of Theorem 1.1 holds.
 Acknowledgments